### v4

# Superquantile-Gibbs Relaxation for Minima-selection in Bilevel Optimization

Abstract
Bilevel optimization (BLO) becomes fundamentally more challenging when the lower-level objective admits multiple minimizers. In contrast to the commonly studied unique-minimizer setting, this regime introduces two additional difficulties: (1) evaluating the hyper-objective $F_{\max}$ requires a minima-selection step, which, in the worst case, amounts to solving an optimization problem over a topologically disconnected set; (2) $F_{\max}$ is in general discontinuous without strong structural assumptions. We show that both issues can be circumvented under a local Polyak–Łojasiewicz (PŁ) property, denoted by , on the lower-level objective. This condition is strictly weaker than the global PŁ property assumed in recent BLO studies and covers problems such as hyperparameter tuning in over-parameterized deep learning. Under , we prove that $F_{\max}$ is Lipschitz continuous and that, for all upper-level variables, the corresponding lower-level minimizer sets are topologically connected and are provably closed embedded submanifolds, all sharing a common intrinsic dimension k. We show that this intrinsic dimension k, rather than the ambient dimension, governs the complexity of BLO: We propose a method that finds an $(\epsilon,\rho)$-Goldstein stationary point of $F_{\max}$ using at most $\tilde{\mathcal{O}}\big(m^{8\texttt{k}+11}\epsilon^{-2}(\epsilon\rho)^{-8\texttt{k}-10}\big)$ queries to the gradient oracle of the lower-level function, where $m$ is the dimension of the upper-level domain. The key ingredient is a Superquantile–Gibbs relaxation that yields an approximate function-evaluation oracle for $F_{\max}$: it turns minima selection into a sampling problem that can be efficiently addressed using standard Langevin dynamics. To the best of our knowledge, this is the first work that rigorously tackles the minima selection step in BLO and explicitly characterizes how the overall complexity of BLO scales with the intrinsic dimensionality of the lower-level problem.
Keywords: Bilevel optimization, Nonconvex lower-level problem, Local PŁ condition, and Nonsmooth optimization.
1 Introduction
We study the pessimistic/optimistic formulation of Bilevel Optimization (BLO), which is a common paradigm in the literature [90, 87, 23, 62, 84], although alternatives to pessimistic/optimistic formulation for BLO also exist (see e.g., [4, 43, 11]). The pessimistic formulation of BLO is defined as follows [^1] All results extend to the optimistic setting where minimizing is considered instead of maximizing $f$ over $x$.:
$\min_{\theta\in\Theta\subset\operatorname{\mathbb{R}}^{m}}F_{\max}(\theta):=\left\{\max_{x\in\mathcal{S}(\theta)\subseteq\mathbb{R}^{d}}f(\theta,x)\right\},\quad\mathcal{S}(\theta):=\operatorname*{arg\,min}_{x\in\mathbb{R}^{d}}g(\theta,x),$ (1)
where $\theta$ and $x$ are the upper- and lower-level variables, $f$ and $g$ are the upper- and lower-level objectives, respectively, and $\Theta$ is a closed compact convex subset of $\mathbb{R}^{m}$. We focus on the setting where the lower-level minima set $\mathcal{S}(\theta)$ is non-singleton. We refer to the maximization of $f$ over $x\in\mathcal{S}(\theta)$ as minima-selection and call $F_{\max}$ the hyper-objective. BLO captures problems with a nested structure, such as hyper-parameter optimization [28], meta-learning [8], and reinforcement learning [39, 61]. In particular, when the lower-level problem pertains to deep learning, the set $\mathcal{S}(\theta)$ is typically nonconvex and non-singleton [58, 48, 69, 31, 27, 55, 80, 75]. Therefore, minima-selection becomes important.
Relaxing strong convexity to the global PŁ condition
Recent works on BLO have studied nonconvex lower-level problems under the global Polyak-Łojasiewicz (PŁ) condition [14, 50, 83, 40, 77, 59]. This condition relaxes the classical strong convexity assumption on $g$, which implies singleton $\mathcal{S}(\theta)$, to cover nonconvex and potentially non-singleton optimal solution sets. Another motivation for adopting the global PŁ condition is to guarantee the continuity of the hyper-objective function $F_{\max}$, which might otherwise become discontinuous and render BLO problems generally intractable [50, 14].
All of the aforementioned work and their results are crucially dependent on the assumption that $g$ is $\mathcal{C}^{2}$ and satisfies the global PŁ condition. Under these conditions, it can be proved that the lower-level solution set $\mathcal{S}(\theta)$ would be limited to either a singleton or an unbounded subset of $\mathbb{R}^{d}$ [20]. More specifically, Huang [40] assume a nondegenerate Hessian of $g$ at global minimizers, implicitly enforcing uniqueness; Liu et al. [59] explicitly assume $\mathcal{S}(\theta)$ is a singleton; Xiao et al. [83] and Shen et al. [77, Sec. 4] assume additional convexity of the lower-level objective, implying that the minimizer set is a linear subspace; and Kwon et al. [50] require $g$ to be coercive which again yields a singleton solution [^2] Kwon et al. [50] assume the proximal error bound for $g$ (with $\sigma=0$), equivalent to the global PŁ condition for smooth $g$ [14, Prop. C.1].. The above observations hint at the necessity of a further relaxation to capture richer landscapes.
Our setting: Relaxing the global PŁ condition to a local PŁ condition
We relax the global PŁ assumption to the condition, which was recently proposed by [35], to accommodate applications of interest such as those in deep learning. For a fixed $\theta$, a lower-level objective $g(\theta,\cdot)$ admits a provably connected optimal set $\mathcal{S}(\theta)$ and it satisfies the PŁ inequality locally in a neighborhood of $\mathcal{S}(\theta)$. Moreover, the condition ensures that $\mathcal{S}(\theta)$ is a $\mathcal{C}^{2}$ embedding submanifold of the ambient space $\operatorname{\mathbb{R}}^{d}$ without boundary [35]. We shall define condition formally in Section˜2. This setting is sufficiently rich because the lower-level problem admits bounded non-singleton minimizers. Hence, the minima-selection step remains relevant. One such example is when $g$ represents the loss function of some deep learning tasks in the over-parameterized regime, where the local minima are observed to be connected [27, 31, 69, 48, 58]. Another advantage of this setting is that, as we prove in Section˜3.1, the hyper-objective under the condition is Lipschitz continuous. However, under the same condition, the hyper-objective function is not necessarily differentiable, and conditions on functions $f$ and $g$ that would ensure the differentiability of the hyper-objective are often restrictive and unrealistic in practice (see the discussion in Section˜A.2).
Solution concept for nonsmooth nonconvex optimization.
Solving BLO problems with a non-singleton lower-level solution set $\mathcal{S}(\theta)$ involves two main challenges: 1) As previously emphasized, for a fixed upper-level variable $\theta$, the minima-selection problem requires maximization over a nonconvex set. Consequently, exact evaluations of the hyper-objective $F_{\max}$ are generally inaccessible. 2) Even when the minima-selection subproblem can be solved exactly, the resulting hyper-objective $F_{\max}$ may still be non-differentiable with respect to the upper-level variable $\theta$. Addressing this issue requires establishing local optimality guarantees for a nonsmooth nonconvex objective, a widely recognized challenge in the literature [46, 44, 78, 47, 93].
To address these issues, we introduce the superquantile–Gibbs relaxation to approximate the hyper-objective, providing a practical surrogate for $F_{\max}$. Using evaluations of this surrogate, we propose a zeroth-order method that seeks a generalized Goldstein stationary point of $F_{\max}$ (see Section˜5.2; cf. [64]).
1.1 Finding a Goldstein Stationary Point of the Hyper-objective
As we shall see in ˜4.5, under the condition the hyper-objective $F_{\max}$ is Lipschitz continuous. A natural choice for approximate stationarity in this setting is Clarke’s notion of stationarity [18]. However, Zhang et al. [93] showed that, for nonsmooth Lipschitz problems, obtaining an $\epsilon$-Clark stationary point in polynomial time is generally impossible, motivating more nuanced concepts. We therefore adopt the notion of $(\epsilon,\rho)$-Goldstein stationarity [34].
For the domain-constrained objective in (1), akin to [64], we measure stationarity by a generalized $(\epsilon,\rho)$-Goldstein criterion which is defined for the gradient mapping [^3] The gradient mapping for $F:\Theta\to\mathbb{R}$ at a point $\theta$ is $\mathcal{G}_{\Theta}(\theta,\nabla F;\eta):=\eta^{-1}(\theta-\text{Proj}_{\Theta}(\theta-\eta\nabla F(\theta)))$., denoted by $(\epsilon,\rho,\eta)$-generalized Goldstein stationary point (see Section˜5.2). As $F_{\max}$ may not be differentiable, we propose a zeroth-order (ZO) method to find such a stationary point of $F_{\max}$.
The major obstacle to applying ZO methods to minimize $F_{\max}$ is that exact function values of this hyper-objective are typically unavailable. To address this, we introduce a tractable surrogate, the superquantile-Gibbs $\tilde{F}_{{\texttt{SQ-G}}}(\theta)$, which we define next and show that it approximates $F_{\max}$ well. Consequently, querying $\tilde{F}_{{\texttt{SQ-G}}}$ in place of $F_{\max}$ provides the finite-difference gradient estimates for our zeroth-order (ZO) method. We prove that the approximation error incurred by replacing $F_{\max}$ with $\tilde{F}_{{\texttt{SQ-G}}}$ does not adversely affect the convergence guarantees of our ZO algorithm for obtaining an $(\epsilon,\rho,\eta)$-Goldstein stationary point.
1.2 Superquantile-Gibbs Relaxation (SQ-G)
In this part, we introduce the superquantile of the random variable $f(\theta,X)$ where $X$ is sampled according to the the Gibbs measure $\sim\exp(-g(\theta,x)/\lambda)$. Our idea is that as $\lambda\rightarrow 0$, the Gibbs measure concentrates on the feasible domain $\mathcal{S}(\theta)$, and the superquantile at a high confidence level would approach the maximum value of $f(\theta,\cdot)$ on $\mathcal{S}(\theta)$. Formally, we define the Superquantile-Gibbs Relaxation (SQ-G)-loss using the variational formulation of the superquantile by [74]:
$(\text{{SQ-G}-loss})$ $\ \tilde{F}_{\texttt{SQ-G}}(\theta)=\min_{\beta\in\mathbb{R}}\phi_{\lambda,\delta}(\theta,\beta)$ (2) with $\ \phi_{\lambda,\delta}(\theta,\beta)=\beta+\frac{1}{\delta}\mathbb{E}_{X\sim\mu^{\lambda}_{\theta}}[\max\{f(\theta,X)-\beta,0\}].$
Here $\lambda>0$ and $\delta>0$ are parameters controlling the level of SQ-G relaxation, with $(1-\delta)$ known as the confidence level, and $\mu^{\lambda}_{\theta}$ denotes the Gibbs measure:
$\mu^{\lambda}_{\theta}(x):=\frac{\exp({-{g(\theta,x)}/{\lambda}})}{Z_{\lambda}(\theta)},$ (3)
where $Z_{\lambda}(\theta):=\int_{\mathbb{R}^{d}}\exp({-{g(\theta,z)}/{\lambda}})dz$ is a normalizing factor. A key contribution of this project is that we characterize how $\lambda$ and $\delta$ impact the accuracy with which $\tilde{F}_{\texttt{SQ-G}}(\theta)$ approximates $F_{\max}$.
1.3 Our Contributions
Our main contributions are as follows.
  • $\bullet$
    To allow for lower-level optimal sets with rich structures, we relax the global PŁ condition, commonly imposed on the lower-level objective, to the weaker condition from [35]. We show that the condition is sufficient to ensure Lipschitz continuity of the hyper-objective function $F_{\max}$.
  • $\bullet$
    We use the Superquantile–Gibbs relaxation (SQ-G) to approximate $F_{\max}$. This converts the crucial minima-selection step in BLO, originally a nonconvex optimization over a nonconvex set, into sampling from a Gibbs distribution, which we implement efficiently via Langevin dynamics.
  • $\bullet$
    Building on the manifold structure of the lower-level optimal set, we derive the following results.
    Result 1 (Approximation).
    Assume that $g$ satisfies the condition (to be stated in Section˜2), along with standard regularity conditions such as continuity of the gradient and Hessian for both $f$ and $g$. As shown in Section˜4, this implies the approximation bound $|F_{\max}(\theta)-\tilde{F}_{{\texttt{SQ-G}}}(\theta)|=\mathcal{O}\left(\delta^{\frac{1}{\texttt{k}}}+\delta^{-1}\lambda^{1/2}\right).$ Here k is the common [^4] In Section 3.2, we show that he manifolds $\{\mathcal{S}(\theta)\}_{\theta\in\Theta}$ admit a common dimension k. dimension of the embedding submanifolds $\{\mathcal{S}(\theta)\}_{\theta\in\Theta}$.
    Result 2 (Optimization).
    Assume that $g$ satisfies and the regularity conditions for $f$ and $g$ (to be stated in Section˜2). As shown in Section˜5, by using a Projected Stochastic Zeroth-Order method with Minima-Selection (see Algorithm˜2) to optimize $F_{\max}$, we obtain an $(\epsilon,\rho,\eta)$-generalized Goldstein stationary point of $F_{\max}$ on average using $\tilde{\mathcal{O}}\big(m^{2\texttt{k}+5}\epsilon^{-2}(\epsilon\rho)^{-2\texttt{k}-4}\big)$ samples from the distribution $\mu^{\lambda}_{\theta}$ for $\texttt{k}\geq 1$. When the samples of $\mu^{\lambda}_{\theta}$ are approximated by the Langevin Monte Carlo method, we show in Section˜5.3 that the same goal can be achieved with $\tilde{\mathcal{O}}\big(m^{8\texttt{k}+11}\epsilon^{-2}(\epsilon\rho)^{-8\texttt{k}-10}\big)$ queries to a gradient oracle for $g$ and $\tilde{\mathcal{O}}\big(m^{2\texttt{k}+5}\epsilon^{-2}(\epsilon\rho)^{-2\texttt{k}-4}\big)$ queries to a value oracle for $f$.
    The precise statements of these results appear in ˜4.5, ˜5.5 and ˜5.7, respectively.
  • $\bullet$
    We evaluate our approach on both synthetic and real-world benchmarks and demonstrate the effect of minima-selection. As shown in Section˜6, our proposed approach outperforms hyper-gradient and penalty-based baselines.
Remark 1.1.
(Exponential dependence on k) The $\epsilon^{-\texttt{k}}$ factor in our results reflects the intrinsic difficulty of the problem. Since evaluating $F_{\max}(\theta)=\max_{y\in S(\theta)}f(\theta,y)$ is equivalent to solving a nonconvex optimization problem over a nonconvex set, a single $\epsilon$-accurate evaluation of $F_{\max}(\theta)$ yields an $\Omega(\epsilon^{-\texttt{k}})$ complexity lower bound. More specifically, consider the instance where $S(\theta)=\mathbb{S}^{\texttt{k}}$ is a hypersphere. Reparameterizing $y$ in polar coordinates yields $F_{\max}(\theta)=\max_{x\in\mathbb{B}_{\texttt{k}}}f(\theta,y(x))$. Here $\mathbb{B}_{\texttt{k}}=[0,\pi]^{\texttt{k}}$ and the point $y(x)\in\mathbb{S}^{\texttt{k}}$ is given by $y_{i}(x)=\sin(x_{1})\cdots\sin(x_{i-1})\cos(2x_{i})$ for $i\in[\texttt{k}]$ and $y_{\texttt{k}+1}(x)=\sin(x_{1})\cdots\sin(x_{\texttt{k}-1})\sin(2x_{\texttt{k}})$. It follows from [67, Section 1.6.2] that the worst-case complexity of finding an $\epsilon$-accurate solution to is $\Omega(\epsilon^{-\texttt{k}})$ for $f(\theta,\cdot)\in\mathcal{C}^{1}$.
1.4 Related Work
Bilevel optimization has a long history in mathematical programming; see, e.g., the early bibliographic review by [81], the monograph [22], the survey [19], the classical text on mathematical programs with equilibrium constraints (MPECs) by [66], and the recent collection [25]. Classical approaches often reduced bilevel problems to single-level programs, either via value-function reformulations or by replacing the lower-level problem with its Karush–Kuhn–Tucker (KKT) optimality system. A related value-function-based reduction for bilevel programs with nonconvex lower-level objectives yields single-level formulations with nonsmooth inequality constraints, for which smoothing and augmented-Lagrangian type algorithms were developed in, e.g., [56, 84]. The latter route yields MPEC formulations in which complementarity conditions between multipliers and constraint violations appear as explicit equilibrium constraints; see, for example, [89, 88, 26, 91]. In practice, these complementarity constraints are frequently handled through mixed-integer or so-called big-$M$ linearizations; however, their performance can be sensitive to the choice of constants, with overly small values excluding feasible bilevel solutions and overly large values leading to weak relaxations and numerical instability [45]. Complementing these algorithmic considerations, recent work has also highlighted inherent geometric and computational hardness in bilevel programming, even under smoothness, which motivates the search for additional structural assumptions and tailored relaxations [12].
Gradient-based bilevel methods with strongly convex lower level.
Modern large-scale machine-learning applications led to gradient-based methods for bilevel empirical risk minimization and hyperparameter optimization. When the lower-level objective is strongly convex in the inner variable and $f,g$ are smooth, the hyper-objective is differentiable and its gradient can be obtained by implicit differentiation. Representative works in this setting include approximate hyper-gradient methods such as HOAG [70], stochastic bilevel schemes [32, 85], and complexity-refined algorithms for empirical risk minimization [21, 15], as well as more abstract treatments of optimization layers via automatic differentiation [1], amortized implicit differentiation [3], and modular implicit differentiation frameworks [10, 37]. In these approaches, computing exact hyper-gradients typically required Hessian or Hessian-inverse information for the lower-level problem, which could be prohibitive in large neural networks. This motivated a substantial line of Hessian-free or fully first-order methods that approximated the hyper-gradient using only first-order information, including BOME [59], Hessian/Jacobian-free stochastic methods with improved sample complexity [86], fully first-order stochastic approximation methods such as F2BO [49], and bilevel stochastic-gradient methods that also accommodate nonlinear lower-level constraints and inexact hyper-gradient computations [33]. A different line of work revisits bilevel learning from a functional viewpoint: Petrulionytė et al. [72] treat the lower-level variable as a prediction function (rather than a parameter vector) and optimize it over a convex function space. Assuming strong convexity in this functional variable yields a unique lower-level solution and enables implicit differentiation via a functional implicit-function theorem, shifting the strong-convexity requirement from parameters to function space. These developments made bilevel optimization scalable in settings with a unique lower-level solution, so the minima-selection problem that we study did not arise.
Relaxing strong convexity to global PŁ.
A complementary line of work relaxed lower-level strong convexity to a global Polyak–Łojasiewicz (PŁ) condition and developed bilevel gradient methods under this assumption [14, 50, 83, 40, 77, 59]. To obtain a smooth and well-behaved hyper-objective, these methods typically combined the global PŁ condition with additional regularity assumptions on the lower-level problem (as discussed earlier), so that minima-selection was effectively avoided. In contrast, our analysis is based on a local PŁ condition and explicitly exploits the geometry of nontrivial minimizer manifolds. In a complementary direction, Chen et al. [13] introduce a set-smoothness property of the lower-level solution mapping and show that it yields weak convexity/concavity of the induced optimistic/pessimistic hyper-objectives, enabling the computation of approximate Clarke hyper-stationary points. Their analysis relies on assumptions such as a uniform error bound together with smoothness, which imply a global PŁ condition for the lower-level problem. When the lower-level objective is $\mathcal{C}^{2}$ in the lower-level variable, this global PŁ property forces $\operatorname*{arg\,min}_{x}g(\theta,x)$ to be either a singleton or an unbounded set [20]. As a consequence, these conditions rule out the bounded non-singleton minimizer manifolds that motivate the minima-selection regime studied in our work.
Alternative formulations of BLO.
Beyond such regularity-based assumptions (strong convexity or global PŁ), recent works proposed alternative formulations of BLO that target nonconvex lower-level problems. For instance, Bolte et al. [11] worked under a (piecewise) Morse-type parametric qualification, showed that the graphs of lower-level critical points and local minima are generically decomposed into finitely many smooth manifolds, and used this structure to define branch-wise hyper-objectives and study gradient methods that updated the upper-level variable along individual branches. Arbel and Mairal [4] introduced bilevel games with selection maps, where the selection map chooses a specific lower-level critical point for each upper-level parameter (e.g., the limit of a gradient flow from a given initialization). They showed that popular unrolled/iterative-differentiation bilevel algorithms with a finite inner budget approximate equilibria of these selection-based games up to approximation errors, and they extended implicit differentiation to characterize the selection and proposed a correction that removes the finite-budget error. Jiang et al. [43] further argued that, for general nonconvex lower levels, the classical hyper-objective may be ill-posed when global minimizers are unattainable; they proposed a correspondence-driven hyper-objective that evaluated the upper-level loss at the output of a prescribed lower-level algorithm, applied Gaussian smoothing to this mapping, and developed a projected stochastic gradient method with oracle-complexity guarantees for the smoothed problem. Another recent line develops Moreau-envelope single-level reformulations as an alternative to classical value-function and KKT-based complementarity-constrained reductions. Bai et al. [5] derive directional necessary optimality conditions in this framework, Gao et al. [30] propose a difference-of-weakly-convex reformulation with a sequentially convergent algorithm, and Lu [65] builds on these ideas to obtain a two-sided smoothed primal–dual method for general nonconvex bilevel problems with KKT-type convergence guarantees. In contrast, our work retains the classical (pessimistic or optimistic) hyper-objective, combines it with a local PŁ condition to obtain a manifold structure for the lower-level minimizers, and uses a superquantile–Gibbs relaxation to convert the minima-selection into sampling from a Gibbs measure.
Pessimistic bilevel optimization.
The approaches discussed above treat lower-level non-uniqueness either by enforcing regularity conditions (strong convexity or global PŁ) or by embedding explicit selection mechanisms (Morse branches, selection maps, or algorithmic correspondences), and thus remain essentially in an optimistic or selection-based setting. By contrast, the pessimistic formulation evaluates each upper-level decision against the worst upper-level value over all lower-level minimizers, which is natural when the follower may act adversarially or unreliably [6]. This risk-averse viewpoint was studied in mathematical programming, including structural and complexity results for general pessimistic models [82], surveys of modelling and solution schemes [60], and necessary optimality conditions based on generalized differentiation and value-function techniques [24]. On the algorithmic side, several works constructed single-level relaxations of the pessimistic problem: Zeng and Ye [92] designed a tight relaxation and a computation scheme for (mainly linear) pessimistic bilevel problems; Benchoouk et al. [7] analyzed a Scholtes-type relaxation of the KKT system; and Antoniou et al. [2] proposed a $\delta$-perturbed formulation that enforced an approximate unique follower response and derived error bounds between the perturbed and original pessimistic problems. Pessimistic ideas also appeared in machine learning: Ustun et al. [79] formulated hyperparameter tuning as a pessimistic bilevel problem over a finite set of tasks and highlighted advantages of this formulation over the optimistic counterpart, and Jiang et al. [42] interpreted adversarial training as a bilevel game in which a learned attacker played the role of the follower. All these approaches, however, either (i) resolved non-uniqueness by modifying the lower-level problem (e.g., $\delta$-perturbations, KKT relaxations) or (ii) specialized to finite or discrete follower responses, rarely addressed the continuous pessimistic setting with a non-singleton solution set. In contrast, we work directly with the pessimistic hyper-objective $F_{\max}$ under a local PŁ condition, model pessimism via a superquantile risk over Gibbs samples on the minimizer manifolds, and obtain complexity results that explicitly depend on the intrinsic dimension k of $\mathcal{S}(\theta)$ rather than on the ambient dimension.
1.5 Organization
The remainder of the paper is organized as follows. Section˜2 introduces the assumption and additional regularity conditions, and discusses their implications for the structure of the solution set as well as for the convergence of the Gibbs measure to its limit. Building on this, Section˜3 develops uniform geometric properties of the minimizer manifolds $\{\mathcal{S}(\theta)\}_{\theta\in\Theta}$, showing in particular that the solution mapping is locally Lipschitz and that these manifolds share a common intrinsic dimension. Section˜4 shows that, under our assumptions, the SQ-G relaxation provides a sharp approximation of $F_{\max}$ (see Section˜4.2). Using this approximation result, in Section˜5 we show our main complexity guarantees: ˜5.5 and ˜5.7 show that the complexity of optimizing $F_{\max}$ via a zeroth-order algorithm (see Algorithm˜2) scales with the common intrinsic dimension of the lower-level solution manifolds. Finally, Section˜6 presents our numerical results: we first study the dependence on the Riemannian dimension in a synthetic example and then compare our method with several competitive baselines on both a toy problem and a data hyper-cleaning task.
1.6 Notations
Unless stated otherwise, $\|\cdot\|$ denotes the Euclidean ($\ell_{2}$) vector norm and the associated operator norm on matrices and higher-order tensors. For a random variable $Z$, $\mathrm{supp}(Z)$ denotes its support. For $d\in\mathbb{N}$, $x\in\mathbb{R}^{d}$, and $r>0$, we write $\mathbb{B}_{d}(x;r):=\{y\in\mathbb{R}^{d}:\|y-x\|\leq r\}$ for the closed $d$-dimensional Euclidean ball with center $x$ and radius $r$. For $p\geq 1$ and probability measures $\mu,\nu$ on $\mathbb{R}^{d}$ with finite $p$-th moments, the $p$-Wasserstein distance is
$\mathbb{W}_{p}(\mu,\nu):=\left[\inf_{\pi\in\Pi(\mu,\nu)}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|x-y\|^{p}\,\mathrm{d}\pi(x,y)\right]^{1/p},$
where $\Pi(\mu,\nu)$ is the set of all couplings of $\mu$ and $\nu$. In particular, $\mathbb{W}_{1}$ is the 1-Wasserstein distance. Given a twice differentiable function $f:\mathbb{R}^{m}\times\mathbb{R}^{d}\to\mathbb{R}$, we write $\partial_{\theta}f(\theta,x)$ and $\partial_{x}f(\theta,x)$ for the gradients of $f$ with respect to its first and second arguments, respectively. The mixed derivative $\partial^{2}_{\theta,x}f(\theta,x):=\partial_{\theta}(\partial_{x}f(\theta,x))$ denotes the Jacobian of the vector $\partial_{x}f(\theta,x)$ with respect to $\theta$, and $\partial^{2}_{\theta}f(\theta,x)$ denotes the Hessian of $f$ with respect to $\theta$. The diameter of a set $\Theta\subset\mathbb{R}^{d}$, is defined as $\text{diam}(\Theta):=\sup_{\theta_{1},\theta_{2}\in\Theta}\|\theta_{1}-\theta_{2}\|$. The class of continuous real-valued functions on $\mathbb{R}^{d}$ is denoted by $\mathcal{C}(\mathbb{R}^{d},\mathbb{R})$. We write $f\in\mathcal{C}^{k}$ if $f$ has continuous derivatives up to order $k$. For a set $A\subset\mathbb{R}^{d}$, we denote its closure by $\overline{A}$. For a scalar random variable $Z$ and $\delta\in(0,1)$, the $(1-\delta)$-quantile is defined as
$\text{Q}_{1-\delta}(Z):=\inf\{t\in\mathbb{R}:\mathbb{P}[Z>t]\leq\delta\}.$
For a closed convex set $\Theta\subset\mathbb{R}^{m}$ and $y\in\mathbb{R}^{m}$, we denote by $\text{Proj}_{\Theta}(y)=\operatorname*{arg\,min}_{z\in\Theta}\|y-z\|$ the (Euclidean) projection of $y$ onto $\Theta$. We use the notation $\mathcal{O}$ to hide constants, and the notation $\tilde{\mathcal{O}}$ to hide both constants and logarithmic factors. We write $a\lesssim b$ to indicate that $a\leq cb$ for some universal positive constant $c>0$ that is independent of the problem parameters. We write $\mathbb{R}_{\alpha}(P\|Q)$ for the Rényi divergence of order $\alpha\in(0,\infty)\setminus\{1\}$, defined by
$\mathbb{R}_{\alpha}(P\|Q)=\begin{cases}\displaystyle\frac{1}{\alpha-1}\log\!\int\!\Big(\frac{\mathrm{d}P}{\mathrm{d}Q}\Big)^{\alpha}\,\mathrm{d}Q,&P\ll Q,\\[4.30554pt] +\infty,&\text{otherwise.}\end{cases}$
We denote $\mathrm{Law}(X)$ as the probability distribution (or law) of a random variable $X$.
2 Summary of Assumptions and Preliminaries
In this work, we focus on the setting where the lower-level objective is coercive and satisfies a local Polyak-Łojasiewicz () condition for all upper-level variables $\theta\in\Theta$, formally stated as follows.
Definition 2.1.
Let $\mathcal{M}$ be the collection of all local minima of a function $\in\mathcal{C}^{1}(\mathbb{R}^{d},\mathbb{R}). gg$ is locally PŁ if there exists a constant $\mu>0$ such that for any connected component $\mathcal{M}^{\prime}\subseteq\mathcal{M}$, there exists an open neighborhood $\mathcal{N}(\mathcal{M}^{\prime})$ containing $\mathcal{M}^{\prime}$ ($\mathcal{N}(\mathcal{M}^{\prime})\supset\mathcal{M}^{\prime}$) such that
$\forall x\in\mathcal{N}(\mathcal{M}^{\prime}),\quad g(x)-\min_{x^{\prime}\in\mathcal{N}(\mathcal{M}^{\prime})}g(x^{\prime})\leq(2\mu)^{-1}\|\nabla g(x)\|^{2}.$ (4)
Definition 2.2.
A function $g:\operatorname{\mathbb{R}}^{d}\rightarrow\operatorname{\mathbb{R}}$ satisfies the condition if the following hold.
  • $\bullet$
    $g$ is a $\mathcal{C}^{2}$ and locally PŁ function, and $g$ is a $\mathcal{C}^{4}$ function in every neighborhood $\mathcal{N}(\mathcal{M}^{\prime})$ [^5] Gong et al. [35] require $g\in\mathcal{C}^{3}$. Here we slightly enhance the assumption to $\mathcal{C}^{4}$ to enable Section 2.1.2..
  • $\bullet$
    Let $\mathcal{N}(\mathcal{M}):=\cup_{\mathcal{M}^{\prime}}\mathcal{N}(\mathcal{M}^{\prime})$. For any $x\in\mathbb{R}^{d}\setminus\mathcal{N}(\mathcal{M})$, if $\nabla g(x)=0$, then $\nabla^{2}g(x)\prec 0$.
  • $\bullet$
    The collection of all local minima $\mathcal{M}$ is contained within a compact set.
A fundamental implication of the condition is the connectivity of the optimal set.
Proposition 2.3 (Connectivity, Proposition 3 in [35]).
For a function, its local minima set $\mathcal{M}$ is path-connected.
Consequently, all local minima of a function are global minima which are all connected. Further, Rebjock and Boumal [73] proved the local manifold structure of $\mathcal{M}$. This was later extend to a global version in [35], where they also provide an explicit statement about the boundary.
Remark 2.4 (No-saddle point property).
The second condition in Section˜2 is used solely to establish the connectedness of the local minima set via the mountain pass theorem (see the proof of Proposition 3 in [35] for details). Our results remain valid if this condition is replaced with the assumption that $\mathcal{S}(\theta)$ is connected. Empirically, such connectedness has been widely observed in over-parameterized deep learning settings; see, e.g., [27, 69, 48].
Proposition 2.5 (Manifold structure, Corollary 1 in [35]).
For a function, its optimal set $\mathcal{M}$ is a compact $\mathcal{C}^{2}$ embedding submanifold of the ambient space $\operatorname{\mathbb{R}}^{d}$ without boundary.
Assumption 2.6.
For every $\theta\in\Theta$, the lower-level objective $g(\theta,\cdot)$ satisfies the condition.
Following [35], we make the next assumption to exclude the manifolds with ill-conditioned charts. We also use it to establish a nonzero lower bound on the injectivity radius, allowing us to analyze the volume-radius relationship of geodesic balls, which is crucial for our approximation results.
Assumption 2.7 (Second-order fundamental form).
For every $\theta\in\Theta$, the manifold $\mathcal{S}(\theta)$ has a well-defined second fundamental form in the sense of Section˜A.1 in Section˜A.1, and its operator norm is uniformly bounded by a constant $\mathsf{C}>0$.
We further need the following regularity conditions, which are standard in the BLO literature.
Assumption 2.8.
The following statements hold for all $\theta\in\Theta$. $\bullet$ $f$ is a $\mathcal{C}^{1}$ function and $L_{f,1}$-Lipschitz, and $L_{f,2}$-smooth w.r.t. $(\theta,x)$. For all $x\in\mathbb{R}^{d}$, $|f(\theta,x)|\leq C_{f}\|x\|^{\mathsf{n}_{f}}+D_{f}$ where $C_{f},D_{f}\geq 0$ are two constants and $\mathsf{n}_{f}\geq 1$ is an integer constant. $\bullet$ $g$ is a $\mathcal{C}^{2}$ function and $L_{g,2}$-smooth w.r.t. $\theta$ for every $x\in\mathbb{R}^{d}$. For all $x\in\mathbb{R}^{d}$, $\|\partial_{\theta}g(\theta,x)\|\leq C_{g}\|x\|^{\mathsf{n}_{g}}+D_{g}$ where $C_{g},D_{g}\geq 0$ are two constants and $\mathsf{n}_{g}\geq 1$ is an integer constant. $\bullet$ The Hessian $\partial_{x}^{2}g(\theta,x)$ is continuous w.r.t. $(\theta,x)$. $\bullet$ Beyond a compact set, $g(\theta,\cdot)$ satisfies the quadratic growth, i.e., $\exists\mathsf{D}>0$ such that $\forall|x|\geq\mathsf{D}$, $\mu_{\mathsf{qg}}\cdot\mathrm{dist}^{2}(x,\mathcal{S}(\theta))/2\leq g(\theta,x)-\min_{z\in\mathbb{R}^{d}}g(\theta,z)$ [^6] W.L.O.G., we assume that $\mathsf{D}$ is sufficiently large so that $\mathcal{S}(\theta)\subseteq\mathbb{B}_{d}(0;\mathsf{D})$ for all $\theta\in\Theta$. This assumption implies the coercivity of function $g(\theta,\cdot)$..
The regularity conditions in Section˜2, such as Lipschitzness and smoothness, are standard in the literature [50, 14]. However, unlike in [50], we do not assume a bounded $x$-domain in the lower-level problem. Instead, we impose quadratic growth beyond a compact set, ensuring $\mu_{g}^{\lambda}$ is well-defined [38]. Additionally, whereas [50] assumes boundedness of $f$, we relax this to a polynomial bound condition. We also assume a bounded $\theta$-domain and polynomial bound on $\partial_{\theta}g$, which we plan to further relax in future work.
Example 2.9.
To justify that the assumptions allow for a nontrivial function class, we provide a simple example. Since $f$ only requires regularity conditions, it can be easily constructed, so we focus on $g$. Consider $g(\theta,x)=\frac{1}{4}(\|x\|^{2}-\theta)^{2}-\frac{1}{2}\|x\|^{2}$ for $0\leq\theta\leq 1$, which satisfies all given assumptions. Here, the lower-level minima set $\mathcal{S}(\theta)=\{x:\|x\|=\theta+1\}$ is non-singleton and nonconvex.
Under Assumption 2, the mappings $(\theta,x)\mapsto f(\theta,x)$ and $(\theta,x)\mapsto\partial_{\theta}g(\theta,x)$ are continuous on the compact set $\Theta\times\mathbb{B}_{d}(0;R)$ with $R>0$. We define the constants
$L_{f,0}(R)\;:=\;\sup_{\theta\in\Theta}\ \sup_{x\in\mathbb{B}_{d}(0;R)}|f(\theta,x)|,\quad L_{g,1}(R)\;:=\;\sup_{\theta\in\Theta}\ \sup_{x\in\mathbb{B}_{d}(0;R)}\|\partial_{\theta}g(\theta,x)\|.$ (5)
By compactness and continuity, these suprema are finite.
2.1 Implications of the Condition
We review implications of Section˜2, in particular how it helps to quantify the convergence of the Gibbs measures $\mu_{\theta}^{\lambda}$ to the limiting measure $\mu_{\theta}^{0}$ under the Wasserstein-1 metric. In this subsection, $\theta$ is considered fixed, and hence for ease of the notation, we omit it and simply write $\mu^{\lambda}$, and $\mu^{0}$.
2.1.1 Implications of the Embedding Submanifold Structure in Section˜2
Following Remark 1 in [35], WLOG and for the ease of notation, we assume that the atlas of the manifold $\mathcal{S}$ consists of a single chart $(\mathcal{U},\phi)$, where $\mathcal{U}=\mathcal{S}$. Let $\Gamma=\phi(\mathcal{U})\subseteq\mathbb{R}^{\texttt{k}}$ denote the coordinate domain [^7] The standard technique of partition of unity can be used to translate the results to the multi-chart case.. The corresponding parametrization (the inverse of the chart) can be represented as $x=\mathsf{S}(u):=[m_{1}(u),\ldots,m_{d}(u)]^{\top}$ for all $u\in\Gamma$, where $m_{j}:\Gamma\to\mathbb{R}$ for $j\in[d]$ are $\mathcal{C}^{2}$ coordinate functions.
Tubular Neighborhood.
For $x\in\mathcal{S}$, the tangent space at $x$ is denoted by $T_{x}(\mathcal{S})$. The orthogonal complement of $T_{x}(\mathcal{S})$ in $\mathbb{R}^{d}$ is denoted by $T_{x}(\mathcal{S})^{\perp}$. The tubular neighborhood $\mathcal{N}(\varepsilon)$ of $\mathcal{S}$ for $\varepsilon>0$ is defined as: $\mathcal{N}(\varepsilon):=\{x+v:x\in\mathcal{S},v\in T_{x}(\mathcal{S})^{\perp},\|v\|<\varepsilon\}.$ A key property of the $\mathcal{C}^{2}$ embedded submanifold is as follows [^8] For simplicity, we assume a single chart and omit the dependence on the inclusion $i_{\mathcal{S}}$. [52, Theorem 6.24]: There exists $\varepsilon_{0}>0$ such that for all $0<\varepsilon\leq\varepsilon_{0}$, there is a $\mathcal{C}^{2}$ diffeomorphism $\Phi$:
$\Phi:\mathcal{S}\times\mathbb{B}_{d-\texttt{k}}(0;\varepsilon)\to\mathcal{N}(\varepsilon),\quad(x,t)\mapsto x+\sum_{j=1}^{d-\texttt{k}}t_{j}e^{(j)}(x),$ (6)
where $\{e^{(1)}(x),\dots,e^{(d-\texttt{k})}(x)\}$ forms an orthonormal basis for the normal space $T_{x}(\mathcal{S})^{\perp}$. Under the local chart $(\Gamma,\phi)$, any $y\in\mathcal{N}(\varepsilon)$ can be written as $y=x+v$, where $x$ is a point on $\mathcal{S}$ and $v\perp\mathcal{S}$ at $x$ with $\|v\|\leq\varepsilon$. The map $y(u,t):=\Phi(\mathsf{S}(u),t)\rightarrow(x,v)$ is a $\mathcal{C}^{2}$ diffeomorphism.
Volume measure.
We define the Riemannian (tensor) metric $\mathsf{g}_{\mathcal{S}}$ on $\mathcal{S}$ as the pullback metric by including map $i_{\mathcal{S}}:\mathcal{S}\hookrightarrow\operatorname{\mathbb{R}}^{d}$, i.e. $\mathsf{g}_{\mathcal{S}}=i^{\ast}_{\mathcal{S}}(\mathsf{g}_{E})$, where $\mathsf{g}_{E}$ is the standard Riemannian metric on $\operatorname{\mathbb{R}}^{d}$ and $i_{\mathcal{S}}^{\ast}$ is the pullback map associated with $i_{\mathcal{S}}$. Now we can define the standard volume form $d\mathcal{M}$ [^9] Note that $\mathrm{d}\mathcal{M}$ stands for $\mathrm{d}\mathcal{M}_{\theta}$ in all the following sections. as $\mathrm{d}\mathcal{M}(u)=\sqrt{\det(\mathsf{g}_{\mathcal{S}})}\mathrm{d}u,\ u\in\Gamma,$ where $\text{det}(\mathsf{g}_{\mathcal{S}})$ is the determinant of $\mathsf{g}_{\mathcal{S}}$.
Poincare constant.
A probability measure $\pi$ over $\mathbb{R}^{d}$ is said to satisfy the Poincaré inequality with inverse Poincaré constant $C_{PI}$ if for all test functions $\varphi\in\mathbb{H}^{1}(\mu)$, the following inequality holds:
$\int\left(\varphi-\int\varphi\mathrm{d}\mu\right)^{2}\mathrm{d}\mu\leq C_{PI}\int\|\nabla\varphi\|^{2}\mathrm{d}\mu,$
where $\mathbb{H}^{1}(\mu)$ denotes the Sobolev space weighted by $\mu$. Importantly, for all functions $g:\mathbb{R}^{d}\rightarrow\mathbb{R}$, the constant $C_{PI}$ for the Gibbs probability measure $\mu^{\lambda}\propto\exp(-g/\lambda)$ is independent of $\lambda$ for all sufficiently small $\lambda\leq\lambda_{0}$, where both $\lambda_{0}$ and an upper bound of $C_{PI}$ are explicitly given in [35].
2.1.2 Limiting Measure of Gibbs Measures
To sample approximately from the optimal set $\mathcal{S}$, we select a temperature parameter $\lambda$ approaching zero, which causes the Gibbs measure (3) to concentrate on $\mathcal{S}$. Let $\mu^{0}:=\lim_{\lambda\to 0}\mu^{\lambda}$ denote the weak limit of the Gibbs measures. By combining the manifold characterization of the optimal set for functions in [35] with the asymptotic analysis in [41, Theorem 3.1], the following lemma establishes that $\mu^{0}$ is well-defined, has support restricted to $\mathcal{S}$, and admits a closed-form density function with respect to the Riemannian volume measure $d\mathcal{M}$ on the manifold.
Proposition 2.10 (Theorem 3.1 in [41]).
Suppose that Section˜2 holds and let k be the dimension of the submanifold $\mathcal{S}$. Let $\mu^{0}:=\lim_{\lambda\rightarrow 0}\mu^{\lambda}$ be the weak limit of the Gibbs measures, then $\mathrm{supp}(\mu^{0})=\mathcal{S}$ and
$\rho(u):=\frac{\mathrm{d}\mu^{0}}{\mathrm{d}\mathcal{M}}=\frac{\left|\partial^{2}_{t}g(y(u,t))_{|t=0}\right|^{-\frac{1}{2}}}{\int_{\mathcal{S}}\left|\partial_{t}^{2}g(y(v,t))_{|t=0}\right|^{-\frac{1}{2}}\mathrm{d}\mathcal{M}(v)}.$ (7)
Here $|\cdot|:\operatorname{\mathbb{R}}^{(d-\texttt{k})\times(d-\texttt{k})}\rightarrow\operatorname{\mathbb{R}}$ denotes the matrix determinant. We write $\mu^{0}(\mathrm{d}u)=\rho(u)\mathrm{d}\mathcal{M}(u)$ [^10] Note that $\mu^{0}$ and $\rho$ stand for $\mu^{0}_{\theta}$ and $\rho_{\theta}$ in all the following sections..
Wasserstein-1 distance between the Gibbs measure and its limit measure.
The following proposition from [38, Theorem 3.6] proves the non-asymptotic convergence of $\mu^{\lambda}$ as $\lambda\to 0$ to $\mu^{0}$ (defined above) in the Wasserstein-1 sense [^11] We use the assumption on $g$ in Section 2 to ensure that the conditions in [38, Theorem 3.6] hold. Further discussion on the validity of their conditions can be found in Appendix A.3.. Establishing convergence in $\mathbb{W}_{1}$ is crucial for our analysis because it directly bounds the approximation error between the superquantile-Gibbs surrogate and the superquantile loss taken over the limiting measure $\mu^{0}$ (to be defined in (11)). This bound is established by exploiting the Lipschitz continuity of the upper-level objective (see Section˜4.2).
Proposition 2.11 (Theorem 3.6 in [38] ).
Under Sections˜2, 2 and 2, $\mathbb{W}_{1}(\mu^{\lambda},\mu^{0})\leq K_{0}\,\lambda^{\frac{1}{2}}$ where $K_{0}$ is $\lambda$-independent.
3 Regularity of the Solution Mapping and Geometric Properties
We first establish that $\mathcal{S}(\theta)$ varies locally Lipschitz in Hausdorff distance under the local condition, and consequently that the hyper-objective $F_{\max}$ is Lipschitz continuous (Corollary 3.1). In Section 3.2, we show that the minimizer manifolds $\{\mathcal{S}(\theta)\}_{\theta\in\Theta}$ share a common intrinsic dimension k (Lemma 3.2), and then show $\theta$–independent geometric bounds such as a uniform diameter, curvature control, and a positive injectivity radius (Proposition 3.2). We also show that the limiting Gibbs density on $\mathcal{S}(\theta)$ is uniformly bounded from below and above (Section˜3.2). These uniform geometric properties will play a role later in our approximation analysis (Section˜4).
3.1 Lipschitz Continuity of the Solution Mapping
In this subsection, our goal (Lemma 3.1) is to show that, under the condition, the set-valued mapping $\theta\mapsto\mathcal{S}(\theta)$ is locally Lipschitz with respect to the Hausdorff distance. We then combine this regularity with the Lipschitz continuity of $f(\theta,x)$ to deduce the Lipschitz continuity of $F_{\max}$ (Corollary 3.1). We begin by recalling the notion of Hausdorff distance.
Definition 3.1 (Hausdorff distance).
Let $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ be two subsets of $\mathbb{R}^{d}$. The Hausdorff distance between $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ is defined as
$\text{dist}(\mathcal{S}_{1},\mathcal{S}_{2}):=\max\left\{\sup_{x_{1}\in\mathcal{S}_{1}}\inf_{x_{2}\in\mathcal{S}_{2}}\|x_{1}-x_{2}\|,\;\sup_{x_{2}\in\mathcal{S}_{2}}\inf_{x_{1}\in\mathcal{S}_{1}}\|x_{1}-x_{2}\|\right\}.$
Since the condition provides only a local PŁ inequality in a neighborhood of the minimizer set $\mathcal{S}(\theta)$, Lipschitz stability of the solution mapping does not follow from standard global error-bound arguments. By contrast, under a global PŁ condition for $g(\theta,\cdot)$, one obtains a global error bound and then deduces Lipschitz continuity of $\theta\mapsto\mathcal{S}(\theta)$ directly from the smoothness of $g$ on $\theta$ (see, e.g., [50, 14]). In our local setting, the obstruction is geometric: a local error bound alone still allows the minimizer set to have multiple disconnected components, and minimizers could jump from one component to another under arbitrarily small changes in $\theta$, precluding any Hausdorff Lipschitz control. This is why we need that $\mathcal{S}(\theta)$ is connected under our assumptions (see Section˜2) to prove a quantitative stability result showing that minimizers cannot jump and must move continuously under small perturbations of $\theta$. The next lemma makes this precise.
Lemma 3.2.
Under Assumption 2 and Assumption 2, the minimizer mapping $\theta\mapsto\mathcal{S}(\theta)$ is locally Lipschitz continuous in the Hausdorff distance sense: there exists $\gamma>0$ such that for every $\theta_{1}$ and $\theta_{2}$ with $\|\theta_{1}-\theta_{2}\|\leq\gamma$, we have
$\operatorname{dist}\bigl(\mathcal{S}(\theta_{1}),\mathcal{S}(\theta_{2})\bigr)\leq\frac{L_{g,2}}{\mu}\|\theta_{1}-\theta_{2}\|.$
Proof.
The proof has two steps. First, the local PŁ condition and $\mathcal{C}^{2}$ regularity imply that there exists a neighborhood $\mathcal{U}(\theta_{1})$ of the minimizer manifold $\mathcal{S}(\theta_{1})$ such that inside $\mathcal{U}(\theta_{1})$, the distance to $\mathcal{S}(\theta_{1})$ is controlled by $\|\partial_{x}g(\theta_{1},\cdot)\|$. Second, we show that for small perturbations $\theta_{2}$ of $\theta_{1}$, every minimizer $x_{2}\in\mathcal{S}(\theta_{2})$ lies in $\mathcal{U}(\theta_{1})$, by comparing $g(\theta_{1},x_{2})$ and $g(\theta_{1},x_{1})$ for $x_{1}\in\mathcal{S}(\theta_{1})$ using Lipschitz continuity in $\theta$. Once $x_{2}\in\mathcal{U}(\theta_{1})$, the error bound and $L_{g,2}$–smoothness in $\theta$ give
$\operatorname{dist}\bigl(x_{2},\mathcal{S}(\theta_{1})\bigr)\;\lesssim\;\|\partial_{x}g(\theta_{1},x_{2})\|\;\lesssim\;L_{g,2}\|\theta_{1}-\theta_{2}\|,$
and taking the supremum over $x_{2}\in\mathcal{S}(\theta_{2})$ yields the desired local Lipschitz bound in Hausdorff distance.
Step 1: local error bound around $\mathcal{S}(\theta)$. Fix $\theta\in\Theta$. By Section˜2, Section˜2 implies that every local minimizer of $g(\theta,\cdot)$ is in fact global, so the local PŁ inequality in Section˜2 holds in a neighborhood of the global minimizer set $\mathcal{S}(\theta)$. In particular, there exists an open neighborhood $\mathcal{N}(\mathcal{S}(\theta))$ of $\mathcal{S}(\theta)$ such that
$g(\theta,x)-\min_{z\in\mathbb{R}^{d}}g(\theta,z)\;\leq\;\frac{1}{2\mu}\,\bigl\|\partial_{x}g(\theta,x)\bigr\|^{2}\qquad\forall x\in\mathcal{N}(\mathcal{S}(\theta)).$ (8)
By the local PŁ condition and $\mathcal{C}^{2}$ regularity, $g(\theta,\cdot)$ satisfies a local error bound in a smaller neighborhood of $\mathcal{S}(\theta)$ [73, Propositions 2.2 & 2.8]: there exists an open set $\mathcal{U}(\theta)\subseteq\mathcal{N}(\mathcal{S}(\theta))$ with $\mathcal{S}(\theta)\subset\mathcal{U}(\theta)$ such that
$\operatorname{dist}\bigl(x,\mathcal{S}(\theta)\bigr)\;\leq\;\mu^{-1}\,\bigl\|\partial_{x}g(\theta,x)\bigr\|\qquad\forall x\in\mathcal{U}(\theta).$ (9)
Moreover, since $\mathcal{S}(\theta)$ is compact and $\mathcal{U}(\theta)$ is open, we can find $\tilde{\epsilon}>0$ such that the $\tilde{\epsilon}-$sub-level set of $g(\theta,\cdot)$ lies inside $\mathcal{U}(\theta)$:
Lemma 3.3.
Under the local error bound for $g(\theta,\cdot)$ with $\mathcal{U}(\theta)$ being as above, there exists $\tilde{\epsilon}>0$ such that
$\Bigl\{x\in\mathbb{R}^{d}:\;g(\theta,x)-\min_{z\in\mathbb{R}^{d}}g(\theta,z)\leq\tilde{\epsilon}\Bigr\}\;\subset\;\mathcal{U}(\theta).$
The proof of Lemma 3.1 is given in Appendix B.1. Intuitively, since $\mathcal{U}(\theta)$ is an open neighborhood of $\mathcal{S}(\theta)$ and $g(\theta,\cdot)$ is continuous, a small enough sub-level set around $\mathcal{S}(\theta)$ must be contained in $\mathcal{U}(\theta)$.
Step 2: stability of $\mathcal{S}(\theta)$ under small perturbations of $\theta$. Let $\mathcal{U}(\theta_{1})$ and $\tilde{\epsilon}>0$ be as above. By Assumption 2 and using Equation˜5, $g$ is $L_{g,1}(\mathsf{D})$–Lipschitz in $\theta$ on a compact domain $\mathbb{B}_{d}(0;\mathsf{D})$ in $x$, so for any $x\in\mathbb{R}^{d}$ and any $\theta_{1},\theta_{2}\in\Theta$,
$\bigl|g(\theta_{1},x)-g(\theta_{2},x)\bigr|\;\leq\;L_{g,1}(\mathsf{D})\;\|\theta_{1}-\theta_{2}\|.$
Let $x_{1}\in\mathcal{S}(\theta_{1})$ and $x_{2}\in\mathcal{S}(\theta_{2})$, where $\theta_{2}$ will be taken close to $\theta_{1}$. Using optimality of $x_{1}$ for $g(\theta_{1},\cdot)$ and of $x_{2}$ for $g(\theta_{2},\cdot)$, we have
$0$ $\leq g(\theta_{1},x_{2})-g(\theta_{1},x_{1})$
$=\bigl[g(\theta_{1},x_{2})-g(\theta_{2},x_{2})\bigr]+\bigl[g(\theta_{2},x_{2})-g(\theta_{2},x_{1})\bigr]+\bigl[g(\theta_{2},x_{1})-g(\theta_{1},x_{1})\bigr]$
$\leq\bigl|g(\theta_{1},x_{2})-g(\theta_{2},x_{2})\bigr|+0+\bigl|g(\theta_{2},x_{1})-g(\theta_{1},x_{1})\bigr|$
$\leq 2L_{g,1}(\mathsf{D})\,\|\theta_{1}-\theta_{2}\|.$ (10)
Hence, if we choose $\gamma\;\leq\;\tilde{\epsilon}/(2L_{g,1}(2\mathsf{D}))$ and $\|\theta_{1}-\theta_{2}\|\leq\gamma$, then (10) implies
$g(\theta_{1},x_{2})-\min_{z}g(\theta_{1},z)\;=\;g(\theta_{1},x_{2})-g(\theta_{1},x_{1})\;\leq\;\tilde{\epsilon}$,
so $x_{2}$ belongs to the sub-level set in Lemma 3.1. Therefore $x_{2}\in\mathcal{U}(\theta_{1})$ whenever $\|\theta_{1}-\theta_{2}\|\leq\gamma$. Then the error bound (9) for $g(\theta_{1},\cdot)$ yields
$\operatorname{dist}\bigl(x_{2},\mathcal{S}(\theta_{1})\bigr)\;\leq\;\mu^{-1}\,\bigl\|\partial_{x}g(\theta_{1},x_{2})\bigr\|.$
Because $x_{2}$ is a minimizer for $g(\theta_{2},\cdot)$, we have $\partial_{x}g(\theta_{2},x_{2})=0$, hence
$\bigl\|\partial_{x}g(\theta_{1},x_{2})\bigr\|$ $=\bigl\|\partial_{x}g(\theta_{1},x_{2})-\partial_{x}g(\theta_{2},x_{2})\bigr\|\leq L_{g,2}\,\|\theta_{1}-\theta_{2}\|,$
where we used the $L_{g,2}$–smoothness of $g$ with respect to $\theta$. Combining the two last inequalities results in
$\operatorname{dist}\bigl(x_{2},\mathcal{S}(\theta_{1})\bigr)\;\leq\;\frac{L_{g,2}}{\mu}\,\|\theta_{1}-\theta_{2}\|.$
By symmetry (interchanging the roles of $\theta_{1}$ and $\theta_{2}$ and repeating the same argument) and recalling the definition of the Hausdorff distance in Section˜3.1, we get
$\operatorname{dist}\bigl(\mathcal{S}(\theta_{1}),\mathcal{S}(\theta_{2})\bigr)\;\leq\;\frac{L_{g,2}}{\mu}\,\|\theta_{1}-\theta_{2}\|\qquad\text{whenever}\quad\|\theta_{1}-\theta_{2}\|\leq\gamma.$
Having established the local Lipschitz continuity of $\mathcal{S}(\theta)$ in Lemma 3.1, we now turn to the regularity of the hyper-objective $F_{\max}$. The following corollary follows by combining the local Lipschitz continuity of $\theta\mapsto\mathcal{S}(\theta)$ with the Lipschitz continuity of $f(\theta,x)$ [^12] A standard chaining argument shows that if $|F_{\max}(\theta_{1})-F_{\max}(\theta_{2})|\leq L\|\theta_{1}-\theta_{2}\|$ whenever $\|\theta_{1}-\theta_{2}\|\leq\gamma$, then the same constant works uniformly on $\Theta$.; see, e.g., [14, Theorem 3.1].
Corollary 3.4.
Under Sections˜2 and 2, $F_{\max}(\theta)$ is $(L_{f,1}+L_{f,1}{L_{g,2}}/{\mu})$-Lipschitz continuous.
For completeness, we provide the proof in Appendix˜B.
3.2 Common Dimension and Geometric Properties of $\{\mathcal{S}(\theta)\}_{\theta\in\Theta}$
For each $\theta\in\Theta$, the local PŁ condition implies that on $\mathcal{S}(\theta)$ the Hessian $\partial_{x}^{2}g(\theta,x)$ is positive definite in the normal directions and degenerate in the tangent directions [73, Proposition 2.7]. In particular, its rank $\texttt{r}(\theta)$ is constant over $x\in\mathcal{S}(\theta)$, and the dimension of the minimizer manifold can be written as $\dim\mathcal{S}(\theta)=\texttt{k}(\theta)=d-\texttt{r}(\theta)$. Moreover, the solution mapping $\theta\mapsto\mathcal{S}(\theta)$ is continuous (Section˜3.1), and $\partial_{x}^{2}g(\theta,x)$ is continuous in $(\theta,x)$. Thus, when $\theta$ is perturbed slightly, points in $\mathcal{S}(\theta)$ move continuously and the corresponding Hessians remain close, so no eigenvalue can cross zero and the rank $\texttt{r}(\theta)$ cannot change. Hence $\texttt{r}(\theta)$, and therefore $\texttt{k}(\theta)$, is locally constant in $\theta$ and we obtain a single intrinsic dimension shared by all manifolds $\{\mathcal{S}(\theta)\}_{\theta\in\Theta}$. We summarize this in the following lemma.
Lemma 3.5.
Under Sections˜2, 2 and 2, the manifolds $\{\mathcal{S}(\theta)\}_{\theta\in\Theta}$ admit a common dimension k.
The proof is given in Section˜B.3. Lemma 3.2 allows us to work with a single geometric parameter k in all later bounds, rather than letting the dimension depend on $\theta$. In particular, when we study the Gibbs relaxation and the superquantile approximation of $F_{\max}$, the scaling with accuracy is expressed in terms of this common dimension k (see ˜4.5). Throughout the paper, we consider $\texttt{k}\geq 1$, since when $\texttt{k}=0$, $\mathcal{S}(\theta)$ is a singleton and minima-selection is redundant. Next, we collect basic geometric properties of $\{\mathcal{S}(\theta)\}_{\theta\in\Theta}$ that will be used in our approximation analysis (see Section˜4). The proofs are based on the uniform compactness of the manifolds.
Proposition 3.6.
Under compactness of $\Theta$ and Sections˜2 and 2, the following holds for all $\theta\in\Theta$.
  • $\bullet$
    There exists a constant $\mathsf{D}$ independent of $\theta$, such that $\text{diam}(\mathcal{S}(\theta))\leq\mathsf{D}$.
  • $\bullet$
    The absolute value of sectional curvature [^13] Refer to Section A.1 for the definition of sectional curvature. of $\mathcal{S}(\theta)$ is bounded by some constant $\mathsf{C}_{0}>0$.
  • $\bullet$
    There is a constant $r_{0}>0$ such that the injectivity radius [^14] Refer to Section A.1 for the definition of injectivity radius. $\mathrm{inj}(p)\geq r_{0}$ for all $p\in\mathcal{S}(\theta)$.
The proof is given in Section˜B.4. Finally, we show a uniform bound on the limiting density $\rho_{\theta}$ defined in eq.˜7.
Proposition 3.7.
Under Sections˜2 and 2, there exists a constant $\kappa>0$ such that $\kappa\leq\rho_{\theta}(x)\leq 1/\kappa$ for all $x\in\mathcal{S}(\theta)$ and $\theta\in\Theta$.
The proof is given in Section˜B.5. This result ensures that integration with respect to the limiting Gibbs measure is uniformly comparable to integration with respect to the volume measure on $\mathcal{S}(\theta)$, which is crucial for our approximation analysis in the next section.
Remark 3.8.
(On the size of the constant $\kappa$) We chose $\kappa\;=\;\mathsf{b}\cdot({\mu}/{\mathsf{M}_{g}})^{\frac{d-\texttt{k}}{2}}$ in the proof of Section˜3.2, where $\mu>0$ is a uniform lower bound on the nonzero normal eigenvalues of $\partial_{x}^{2}g(\theta,x)$, $\mathsf{M}_{g}>0$ is a uniform upper bound on $\|\partial_{x}^{2}g(\theta,x)\|$ over $\mathcal{S}(\theta)$, and the constant $\mathsf{b}>0$ comes from the volume bounds on $\mathcal{S}(\theta)$ in Section˜B.5 in Section˜B.5. In the worst case, the ratio $\mu/\mathsf{M}_{g}$ can be very small, so $\kappa$ may be exponentially small. In contrast, in an idealized case where in normal coordinates $g(\theta,\cdot)$ behaves like a pure squared distance to $\mathcal{S}(\theta)$, we have $\mu=\mathsf{M}_{g}$, so $\kappa=\mathsf{b}$.
4 Approximation Guarantees with SQ-G Relaxation
In this section, we show that the superquantile-Gibbs relaxation $\tilde{F}_{{\texttt{SQ-G}}}$, defined in Equation˜2, provides a valid approximation to $F_{\max}$ under Sections˜2, 2 and 2, given suitable choices of the parameters $\delta$ and $\lambda$. Our approximation result justifies using $\tilde{F}_{{\texttt{SQ-G}}}$ as a surrogate for $F_{\max}$ when applying zeroth-order methods to optimize the latter. We establish the result through the following steps:
  1. 1.
    The superquantile loss (SQ-loss), defined as
    $\tilde{F}_{\texttt{SQ}}(\theta):=\min_{\beta}\left\{\beta+\frac{1}{\delta}\mathbb{E}_{X\sim\mu^{0}_{\theta}}[\max\{f(\theta,X)-\beta,0\}]\right\},$ (11)
    closely approximates the hyper-objective $F_{\max}(\theta)$ point-wisely.
  2. 2.
    The difference between $\tilde{F}_{\texttt{SQ}}$ and $\tilde{F}_{{\texttt{SQ-G}}}$ is bounded by the Wasserstein-1 distance between $\mu^{\lambda}_{\theta}$ and $\mu^{0}_{\theta}$: $|\tilde{F}_{\texttt{SQ}}(\theta)-\tilde{F}_{{\texttt{SQ-G}}}(\theta)|\leq\mathcal{O}(\delta^{-1}\mathbb{W}_{1}(\mu^{\lambda}_{\theta},\mu^{0}_{\theta})).$
Below we formally present these results.
4.1 Approximation of $F_{\max}$ with $\tilde{F}_{\texttt{SQ}}$
Recall that the $(1-\delta)$-superquantile of the random variable $Y=f(\theta,X)$ with $X\sim\mu_{\theta}^{0}$ is the average of the values of $Y$ that lie above its $(1-\delta)$-quantile threshold. Our goal is to show that, for small $\delta$, the superquantile $\tilde{F}_{\texttt{SQ}}(\theta)$ provides a good approximation of $F_{\max}(\theta)$. To build intuition, fix $\theta$ and suppose that the function $x\mapsto f(\theta,x)$ attains its maximum over $\mathcal{S}(\theta)$ at a unique point $x^{*}(\theta)\in\mathcal{S}(\theta)$. Consider the high-value set $\mathfrak{X}_{\delta}(\theta):=\{x\in\mathcal{S}(\theta):f(\theta,x)\text{ is above the $(1-\delta)$-quantile threshold}\}$. This set forms a neighborhood of $x^{*}(\theta)$. Since $x^{*}(\theta)\in\mathfrak{X}_{\delta}(\theta)$ and $f$ is Lipschitz, the approximation error $|\tilde{F}_{\texttt{SQ}}(\theta)-F_{\max}(\theta)|$ is controlled by how far one can move away from $x^{*}(\theta)$ while staying in $\mathfrak{X}_{\delta}(\theta)$. Denoting by $r$ the minimal geodesic distance from $x^{*}(\theta)$ to the boundary of $\mathfrak{X}_{\delta}(\theta)$, the Lipschitz continuity of $f$ then bounds the error in terms of $r$.
Thus, to bound the approximation error, it suffices to bound $r$. To bound this radius $r$ given our control over the probability mass $\delta$, we proceed in two steps: 1) Probability to Volume: We first translate probability mass $\delta$ into Riemannian volume. This is possible because the density of the limiting measure $\mu_{\theta}^{0}$ is strictly lower-bounded (see Section˜3.2), ensuring that a small probability mass implies a small geometric volume. 2) Volume to Radius: We then translate Riemannian volume into the radius $r$. By establishing that the Riemannian volume scales similarly to Euclidean volume ($\propto r^{\texttt{k}}$), we deduce that the radius shrinks at the rate $r\lesssim\delta^{1/\texttt{k}}$.
Classical results such as [36] give a pointwise asymptotic expansion of the volume of a small geodesic ball, showing that $\mathcal{M}(B_{\mathsf{g}}(p;r))/r^{\texttt{k}}\to c(p)\,$ as $r\to 0$ with a coefficient depending on the local curvature at $p$. For our purposes, this is not sufficient: we will see that in the proof of ˜4.2 we need a uniform, non-asymptotic lower bound that holds simultaneously for all centers $p\in\mathcal{S}(\theta)$ and all radii $0<r\leq r_{1}$, with $r_{1}$ depending only on the global bounds on sectional curvature and injectivity radius from Section˜3.2. Section˜4.1 provides exactly this quantitative version by turning the asymptotic comparison into an explicit estimate with controlled radius and constants.
Lemma 4.1.
Under Sections˜2 and 2, there exist constants $\mathsf{c}_{H},\mathsf{c}_{L}>0$, and $c_{1}>0$ such that for every $\theta\in\Theta$, for any $p\in\mathcal{S}(\theta)$, and any radius $r$ satisfying $0<r<\min\{r_{0},c_{1}/\sqrt{2\mathsf{C}_{0}}\}$, the Riemannian volume of the geodesic ball $\mathbb{B}_{\mathsf{g}}(p;r)$ satisfies
$\mathsf{c}_{L}\cdot{\texttt{Leb}}_{\texttt{k}}\bigl(\mathbb{B}_{\texttt{k}}(p;r)\bigr)\leq\mathcal{M}\bigl(\mathbb{B}_{\mathsf{g}}(p;r)\bigr)\leq\mathsf{c}_{H}\cdot{\texttt{Leb}}_{\texttt{k}}\bigl(\mathbb{B}_{\texttt{k}}(p;r)\bigr),$
where $r_{0}$ and $\mathsf{C}_{0}$ are defined in Section˜3.2, ${\texttt{Leb}}_{\texttt{k}}\bigl(\mathbb{B}_{\texttt{k}}(p;r)\bigr)$ is the standard k-dimensional Lebesgue volume of a ball of radius $r$ in $\mathbb{R}^{\texttt{k}}$, and $\mathcal{M}$ is the Riemannian volume measure on $(\mathcal{S}(\theta),\mathsf{g})$.
The proof of Section˜4.1 is given in Section˜C.1. Based on Section˜4.1, the Riemannian volume of a geodesic ball with radius $r$, $\mathcal{M}\bigl(\mathbb{B}_{\mathsf{g}}(p;r)\bigr)$, scales as $r^{\texttt{k}}$. At a high level, the lemma relies on two geometric ingredients: (i) The injectivity radius bound $r_{0}>0$ guarantees that, around every $p\in\mathcal{S}$, we can introduce normal coordinates on the whole geodesic ball $\mathbb{B}_{\mathsf{g}}(p;r)$ for all $r\leq r_{0}$. (ii) The curvature bound $|\mathrm{Sec}|\leq\mathsf{C}_{0}$ then implies that, in these normal coordinates and for radii $r$ of order $1/\sqrt{\mathsf{C}_{0}}$, at a given point $v\in\mathbb{B}_{\mathsf{g}}(p;r)$, the matrix $(\mathsf{g}_{ij}(v))_{i,j=1}^{\texttt{k}}$ representing the Riemannian metric remains uniformly close to the Euclidean metric $(\delta_{ij}(v))_{i,j=1}^{\texttt{k}}$ [^15] Here $\delta_{ij}$ denotes the Kronecker delta: $\delta_{ij}=1$ if $i=j$ and $\delta_{ij}=0$ otherwise.. Consequently, the Riemannian volume element $\sqrt{\det(\mathsf{g}_{\mathcal{S}}(v))}\,dv$ is uniformly comparable from below and above to the Euclidean volume element $dv$ on such balls. Integrating this comparison over $\mathbb{B}_{\texttt{k}}(0;r)$ yields the stated bounds on the Riemannian volume $\mathcal{M}\bigl(\mathbb{B}_{\mathsf{g}}(p;r)\bigr)$. We now state the following key result.
Theorem 4.2 ($\tilde{F}_{\texttt{SQ}}$ approximates $F_{\max}$).
Recall $\tilde{F}_{\texttt{SQ}}(\theta):=\min_{\beta}\phi_{0,\delta}(\theta,\beta)$. Under Sections˜2, 2 and 2, for every $\theta\in\Theta$ and for
$0<\delta\leq\mathsf{c}_{L}\cdot C_{\texttt{k}}\cdot\kappa\cdot\left(4^{-1}\min\{c_{1}/\sqrt{2\mathsf{C}_{0}},r_{0}\}\right)^{\texttt{k}},$
we have
$|F_{\max}(\theta)-\tilde{F}_{\texttt{SQ}}(\theta)|\leq\frac{4L_{f,1}\delta^{\frac{1}{\texttt{k}}}}{(\kappa\cdot\mathsf{c}_{L}\cdot C_{\texttt{k}})^{\frac{1}{\texttt{k}}}},$
where $\mathsf{c}_{L}$ and $c_{1}$ are defined in Section˜4.1, $\kappa$ in Section˜3.2, k is the dimension of $\mathcal{S}(\theta)$, $C_{\texttt{k}}=\pi^{\texttt{k}/2}/\Gamma(\texttt{k}/2+1)$, and $r_{0},\mathsf{C}_{0}$ are from Section˜3.2.
Proof.
Let $Z:=f(\theta,X)$ where $X\sim\mu_{g}^{0}(\theta)$. Since $Z$ takes values in $\{f(\theta,x):x\in\mathcal{S}(\theta)\}$, we have the elementary bounds
$\max_{x\in\mathcal{S}(\theta)}f(\theta,x)\;\geq\;\tilde{F}_{\texttt{SQ}}(\theta)\;\geq\;\text{Q}_{1-\delta}(Z),$
hence
$|F_{\max}(\theta)-\tilde{F}_{\texttt{SQ}}(\theta)|$ $=F_{\max}(\theta)-\tilde{F}_{\texttt{SQ}}(\theta)$
$\leq F_{\max}(\theta)-\text{Q}_{1-\delta}\bigl(f(\theta,X)\bigr).$ (12)
Thus it remains to bound $F_{\max}(\theta)-\text{Q}_{1-\delta}\bigl(f(\theta,X)\bigr)$.
We fix $\theta$ and drop it from the notation. Write $\mathcal{S}=\mathcal{S}(\theta)$, $f=f(\theta,\cdot)$, $f^{*}=\max_{x\in\mathcal{S}}f(x)$, $\mathcal{S}_{f}=\arg\max f$ (nonempty by compactness), and draw $Y\sim\mu^{0}$ as in Section˜2.1.2. Denote by $\mathrm{dist}_{\mathsf{g}}$ the geodesic distance on $\mathcal{S}$, and write $\mathbb{B}_{\mathsf{g}}(x;r):=\{u\in\mathcal{S}:\mathrm{dist}_{\mathsf{g}}(u,x)<r\}$ for the geodesic ball (refer to (35) in Section˜A.1). While $\mathcal{S}_{f}$ may generally consist of multiple connected components, we assume without loss of generality that it is connected, as the extension of our analysis below to the general case is straightforward.
Proof outline.
(i) Let $\varepsilon:=f^{*}-\text{Q}_{1-\delta}(f(Y))$ and $A_{\varepsilon}:=\{u\in\mathcal{S}:\,f(u)\geq f^{*}-\varepsilon\}$. Set $r:=\mathrm{dist}_{\mathsf{g}}\bigl(\overline{\mathcal{S}\setminus A_{\varepsilon}},\mathcal{S}_{f}\bigr)$. A minimizing geodesic between these sets has a midpoint $x$ with a geodesic ball $\mathbb{B}_{\mathsf{g}}(x;r/4)\subset A_{\varepsilon}\setminus\mathcal{S}_{f}$. (ii) By Section˜3.2, $\mu^{0}(\cdot)$ is comparable to $\mathcal{M}(\cdot)$; since $\mu^{0}(A_{\varepsilon})=\delta$, we get $\mathcal{M}(\mathbb{B}_{\mathsf{g}}(x;r/4))\lesssim\delta$. Under Section˜3.2 and by Section˜4.1, $\mathcal{M}(\mathbb{B}_{\mathsf{g}}(x;r/4))\gtrsim(r/4)^{\texttt{k}}$. Hence $r\lesssim\delta^{1/\texttt{k}}$ (for small $\delta$). (iii) Lipschitzness of $f$ yields $f^{*}-\text{Q}_{1-\delta}(f(Y))\leq L_{f,1}\,r$. Combining with $r\lesssim\delta^{1/\texttt{k}}$ and the reduction in (12) gives the claimed rate.
(i) Near–optimal set contains a geodesic ball.
For any set $A\subseteq\mathcal{S}$, we write $\overline{A}$ for the closure of $A$ in $\mathcal{S}$, i.e., with respect to the topology induced by the Riemannian metric on $\mathcal{S}$. We then define the geodesic set–set distance using the closure
$r:=\mathrm{dist}_{\mathsf{g}}\bigl(\overline{\mathcal{S}\setminus A_{\varepsilon}},\,\mathcal{S}_{f}\bigr)=\min_{z\in\overline{\mathcal{S}\setminus A_{\varepsilon}}}\ \min_{y\in\mathcal{S}_{f}}\ \mathrm{dist}_{\mathsf{g}}(z,y).$
Since $\overline{\mathcal{S}\setminus A_{\varepsilon}}$ and $\mathcal{S}_{f}$ are compact, minimizers $z\in\overline{\mathcal{S}\setminus A_{\varepsilon}}$ and $y\in\mathcal{S}_{f}$ exist with $\mathrm{dist}_{\mathsf{g}}(z,y)=r$. Let $\gamma$ be a minimizing geodesic from $y$ to $z$ and set the geodesic midpoint $x:=\gamma(r/2)$. We claim
$\mathbb{B}_{\mathsf{g}}(x;r/4)\subset A_{\varepsilon}\setminus\mathcal{S}_{f}.$
Indeed, for any $w\in\mathbb{B}_{\mathsf{g}}(x;r/4)$, using that the map $x\mapsto\mathrm{dist}_{\mathsf{g}}(x,\mathcal{S}_{f})$ is $1$-Lipschitz [^16] For any $a\in\mathcal{S}_{f}$, the triangle inequality gives $\mathrm{dist}_{\mathsf{g}}(w,a)\leq\mathrm{dist}_{\mathsf{g}}(w,x)+\mathrm{dist}_{\mathsf{g}}(x,a)$. Taking $\inf_{a\in\mathcal{S}_{f}}$ yields $\mathrm{dist}_{\mathsf{g}}(w,\mathcal{S}_{f})\leq\mathrm{dist}_{\mathsf{g}}(w,x)+\mathrm{dist}_{\mathsf{g}}(x,\mathcal{S}_{f})$. Swapping $w$ and $x$ gives the reverse inequality, hence $\bigl|\mathrm{dist}_{\mathsf{g}}(w,\mathcal{S}_{f})-\mathrm{dist}_{\mathsf{g}}(x,\mathcal{S}_{f})\bigr|\leq\mathrm{dist}_{\mathsf{g}}(w,x)$. we have
$\mathrm{dist}_{\mathsf{g}}(w,\mathcal{S}_{f})\ \geq\ \mathrm{dist}_{\mathsf{g}}(x,\mathcal{S}_{f})-\mathrm{dist}_{\mathsf{g}}(w,x)\ \geq\ r/2-r/4\ =\ r/4\ >0,$
so $w\notin\mathcal{S}_{f}$; and
$\mathrm{dist}_{\mathsf{g}}(w,\mathcal{S}_{f})\leq\mathrm{dist}_{\mathsf{g}}(w,x)+\mathrm{dist}_{\mathsf{g}}(x,y)\leq r/4+r/2=3r/4<r.$
Because $r=\min_{u\in\overline{\mathcal{S}\setminus A_{\varepsilon}}}\mathrm{dist}_{\mathsf{g}}(u,\mathcal{S}_{f})$, the inequality $\mathrm{dist}_{\mathsf{g}}(w,\mathcal{S}_{f})<r$ implies $w\notin\overline{\mathcal{S}\setminus A_{\varepsilon}}$, hence $w\in A_{\varepsilon}$.
(ii) From probability mass to a bound on $r$.
By Section˜3.2, the density $\rho$ of $\mu^{0}$ w.r.t. the Riemannian volume $\mathcal{M}$ satisfies $\kappa\leq\rho\leq\kappa^{-1}$ on $\mathcal{S}$, hence for any Borel $A\subset\mathcal{S}$,
$\kappa\,\mathcal{M}(A)\leq\mu^{0}(A)\leq\kappa^{-1}\mathcal{M}(A).$
Using $\mathbb{B}_{\mathsf{g}}(x;r/4)\subset A_{\varepsilon}$ and $\mu^{0}(A_{\varepsilon})=\delta$, we obtain
$\mathcal{M}\!\bigl(\mathbb{B}_{\mathsf{g}}(x;r/4)\bigr)\ \leq\ \mathcal{M}(A_{\varepsilon})\ \leq\ \kappa^{-1}\mu^{0}(A_{\varepsilon})\ =\ \kappa^{-1}\delta.$
On the other hand, the injectivity–radius and curvature bounds in Section˜3.2 allow us to apply Section˜4.1. Hence, for $r\leq\min\{r_{0},\,1/\sqrt{2K}\}$, we have
$c\,C_{\texttt{k}}\,(r/4)^{\texttt{k}}=c\,{\texttt{Leb}}_{\texttt{k}}\!\bigl(\mathbb{B}_{\texttt{k}}(0;r/4)\bigr)\leq\mathcal{M}\!\bigl(\mathbb{B}_{\mathsf{g}}(x;r/4)\bigr),$
where $C_{\texttt{k}}=\pi^{\texttt{k}/2}/\Gamma(\texttt{k}/2+1)$. Combining the last two inequalities yields
$c\,C_{\texttt{k}}\,(r/4)^{\texttt{k}}\ \leq\ \kappa^{-1}\delta\quad\Longrightarrow\quad r\ \leq\ 4\Bigl(\delta\,\kappa^{-1}[c\,C_{\texttt{k}}]^{-1}\Bigr)^{\frac{1}{\texttt{k}}},$ (13)
provided $\delta\leq c\,C_{\texttt{k}}\,\kappa\,[4^{-1}\min\{r_{0},1/\sqrt{2\mathsf{C}_{0}}\}]^{\texttt{k}}$ so that the small–radius condition of Section˜4.1 holds.
(iii) Lipschitz control of the value gap.
Since $f$ is $L_{f,1}$-Lipschitz (in the ambient norm) and $\mathrm{dist}_{\mathsf{g}}\geq\|\cdot\|$ [^17] Here we use the fact that the geodesic distance $\mathrm{dist}_{\mathsf{g}}(x,y)$ is at least as large as the Euclidean distance $\|x-y\|$ in the ambient space.,
$f^{*}-\text{Q}_{1-\delta}(f(Y))$ $=\varepsilon\ \leq\ f^{*}-f(z)\ \leq\ L_{f,1}\,\|y-z\|\ \leq\ L_{f,1}\,\mathrm{dist}_{\mathsf{g}}(y,z)$
$=L_{f,1}\,r.$ (14)
Together with the bound on $r$ in (13), we conclude
$f^{*}-\text{Q}_{1-\delta}(f(Y))\ \leq\ 4\,L_{f,1}\,\Bigl(\delta\,\kappa^{-1}[c\,C_{\texttt{k}}]^{-1}\Bigr)^{\frac{1}{\texttt{k}}},$
which completes the proof. ∎
Remark 4.3.
(Sharpness of the $\delta^{1/\texttt{k}}$) The $\delta^{1/\texttt{k}}$ dependence in ˜4.2 is dictated by the k-dimensional volume growth of small balls on $\mathcal{S}(\theta)$ and cannot, in general, be improved. Indeed, fix $\theta$ and a point $x^{\ast}(\theta)\in\mathcal{S}(\theta)$, and consider the toy function $f(\theta,x):=-\text{dist}_{\mathsf{g}}(x,x^{\ast}(\theta))$. Clearly, $F_{\max}(\theta)=0$. For $r>0$, the sublevel set $A_{r}(\theta)=\{x\in\mathcal{S}(\theta):-f(\theta,x)\leq r\}$ is exactly the geodesic ball $B_{g}(x^{\ast}(\theta);r)$. Let us pick $r=-\text{Q}_{1-\delta}(f(\theta,Y))$ where $Y\sim\mu_{\theta}^{0}$ and $\delta$ sufficiently small as stated in ˜4.2. By Section˜3.2 and Section˜4.1, we have
$\delta$ $=\mu_{0}(A_{r}(\theta))=\mu_{0}\bigl(\mathbb{B}_{\mathsf{g}}(x^{\ast}(\theta);r)\bigr)\leq\kappa^{-1}\cdot\mathcal{M}(\mathbb{B}_{\mathsf{g}}(x^{\ast}(\theta);r))$
$\leq\kappa^{-1}\cdot\mathsf{c}_{H}\cdot{\texttt{Leb}}_{\texttt{k}}(\mathbb{B}_{\mathsf{g}}(x^{\ast}(\theta);r))\lesssim r^{\texttt{k}}.$ (15)
Then $\delta^{1/\texttt{k}}\lesssim r=F_{\max}(\theta)-\text{Q}_{1-\delta}(f(\theta,Y))$. Thus one cannot, in general, replace the factor $\delta^{1/\texttt{k}}$ in ˜4.2 by a smaller power of $\delta$.
4.2 Approximation of SQ-Loss with SQ-G-Loss
The variational formulations $\phi_{0,\delta}$ and $\phi_{\lambda,\delta}$ involve expectations of $\xi(x):=\max\{f(\theta,x)-\beta,0\}$ over $\mu_{\theta}^{0}$ and $\mu_{\theta}^{\lambda}$, respectively. Using this expectation structure, we bound the difference between $\tilde{F}_{\texttt{SQ}}(\theta)$ and $\tilde{F}_{{\texttt{SQ-G}}}(\theta)$ by the Wasserstein-1 distance between $\mu^{\lambda}_{\theta}$ and $\mu_{\theta}^{0}$: We use the Kantorovich–Rubinstein duality for the Wasserstein-1 distance with $\xi(x)$ as a test function.
Lemma 4.4.
Under Section˜2, for every $\theta\in\Theta$, we have,
$|\tilde{F}_{{\texttt{SQ-G}}}(\theta)-\tilde{F}_{\texttt{SQ}}(\theta)|\leq\frac{L_{f,1}}{\delta}\cdot\mathbb{W}_{1}(\mu_{\theta}^{\lambda},\mu_{\theta}^{0}).$ (16)
The proof of Section˜4.2 is given in Section˜C.2. As a consequence of ˜4.2, Section˜4.2, and Section˜2.1.2, we have that $\tilde{F}_{{\texttt{SQ-G}}}(\theta)$ approximates $F_{\max}(\theta)$ by choosing $\delta$ and $\lambda$ properly.
Theorem 4.5.
Under Sections˜2, 2 and 2, for $\delta<c\cdot C_{\texttt{k}}\cdot\kappa\cdot\left(4^{-1}\min\{1/\sqrt{2\mathsf{C}_{0}},r_{0}\}\right)^{\texttt{k}}$, we have $\left|F_{\max}(\theta)-\tilde{F}_{{\texttt{SQ-G}}}(\theta)\right|\leq\frac{4L_{f,1}\delta^{\frac{1}{\texttt{k}}}}{(\kappa\cdot\mathsf{c}\cdot C_{\texttt{k}})^{\frac{1}{\texttt{k}}}}+\frac{L_{f,1}}{\delta}\cdot K_{0}\lambda^{\frac{1}{2}}.$ (17)
The bound in (17) makes the role of the manifold dimension k explicit: the bias term scales as $\delta^{1/\texttt{k}}$, so larger k requires smaller $\delta$ to reach the same accuracy. This bound also yields a simple rule for choosing the parameters to reach any target accuracy: pick $\delta=\epsilon_{v}^{\texttt{k}}$ and $\lambda=\epsilon_{v}^{2(\texttt{k}+1)}$, where $0<\epsilon_{v}<(\mathsf{c}\cdot C_{\texttt{k}}\cdot\kappa)^{\frac{1}{\texttt{k}}}\cdot 4^{-1}\min\{c_{1}/\sqrt{2\mathsf{C}_{0}},r_{0}\}$, then
$\left|F_{\max}(\theta)-\tilde{F}_{{\texttt{SQ-G}}}(\theta)\right|=\mathcal{O}(\epsilon_{v}).$
This provides a clear, dimension-aware recipe for setting $(\delta,\lambda)$ that we use in Section˜5.
5 Oracle Complexity of Optimizing $F_{\max}$
In this section, we develop a zeroth-order algorithm to find a Goldstein stationary point of the hyper-objective $F_{\max}$. Zeroth-order methods require pointwise evaluations of the objective, but direct evaluation of $F_{\max}$ is intractable because it involves an inner maximization over the lower-level solution set. To obtain a tractable objective oracle, we replace $F_{\max}$ by its Gibbs–superquantile surrogate $\tilde{F}_{{\texttt{SQ-G}}}$ defined in (2), which can be estimated from Gibbs samples. As shown in ˜4.5, suitable choices of $\lambda$ and $\delta$ ensure that $\tilde{F}_{{\texttt{SQ-G}}}$ uniformly approximates $F_{\max}$. We then query $\tilde{F}_{{\texttt{SQ-G}}}$ at multiple points to construct a finite-difference estimator of a proxy gradient of $F_{\max}$, and use this estimator within a zeroth-order scheme to obtain a Goldstein stationary point of the original hyper-objective.
To evaluate $\tilde{F}_{{\texttt{SQ-G}}}(\theta)$, we use its variational representation in (2), which involves solving the inner minimization problem $\min_{\beta\in\mathbb{R}}\phi_{\lambda,\delta}(\theta,\beta)$. Since $\phi_{\lambda,\delta}(\theta,\cdot)$ is convex but possibly nonsmooth, a natural choice is to apply the subgradient method [^18] Because the problem is one-dimensional, the bisection method is also applicable, provided the domain is bounded.. To ensure convergence, the nonsmooth optimization must be restricted to a bounded interval. In Section˜5.1, we show that the optimal set $\operatorname*{arg\,min}_{\beta}\phi_{\lambda,\delta}(\theta,\beta)$ lies within $\mathrm{B},\mathrm{B}]$, so that
$\tilde{F}_{{\texttt{SQ-G}}}(\theta)=\min_{\beta\in[-\mathrm{B},\mathrm{B}]}\phi_{\lambda,\delta}(\theta,\beta).$
Given that (i) $\phi_{\lambda,\delta}(\theta,\cdot)$ is convex and Lipschitz continuous, (ii) the feasible set is bounded, and (iii) the stochastic subgradient $\tilde{\partial}_{\beta}\phi_{\lambda,\delta}(\theta,\beta;X)$ satisfies $|\tilde{\partial}_{\beta}\phi_{\lambda,\delta}(\theta,\beta;X)|\leq\sigma:=1+\delta^{-1}$, we invoke the classical convergence guarantees for convex Lipschitz objectives [51, Proposition 2]. Accordingly, the output $\hat{\beta}$ of projected stochastic gradient descent (PSGD) (see Algorithm˜1) satisfies
$\phi_{\lambda,\delta}(\theta,\hat{\beta})-\min_{\beta}\phi_{\lambda,\delta}(\theta,\beta)\leq\hat{\epsilon},\quad\text{with probability at least }1-\Delta,$ (18)
after at most $\mathcal{O}(\mathrm{B}^{2}\sigma^{2}\hat{\epsilon}^{-2}\log^{2}(\Delta^{-1}))$ iterations.
[…]
Below, we analyze in detail the convergence of Algorithm˜1 (inner minimization of the Gibbs–superquantile) and Algorithm˜2 (outer minimization of the hyper-objective).
5.1 Inner Minimization of Gibbs-superquantile
[…]
We first briefly introduce the PSGD (Algorithm˜1) to solve $\min_{\beta\in[-\mathrm{B},\mathrm{B}]}\phi_{\lambda,\delta}(\theta,\beta)$. Specifically, for a given $\theta\in\Theta$, we have
$\beta^{l+1}\leftarrow\text{Proj}_{[-\mathrm{B},\mathrm{B}]}(\beta^{l}-\eta_{l}\tilde{\partial}_{\beta}\phi_{\lambda,\delta}(\theta,\beta^{l};\tilde{X}^{(l)})),$ (19)
where $\tilde{\partial}_{\beta}\phi_{\lambda,\delta}(\theta,\beta;\tilde{X}^{(l)})$ represents a sub-gradient of the noisy sample of $\phi_{\lambda,\delta}(\theta,\beta)$ as follows:
$\phi_{\lambda,\delta}(\theta,\beta;\tilde{X}^{(l)}):=\beta+\frac{1}{\delta}\max\{f(\theta,\tilde{X}^{(l)})-\beta,0\},$ (20)
and $\{\tilde{X}^{(l)}\}_{l=0}^{L-1}$ are independent and identically distributed (i.i.d.) samples drawn from $\mu_{\theta}^{\lambda}$ and $\mathrm{B}$ is a positive constant. A natural choice for the PSGD output at a fixed $\theta$ is the empirical average $L^{-1}\sum_{i=1}^{L}\beta^{i}$.
We next show in Section˜5.1 that the minimizers $\beta^{\star}(\theta)\in\arg\min_{\beta}\phi_{\lambda,\delta}(\theta,\beta)$ are uniformly bounded. For the upper bound on $\beta^{\ast}(\theta)$, recall the definition of $\tilde{F}_{\texttt{SQ}}(\theta)$ in (11); then
$\beta^{\ast}(\theta)$ $\leq\beta^{\ast}(\theta)+\tfrac{1}{\delta}\,\mathbb{E}_{X\sim\mu_{\theta}^{\lambda}}\!\big[\max\{f(\theta,X)-\beta^{\ast}(\theta),0\}\big]=\phi_{\lambda,\delta}(\theta,\beta^{\ast}(\theta))$
$=\tilde{F}_{{\texttt{SQ-G}}}(\theta)\leq\tilde{F}_{\texttt{SQ}}(\theta)+\big|\tilde{F}_{{\texttt{SQ-G}}}(\theta)-\tilde{F}_{\texttt{SQ}}(\theta)\big|$
$\leq\tilde{F}_{\texttt{SQ}}(\theta)+\tfrac{L_{f,1}}{\delta}\,K_{0}\,\lambda^{1/2}$ (21)
$\leq L_{f,0}(\mathsf{D})+\tfrac{L_{f,1}}{\delta}\,K_{0}\,\lambda^{1/2},$ (22)
where (21) applies Section˜2.1.2 to Section˜4.2, and (22) uses
$\tilde{F}_{\texttt{SQ}}(\theta)\leq\max_{x\in\mathcal{S}(\theta)}f(\theta,x)\leq\max_{x\in\mathbb{B}_{d}(0;\mathsf{D})}f(\theta,x)\leq L_{f,0}(\mathsf{D}),$
with $L_{f,0}(\cdot)$ defined in (5) and $\mathsf{D}$ large enough that $\mathcal{S}(\theta)\subseteq\mathbb{B}_{d}(0;\mathsf{D})$ (see Section˜3.2).
A key part of the proof is a uniform lower bound on $\beta^{\ast}(\theta)$: since $\beta^{*}(\theta)$ is a $(1-\delta)$-quantile, we have $\mathbb{P}\{f(\theta,X)\leq\beta^{\ast}(\theta)\}\geq 1-\delta$ for $X\sim\mu^{\lambda}_{\theta}$. We then prove a tail bound for the Gibbs measure by using the quadratic growth of $g$ beyond $\mathbb{B}_{d}(0;\mathsf{D})$ in Section˜2, which ensures $\mu^{\lambda}_{\theta}(B(0;2\mathsf{D}))\geq\delta_{0}$. Combining both and taking $\delta<\delta_{0}$, it yields
$\mu^{\lambda}_{\theta}\big(\{x:\,f(\theta,x)\leq\beta^{\ast}(\theta)\}\cap B(0;2\mathsf{D})\big)\ \geq\ \delta_{0}-\delta\ >\ 0,$
so there exists $x_{0}\in B(0;2\mathsf{D})$ with $f(\theta,x_{0})\leq\beta^{\ast}(\theta)$. Since $|f(\theta,x)|\leq L_{f,0}(2\mathsf{D})$ on $B(0;2\mathsf{D})$, we conclude $\beta^{\ast}(\theta)\geq-L_{f,0}(2\mathsf{D})$. Together with (22), this shows $\beta^{\ast}(\theta)\in[-\mathrm{B},\mathrm{B}]$ for a suitable $\mathrm{B}<\infty$. Now we state the result in the following lemma. The complete proof is deferred to Section˜D.1.
Lemma 5.1.
Suppose that Section˜2 holds and the condition $\delta\leq 1/2$ hold. For any minimizer $\beta^{*}(\theta)\in\operatorname*{arg\,min}_{\beta}\phi_{\lambda,\delta}(\theta,\beta)$, we have
$|\beta^{*}(\theta)|\leq\mathrm{B}:=\max\left\{L_{f,0}(2\mathsf{D}),L_{f,0}(\mathsf{D})+\frac{L_{f,1}}{\delta}\cdot K_{0}\lambda^{\frac{1}{2}}\right\}.$ (23)
[…]
$\tilde{\psi}(\theta):=\left\{\phi_{\lambda,\delta}(\theta,\hat{\beta};\hat{X}_{I})=\hat{\beta}+\frac{1}{\delta|I|}\sum_{i\in I}\max\{f(\theta,\hat{X}^{(i)})-\hat{\beta},0\}\right\}.$ (24)
Here $\hat{X}_{I}=\{\hat{X}^{(i)}\}_{i\in I}$ denote $M$ i.i.d. samples from the sampling oracle of $\mu^{\lambda}_{\theta}$.
Algorithm 1 $\mathrm{PSGD}(\theta,\beta_{0};(\lambda,\delta))$
Input: $\theta$, $\beta_{0}$, $\lambda$, $\delta$, and $L$
1:$l=0$
2:while $\lambda>0$ and $l\leq L$ do
3:  $\beta^{l+1}=\text{Proj}_{[-\mathrm{B},\mathrm{B}]}(\beta^{l}-\eta_{l}\tilde{\partial}_{\beta}\phi_{\lambda,\delta}(\theta,\beta^{l};\tilde{X}^{(l)}))$ $\{\tilde{X}^{(l)}\}_{l=0}^{L-1}$ are i.i.d samples of $\mu^{\lambda}_{\theta}$.]
4:  $l=l+1$
5:end while
[…]
The next lemma shows that the quantity $\tilde{\psi}(\theta)$ defined in (24) closely approximates $\tilde{F}_{{\texttt{SQ-G}}}(\theta)$. The proof is deferred to Section˜D.2. Before stating it, we briefly explain the role of the high-probability guarantee in (18). The convergence analysis of the outer zeroth-order algorithm is often based on an $L^{2}$-type control of the oracle error, i.e., $\mathbb{E}[|\tilde{\psi}(\theta)-F_{\max}(\theta)|^{2}\mid\theta]$, rather than just on its expectation. An in-expectation convergence result for PSGD would only control the mean optimality gap and does not, in general, translate into a sufficiently tight second-moment bound. In contrast, the high-probability estimate in (18) provides a tail bound on the PSGD optimization error; by choosing the target accuracy $\hat{\epsilon}$ and the failure probability $\Delta$ appropriately, this tail bound can be converted into the required $L^{2}$ control on the oracle error used in Lemma 5.1.
Lemma 5.2.
Suppose that Section˜2 holds. Given $\theta\in\Theta$, by using Algorithm˜1 ($\mathrm{PSGD}(\theta,\beta_{0};(\lambda,\delta))$), with the number of iterations $L=\mathcal{O}\left(\texttt{k}^{2}\mathrm{B}^{2}\epsilon_{v}^{-2-2\texttt{k}}\log^{2}\left(\epsilon_{v}^{-1}\right)\right)$, and $|I|\geq 4\tilde{C}_{f}\epsilon_{v}^{-2-2\texttt{k}}$, we obtain
$\mathbb{E}[|\tilde{\psi}(\theta)-\tilde{F}_{{\texttt{SQ-G}}}(\theta)|^{2}\mid\theta]\leq\epsilon_{v}^{2}.$ (25)
Here $\mathrm{B}$ is defined in Section˜5.1 and $\tilde{C}_{f}$ is a constant.
We now simply show that $\tilde{\psi}(\theta)$ provides a good approximation to $F_{\max}(\theta)$. Applying Lemma 5.1 together with ˜4.5 yields
$\mathbb{E}\bigl[\,\lvert\tilde{\psi}(\theta)-F_{\max}(\theta)\rvert^{2}\bigm|\theta\bigr]\leq\mathbb{E}\bigl[\,\lvert\tilde{\psi}(\theta)-\tilde{F}_{{\texttt{SQ-G}}}(\theta)\rvert^{2}\bigm|\theta\bigr]+\lvert\tilde{F}_{{\texttt{SQ-G}}}(\theta)-F_{\max}(\theta)\rvert^{2}$
$\leq\mathbb{E}\bigl[\,\lvert\phi_{\lambda,\delta}(\theta,\hat{\beta};\hat{X}_{I})-\tilde{F}_{{\texttt{SQ-G}}}(\theta)\rvert^{2}\bigm|\theta\bigr]+\mathcal{O}\bigl(\delta^{2/\texttt{k}}+\lambda\delta^{-2}\bigr)$
$=\mathcal{O}(\epsilon_{v}^{2}),$ (26)
where, in the final bound, we choose $\delta=\epsilon_{v}^{\texttt{k}}$ and $\lambda=\epsilon_{v}^{2(\texttt{k}+1)}$. The Equation (26) will serve as a key ingredient in our later convergence analysis.
5.2 Projected Zeroth-order Algorithm with Minima-Selection
In this subsection, we analyze the projected stochastic zeroth-order method with minima-selection (PSZO-MinSel) introduced in Algorithm 2. Our goal is to quantify how many oracle calls are needed for PSZO-MinSel to reach an approximate Goldstein stationary point of $F_{\max}$.
Algorithm 2 Projected Stochastic Zeroth-order with Minima-selection (PSZO-MinSel)
Initialization: $\beta_{0}$, $\theta_{0}$, $\lambda,\delta<1/2$, and integers $M,N$.
1:for $n$ = 0 to N1N-1 do
2:  Query $M$ i.i.d. samples from the sampling oracle of $\mu_{\theta_{n}}^{\lambda}$, denoted by $\hat{X}_{I}=\{\hat{X}^{(i)}\}_{i\in I}$.
3:  Sample $\{u_{n,t}\}_{t=1}^{b_{u}}$ independently and uniformly from a unit sphere.
4:  $\beta_{n,t}^{+}=\mathrm{PSGD}(\theta_{n}+\rho u_{n,t},\beta_{0};(\lambda,\delta)),$ $\forall t\in[1:b_{u}]$.
5:  $\beta_{n,t}^{-}=\mathrm{PSGD}(\theta_{n}-\rho u_{n,t},\beta_{0};(\lambda,\delta)),$ $\forall t\in[1:b_{u}]$.
6:  $\tilde{\psi}(\theta_{n}+\rho u_{n,t})=\phi_{\lambda,\delta}(\theta_{n}+\rho u_{n,t},\beta_{n,t}^{+};\hat{X}_{I})$.
7:  $\tilde{\psi}(\theta_{n}-\rho u_{n,t})=\phi_{\lambda,\delta}(\theta_{n}-\rho u_{n,t},\beta_{n,t}^{-};\hat{X}_{I})$
8:  $\tilde{\mathsf{g}}_{\rho}(\theta_{n})=\frac{m}{2\rho}\cdot\frac{1}{b_{u}}\sum_{t=1}^{b_{u}}\left[\tilde{\psi}(\theta_{n}+\rho u_{n,t})-\tilde{\psi}(\theta_{n}-\rho u_{n,t})\right]u_{n,t}$
9:  $\theta_{n+1}=\text{Proj}_{\Theta}(\theta_{n}-\tilde{\eta}_{n}\tilde{\mathsf{g}}_{\rho}(\theta_{n}))$
10:end for
In Algorithm˜2, we use the stochastic zeroth-order update
$\theta_{n+1}\leftarrow\text{Proj}_{\Theta}\!\left(\theta_{n}-\tilde{\eta}_{n}\,\tilde{\mathsf{g}}_{\rho}(\theta_{n})\right),$
where $\tilde{\mathsf{g}}_{\rho}(\theta_{n}):=b_{u}^{-1}\sum_{t=1}^{b_{u}}\tilde{\mathsf{g}}_{\rho}(\theta_{n},u_{n,t}),$ and $\tilde{\mathsf{g}}_{\rho}(\theta_{n},u_{n,t}):=\frac{m}{2\rho}\big[\tilde{\psi}(\theta_{n}+\rho u_{n,t})-\tilde{\psi}(\theta_{n}-\rho u_{n,t})\big]\,u_{n,t}$. Here $\tilde{\psi}$ is defined in (24), and $\{u_{n,t}\}_{t=1}^{b_{u}}$ are i.i.d. draws from the uniform distribution on the unit sphere $\mathbb{S}^{m-1}$. The complete procedure is summarized in Algorithm 2. Concretely, lines 4–5 invoke PSGD to approximate the minimisers of the perturbed sub-problems $\min_{\beta}\phi_{\lambda,\delta}(\theta_{n}\!+\!\rho u,\beta)$ and $\min_{\beta}\phi_{\lambda,\delta}(\theta_{n}-\rho u,\beta)$ [^19] If $\tilde{\psi}$ is defined only on $\Theta$, then the perturbed evaluations $\tilde{\psi}(\theta_{n}\pm\rho u_{n,t})$ can be ill-posed when $\theta_{n}\pm\rho u_{n,t}\notin\Theta$. A standard workaround is to use the interiorized feasible set $\Theta_{\rho}:=\{\theta\in\Theta:\theta+\rho\mathbb{B}_{m}(0;1)\subseteq\Theta\}$ and replace $\theta_{n}$ in the oracle queries by $\bar{\theta}_{n}:=\text{Proj}_{\Theta_{\rho}}(\theta_{n})$, so that $\bar{\theta}_{n}\pm\rho u_{n,t}\in\Theta$.. These two inner solutions are combined in line 8 to form the standard two-point stochastic gradient estimator of $\tilde{F}_{{\texttt{SQ-G}}}$, which then drives the projected descent update for $\theta$ in line 9.
To state our main result, we first recall the definitions of the Goldstein $\rho$-subdifferential set and the generalized Goldstein stationary point, as these will be used to formulate the convergence result in Goldstein stationary sense.
Definition 5.3 ([34]).
Denote $\mathbb{B}_{m}(\theta;\rho)=\{v\in\mathbb{R}^{m}:\|v-\theta\|\leq\rho\}$. Given a Lipschitz function $f:\mathbb{R}^{m}\to\mathbb{R}$, a point $\theta\in\mathbb{R}^{m}$ and $\rho\geq 0$, the Goldstein $\rho$-subdifferential of $f$ at $\theta$ is defined as
$\partial_{\rho}f(\theta):=\text{conv}\left(\bigcup_{v\in\mathbb{B}_{d}(\theta;\rho)}\partial f(v)\right).$
Definition 5.4 ([64]).
We say the point $\theta\in\Theta$ is a $(\epsilon,\rho,\eta)$-generalized Goldstein stationary point (GGSP) of the problem (1) if it satisfies
$\min_{g\in\partial_{\rho}f(x)}\|\mathcal{G}_{\Theta}(x,g;\eta)\|\leq\epsilon.$
Here the gradient mapping for a vector $g\in\partial_{\rho}f(x)$ at a point $\theta$ is $\mathcal{G}_{\Theta}(\theta,g;\nu):=\nu^{-1}(\theta-\text{Proj}_{\Theta}(\theta-\nu g))$.
Our main result in this section is summarized in the following theorem; we show that Algorithm˜2 converges to a generalized Goldstein stationary point of $F_{\max}$. More specifically, using the error bound in (26), we show that the error incurred by approximating $F_{\max}$ with $\tilde{\psi}$ in Algorithm˜2 can be efficiently controlled. The proof of ˜5.5 is deferred to Section˜D.3.
Theorem $5.$
Let $\{\theta_{i}\}_{i=1}^{N}$ be the sequence generated by Algorithm˜2. Under Sections˜2, 2 and 2, we obtain $\min_{i\in[1:N]}\mathbb{E}\left[\min_{g\in\partial_{\rho}F_{\max}(\theta_{i})}\|\mathcal{G}_{\Theta}(\theta_{i},g;\eta)\|\right]=\mathcal{O}(\epsilon),$ (27) with $N\geq\mathcal{O}(m\epsilon^{-2}\rho^{-2})$, $b_{u}\geq\mathcal{O}(m^{2}\epsilon^{-2})$, $|I|\geq\tilde{C}_{f}m^{2\texttt{k}+2}(\rho\epsilon)^{-2\texttt{k}-2}$, $\eta\leq\mathcal{O}(m^{-1/2}\rho)$, and $L\geq\mathcal{O}\left(\frac{\texttt{k}^{2}m^{2\texttt{k}+2}\mathrm{B}^{2}}{(\rho\epsilon)^{2\texttt{k}+2}}\cdot\log^{2}\left(m(\rho\epsilon)^{-1}\right)\right).$
Based on ˜5.5, to obtain (27), the total number of samples from $\{\mu_{\theta_{i}}^{\lambda}\}_{i=1}^{N}$ can be computed as follows:
$N\times(b_{u}L+b_{u}|I|)=\mathcal{O}\left(\frac{m^{5+2\texttt{k}}\cdot(\texttt{k}^{2}\mathrm{B}^{2}+\tilde{C}_{f})\cdot\log^{2}\left(m\epsilon^{-1}\rho^{-1}\right)}{\epsilon^{2\texttt{k}+6}\rho^{2\texttt{k}+4}}\right).$ (28)
5.3 Gradient Oracle Complexity under LMC-based Implementation
In the previous subsections, we analyze the sampling oracle complexity to identify an $(\epsilon,\rho,\eta)$-GGSP. For a more concrete complexity result, we analyze the lower-level gradient oracle complexity when using Langevin Monte Carlo (LMC) to implement the Gibbs sampling oracle: To sample from the Gibbs measure in Equation˜3, a common choice is the Langevin dynamics
$\mathrm{d}X(t)=-\partial_{x}g(\theta,X(t))\mathrm{d}t+\sqrt{2\lambda}\mathrm{d}W(t),$ (29)
where $W(t)$ denotes a $d$-dimensional Brownian motion. Note that the Gibbs measure in Equation˜3 is the unique invariant distribution of the above dynamics under mild conditions [17]. In practice, one uses the Euler-Maruyama discretization of the above dynamics, which is also known as the unadjusted LMC method:
$\hat{X}^{k+1}=\hat{X}^{k}-h_{k}\partial_{x}g(\theta,\hat{X}^{k})+\sqrt{2\lambda h_{k}}\,\xi_{k},$ (30)
where $\{\xi_{k}\}_{k=0}^{\infty}$ is a sequence of i.i.d. standard Gaussian random vectors in $\mathbb{R}^{d}$ and $\{h_{k}\}_{k=0}^{\infty}$ is a sequence of step-sizes. We denote by $\hat{\mu}_{n}$ the law of the $n$-th iterate, i.e., $\mathrm{Law}(\hat{X}^{n}):=\hat{\mu}_{n}$.
Our idea is to replace the ideal Gibbs oracle in Algorithms˜1 and 2 by the output of LMC algorithm. While the analyses of this more concrete implementation closely follow the ones in the previous two subsections, we further need to ensure the following two points hold:
  • $\bullet$
    For a sufficiently large $n$, $\hat{\mu}_{n}$ is close to $\mu_{\theta}^{\lambda}$ in Wasserstein-1 distance;
  • $\bullet$
    Conditioned on the first point, replacing Gibbs samples with LMC samples introduces only a controlled bias in the Gibbs–superquantile estimator.
To fulfill the first point, we first use a general inequality relating $\mathbb{W}_{1}$, $\mathbb{W}_{2}$, and the Rényi divergence. By [63, Theorem 1.1],
$\mathbb{W}_{1}(\hat{\mu}_{n},\mu_{\theta}^{\lambda})\;\leq\;\mathbb{W}_{2}(\hat{\mu}_{n},\mu_{\theta}^{\lambda})\;\leq\;2\,C_{\mathrm{PI}}^{1/2}\bigl(e^{\mathbb{R}_{2}(\hat{\mu}_{n},\mu_{\theta}^{\lambda})}-1\bigr)^{1/2},$
where $\mathbb{R}_{2}$ is the order-2 Rényi divergence and $C_{\mathrm{PI}}$ is the Poincaré constant of $\mu_{\theta}^{\lambda}$. To bound $\mathbb{R}_{2}(\hat{\mu}_{n},\mu_{\theta}^{\lambda})$ for the LMC iterates, we invoke the following result.
Proposition 5.6.
[16, Theorem 6.2.9] Consider the LMC iteration
$X^{k+1}=X^{k}-h\nabla G(X^{k})+\sqrt{2h}\,\xi_{k}.$
Assume that $\nabla G$ is $L_{G,2}$-Lipschitz and $\pi\propto\exp(-G)$ satisfies a $C_{PI}$-Poincaré inequality. Then for a sufficiently small target accuracy $\varepsilon$ and an appropriate choice of $h$, we obtain $\mathbb{R}_{2}(\hat{\mu}_{n}\,\|\,\pi)\leq\varepsilon^{2},$ with
$n=\Theta\left(\frac{C_{PI}^{2}L_{G,2}^{2}d}{\varepsilon^{2}}\left(\mathbb{R}_{3}(\hat{\mu}_{0}\,\|\,\pi)^{2}+\log^{2}\frac{1}{\varepsilon}\right)\right)\quad\text{iterations,}$
where $\mathbb{R}_{3}$ denotes the order-3 Rényi divergence.
For the update (30), we pick $G(x):=g(\theta,x)/\lambda$ and $h_{k}=h/\lambda$. Then one has $\mathbb{R}_{2}(\hat{\mu}_{n},\mu_{\theta}^{\lambda})\!\leq\!\varepsilon^{2}$ with $n=\tilde{\mathcal{O}}(dC_{PI}^{2}\lambda^{-2}\varepsilon^{-2})$. Combining this with the preceding Wasserstein bound shows that $\mathbb{W}_{1}(\hat{\mu}_{n},\mu^{\lambda}_{\theta})\leq 2C_{PI}^{\frac{1}{2}}C^{\frac{1}{2}}\varepsilon,$ where $C=2(e^{1/2}-1)$. Moreover, as discussed in the end of Section˜2.1.1, the inverse Poincaré constant $C_{PI}$ in our context is $\mathcal{O}(1)$, independent of $\lambda$ [35].
To establish the second point, we show in Appendix˜E that the additional bias resulting from replacing the ideal Gibbs oracle in Algorithms 1 and 2 by the output of an $n$-step LMC run is bounded by $\delta^{-1}\mathbb{W}_{1}(\hat{\mu}_{n},\mu_{\theta}^{\lambda})$. We then choose $n$ large enough so that this LMC-induced error is at most a constant multiple of the accuracy $\epsilon_{v}$ used in Section˜5.1. With this choice, the total oracle error remains of the same order as in the ideal Gibbs-sampling analysis, so the convergence guarantee of ˜5.5 carries over to the LMC-based implementation with only constant-factor changes. This yields a concrete lower-level gradient oracle complexity, summarized in the following theorem.
Theorem 5.7.
Let $\{\theta_{i}\}_{i=1}^{N}$ be the sequence generated by Algorithm˜4 (refer to Section˜E.2). Under Sections˜2, 2 and 2, we obtain
$\min_{i\in[1:N]}\mathbb{E}\left[\min_{g\in\partial_{\rho}F_{\max}(\theta_{i})}\|\mathcal{G}_{\Theta}(\theta_{i},g;\eta)\|\right]=\mathcal{O}(\epsilon),$ (31)
with $N\geq\mathcal{O}(m\epsilon^{-2}\rho^{-2})$, $b_{u}\geq\mathcal{O}(m^{2}\epsilon^{-2})$, $|I|\geq\tilde{C}_{f}m^{2\texttt{k}+2}(\rho\epsilon)^{-2\texttt{k}-2}$, $\eta\leq\mathcal{O}(m^{-1/2}\rho)$,
$L=\mathcal{O}\left(\frac{\texttt{k}^{2}m^{2\texttt{k}+2}\mathrm{B}^{2}}{(\rho\epsilon)^{2\texttt{k}+2}}\log^{2}(m\epsilon^{-1}\rho^{-1})\right),$
and the number of LMC steps $n=\tilde{\mathcal{O}}(dm^{6+6\texttt{k}}(\epsilon\rho)^{-6-6\texttt{k}})$. Moreover, the total number queries to the lower-level gradient oracle and the upper-level function oracle are
$N\times n\times(b_{u}L+b_{u}|I|)$ $=\mathcal{O}\!\left(\frac{d\,m^{11+8\texttt{k}}\,[\texttt{k}^{2}\,\mathrm{B}^{2}+\tilde{C}_{f}]\log^{2}(m\epsilon^{-1}\rho^{-1})}{\epsilon^{8\texttt{k}+12}\rho^{8\texttt{k}+10}}\right),$
$N\times b_{u}\times|I|$ $=\mathcal{O}\left(\frac{m^{2\texttt{k}+5}}{\rho^{2\texttt{k}+4}\epsilon^{2\texttt{k}+6}}\right),$
respectively.
The LMC layer only changes the cost of each oracle call. The theorem above shows that Algorithm˜2 preserves its intrinsic-dimension complexity guarantees when the ideal Gibbs oracle is replaced by an LMC-based implementation, while making explicit the additional polynomial dependence on the ambient dimension $d$ and the factors $m^{\texttt{k}}$ and $(\rho\epsilon)^{-\texttt{k}}$.
6 Experiments
Before detailing the experiments, we outline our adaptation of Algorithm˜2 to the optimistic formulation (see footnote˜1), the standard in hyperparameter optimization. We will compare this optimistic variant against existing baselines in both synthetic and data hyper-cleaning setups. To approximate $F_{\min}(\theta)\coloneqq\min_{x\in\mathcal{S}(\theta)}f(\theta,x),$ we use the lower-tail superquantile
$\tilde{F}_{\mathrm{LSQ}}(\theta)\coloneqq\max_{\beta}\Bigl\{\,\beta-\delta^{-1}\,\mathbb{E}\bigl[\max\{\beta-f(\theta,X),\,0\}\bigr]\Bigr\}.$
In our implementation of Algorithm˜2, we replace $\phi_{\lambda,\delta}(\theta,\beta)$ with the surrogate
$\psi_{\lambda,\delta}(\theta,\beta)\coloneqq\beta-\delta^{-1}\,\mathbb{E}_{\mu_{g}^{\lambda}(\theta)}\!\left[\max\{\beta-f(\theta,X),\,0\}\right].$
Additionally, in Algorithm˜1 we change the projected gradient descent step in line 2 to a projected gradient ascent step.
Compute budget and reporting protocol.
Our method is zeroth-order and each outer update invokes a Gibbs–superquantile oracle implemented via PSGD and (parallel-chain) Langevin Monte Carlo, so the wall-clock cost per outer iteration can be higher than that of gradient-based baselines. Accordingly, in the main text we report iteration-based results to illustrate algorithmic behavior and minima-selection effects, and in Appendix F we additionally provide compute-matched comparisons under a fixed wall-clock time budget (including per-iteration timings for all methods).
Experiments are run on a single machine with CPU: Apple M1 Pro (10 cores), RAM: 16 GB with no discrete GPU. Code to reproduce all experiments is released at GitHub. Toy-example methods are implemented in JAX, while data hyper-cleaning methods are implemented in PyTorch.
6.1 Toy Example
Optimistic formulation:
We evaluate our algorithm along with three more algorithms in the literature on the optimistic BLO problem as follows
$\min_{\theta\in[-\pi,\pi]}\;\min_{x\in\mathcal{S}(\theta)}\;f(\theta,x),$
where $f(\theta,x)=2\|x\|+\langle x,u(\theta)\rangle$ and
$g_{\texttt{k}}(\theta,x)=\frac{1}{4}\Bigl(\sum_{i=1}^{\texttt{k}+1}x_{i}^{2}-\theta^{2}\Bigr)^{2}-\frac{1}{2}\,\|x\|^{2}.$ (32)
and $u(\theta)=[\cos\theta,\sin\theta,0,\dots,0]^{\top}\in\operatorname{\mathbb{R}}^{d}$. The solution set is the k-dimensional sphere of radius $\sqrt{1+\theta^{2}}$ embedded in $\mathbb{R}^{d}$ along the first $\texttt{k}+1$ coordinates:
$\mathcal{S}(\theta)=\Bigl\{\,x\in\mathbb{R}^{d}:\sum_{i=1}^{\texttt{k}+1}x_{i}^{2}=1+\theta^{2},\;x_{\texttt{k}+2}=\cdots=x_{d}=0\,\Bigr\}.$
So the induced hyper-objective
$F_{\min}(\theta)=\sqrt{1+\theta^{2}}$
has the unique global minimizer $\theta^{*}=0$.
Throughout the experiments, we compare the following algorithms:
  • $\bullet$
    PSZO-MinSel (Algorithm˜2)  adapted to $\min_{x\in\mathcal{S}(\theta)}f(\theta,x)$ instead of pessimistic formulation.
  • $\bullet$
    V-PBGD [77]: joint gradient descent on the penalized objective $\Phi(\theta,x)=f(\theta,x)+\gamma(g(\theta,x)-g^{*}(\theta))$,
  • $\bullet$
    G-PBGD [77]: joint gradient descent on the penalized objective $\Phi(\theta,x)=f(\theta,x)+\tfrac{\gamma}{2}\|\nabla_{x}g(\theta,x)\|^{2}$,
  • $\bullet$
    RHG [29]: truncated reverse‐mode hypergradient with re‐initialization of $\theta$ every outer step,
  • $\bullet$
    IAPTT-GM [62]: unroll $T$ inner steps $x_{k+1}=x_{k}-\beta\nabla_{x}g(\theta,x_{k})$ from $x_{0}=z$, pick the pessimistic truncation $k_{\star}=\arg\max_{k}F(\theta,x_{k})$, and update $\theta,z$ by gradient descent on the truncated objective $F(\theta,x_{k_{\star}})$.
In this part, all methods are implemented in JAX with identical initialization $\theta_{0}=1.0$, $x_{0}\sim\mathcal{N}(0,I_{d})$, and random seed.
Figure 1: Comparison of $F(\theta_{n})-F^{*}$ versus iteration for $d=2$. Our algorithm converges faster with the chosen hyperparameters $\delta=0.1$ and $\lambda=0.01$ in optimistic formulation, while V-PBGD, G-PBGD, RHG, and IAPTT-GM obtain higher values.
Figure 1 plots $F(\theta_{n})-F^{*}$ over iterations for $d=2$. The zeroth-order method rapidly converges toward zero, while the other methods tend to plateau or oscillate around suboptimal objective values, primarily due to the absence of a minima-selection phase.
6.1.1 Dimension Sweep and Minima‐selection
$d$ V-PBGD G-PBGD RHG IAPTT-GM PSZO-MinSel
2 0.297 0.160 0.304 0.152 $\mathbf{0.009\pm 0.001}$
5 0.273 0.160 0.302 0.151 $\mathbf{0.015\pm 0.012}$
10 0.169 0.153 0.303 0.152 $\mathbf{0.023\pm 0.017}$
20 0.113 0.174 0.302 0.154 $\mathbf{0.043\pm 0.025}$
Table 1: Best-so-far absolute error $|\hat{\theta}_{N}-\theta^{*}|$ after $N=500$ outer iterations, where $\hat{\theta}_{N}$ is the selected iterate defined by the best objective value along the trajectory. PSZO-MinSel reports mean $\pm$ 95% CI over 10 runs.
In this experiment, we examine the setting where the intrinsic dimension scales with the ambient dimension, specifically setting $\texttt{k}=d-1$. This simplifies the lower-level objective to $g(\theta,x)=\frac{1}{4}(\|x\|^{2}-\theta^{2})^{2}-\frac{1}{2}\|x\|^{2}$. Table˜1 reports the best-so-far absolute error $|\hat{\theta}_{N}-\theta^{*}|$ after $N=500$ outer iterations across varying dimensions $d$, where $\hat{\theta}_{N}$ denotes the selected iterate with the smallest achieved hyper-objective value along the run. Despite this scaling, our algorithm consistently outperforms the three baseline methods, achieving a significantly closer approximation to the global minimizer $\theta^{\ast}=0$. For robustness, we report the mean and confidence intervals over $10$ independent runs for our randomized method, compared to single runs for the deterministic baselines. Baselines are deterministic given initialization; repeated runs match up to numerical tolerance.
6.1.2 Constant Manifold Dimension
We next investigate the distinction between intrinsic and ambient dimensionality by fixing the manifold dimension at $\texttt{k}=1$ while varying the ambient dimension $d$, using the corresponding instance of $g_{\texttt{k}}(\theta,x)$ as the lower-level objective. To align the implementation with our theoretical bounds, we chose the number of Langevin Monte Carlo (LMC) iterations and the batch size $M$ to closely approximate the target Gibbs distribution. As demonstrated in Table˜2, the best-so-far error of Algorithm˜2 remains nearly constant as $d$ increases, confirming that the complexity is almost independent of the ambient dimension.
This outcome, when combined with the finding that error increases with k, validates our central theoretical prediction that the final optimization error is governed solely by the intrinsic dimensionality of the solution set.
$d$ V-PBGD G-PBGD RHG IAPTT-GM PSZO-MinSel
5 0.495 0.313 0.302 0.119 $\mathbf{0.015\pm 0.014}$
10 0.499 0.313 0.316 0.114 $\mathbf{0.018\pm 0.012}$
20 0.511 0.315 0.307 0.121 $\mathbf{0.018\pm 0.011}$
30 0.532 0.313 0.311 0.121 $\mathbf{0.02\pm 0.02}$
Table 2: Best-so-far absolute error $|\hat{\theta}_{N}-\theta^{*}|$ after $N=500$ outer iterations, with a fixed intrinsic dimension $\texttt{k}=1$.
Pessimistic formulation.
To show that our method is able to handle both the optimistic and pessimistic settings, we solve the previous synthetic problem also in the latter setup. To our knowledge, existing hypergradient and penalty-based methods do not model the minimizer-selection step. They either (i) assume the lower-level solution set is a singleton or (ii) return points satisfying KKT conditions, without guaranteeing that these correspond to a selected minimizer of the original bilevel objective. Thus, the connection between their outputs and the true BLO solution is unclear. We also observe that their algorithms are identical under the optimistic and pessimistic formulations; when the lower-level solution is non-unique, comparing them in both settings is hence not meaningful. Because these methods are designed for the optimistic case, we benchmark against them only in the optimistic setting. For the pessimistic formulation, we report results only for our method.
We now study the following pessimistic toy problem
$\min_{\theta\in[-\pi,\pi]}\;\max_{x\in\mathcal{S}(\theta)}\;f(\theta,x),$ (33)
with $f$ and $g_{\texttt{k}}$ identical to the previous setup. The induced hyper-objective $F_{\max}(\theta)=3\sqrt{1+\theta^{2}}$ admits the unique global minimizer $\theta^{*}=0$.
We also repeat the experiment under a fixed manifold dimension k, using the function instance in (32) as the lower-level objective. In Table˜3 we replicate the last set of experiments exclusively for Algorithm˜2 in the pessimistic setting. In both choices of the lower-level function, we observe a clear dependence on the manifold dimension of the solution set.
$d$ $L$: $g_{d-1}(\theta,x)$ (32) $L$: $g_{1}(\theta,x)$ (32)
5 $\mathbf{0.011\pm 0.010}$ $\mathbf{0.021\pm 0.011}$
10 $\mathbf{0.025\pm 0.020}$ $\mathbf{0.023\pm 0.024}$
20 $\mathbf{0.030\pm 0.025}$ $\mathbf{0.026\pm 0.023}$
30 $\mathbf{0.045\pm 0.029}$ $\mathbf{0.027\pm 0.021}$
Table 3: Best-so-far absolute error $|\hat{\theta}_{N}-\theta^{*}|$ after $N=500$ outer iterations, evaluated using Algorithm˜2 in the pessimistic setting. The first column corresponds to (33) with lower-level $g_{d-1}$ from (32); the second uses the same problem with lower-level $g_{1}$ from (32).
6.2 Data Hyper-cleaning
In this section, we evaluate our method on the data hyper-cleaning benchmark [29, 76], a canonical BLO task widely used in the machine learning community. Given a set of corrupted [^20] A sample is clean if its label equals the true class; otherwise it is corrupted. In MNIST, we corrupt by flipping the label of a fraction $p$ of training points uniformly to a different (incorrect) class. $p$ is the pollute-rate. training samples $\mathcal{D}_{\text{tr}}=\{(d_{i}^{\text{tr}},\,\ell_{i}^{\text{tr}})\}_{i=1}^{N}$, a small clean validation set $\mathcal{D}_{\text{val}}=\{(d_{i}^{\text{val}},\,\ell_{i}^{\text{val}})\}_{i=1}^{M}$, and a clean test set, the goal is to learn a data cleaner that down-weights mislabeled training points so as to improve generalization on unseen data. Let $\theta$ denote the data-cleaner parameters (producing per-example weights) and $x$ the model parameters. We use a sigmoid parameterization for the weights, $\omega_{i}(\theta)\;=\;(1+\exp(-\theta_{i}))^{-1},\ (i=1,\dots,N),$ and constrain $\|\theta\|\leq B$ to avoid exploding weights. The task is cast as the optimistic bilevel problem in the following:
$\min_{\theta,\,x}$ $-\frac{1}{M}\sum_{i=1}^{M}\log P\!\left(\ell_{i}^{\text{val}}\mid d_{i}^{\text{val}};\,x\right)$ (34)
s.t. $\|\theta\|\leq B,\;\;x\in\arg\min_{x^{\prime}\in\mathbb{R}^{d_{y}}}\Bigg[-\frac{1}{N}\sum_{i=1}^{N}\omega_{i}(\theta)\,\log P\!\left(\ell_{i}^{\text{tr}}\mid d_{i}^{\text{tr}};\,x^{\prime}\right)+\frac{\eta}{2}\,\|x^{\prime}\|^{2}\Bigg].$
Setup.
We use MNIST with N=5, [^0] N=5{,}000 training, $M=100$ validation, and 10,00010{,}000 test examples, and uniformly corrupt a fraction $p\in\{0.4,0.6,0.8\}$ of the training labels. In practice, the corrupted/imprecise data are widely available, but clean/precise data are scarce. Thus we intentionally adopt a very small validation set. Much of the literature uses roughly equal training/validation splits, in which case the effect of the corruption rate in the uncleaned training data during training is often not apparent. We compare our zeroth-order method (Algorithm˜2) against V-PBGD and G-PBGD [77], RHG [29], and IAPTT-GM [62], using both a linear model and a two-layer MLP which are standard Neural network models in data hyper-cleaning tasks.
Results.
Table˜4 reports solution quality (test accuracy and F1 score [^21] F1-score is defined as follows: We take clean samples as the positive class and corrupted samples as the negative class. Each sample $i$ is assigned a score $s_{i}=\sigma(\theta_{i})$. We predict “clean” for the top (1p)(1-p)-fraction by $s_{i}$. We compute $\text{Precision}=\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FP}},\ \text{Recall}=\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FN}},$ where $\mathrm{TP}$, $\mathrm{FP}$, and $\mathrm{FN}$ denote true positives, false positives, and false negatives, respectively. The F1 score is the harmonic mean of Precision and Recall, $\text{F1}=\frac{2\,\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}.$ Note that this F1 measures the cleaner’s quality, not the classifier’s accuracy. , averaged over $10$ runs with $95\%$ confidence intervals). Our zeroth-order method (see Algorithm˜2) adapted to the optimistic formulation attains higher accuracy and F1-score in three different pollute rates ($p=40\%,60\%,80\%$). Our method’s high performance stems from the minima-selection step: by choosing the iterate with the lowest validation loss along the trajectory, it produces higher-quality cleaner parameters $\theta$, yields better model parameters $x$, and ultimately improves generalization on the held-out test set. Appendix F additionally reports compute-matched comparisons under a fixed wall-clock time budget.
Table 4: Data hyper-cleaning on MNIST across three noise phases (pollute rate $p\in\{40,60,80\}\%$): test accuracy (%) and F1 score (%) over $10$ runs ($\pm$ 95% CI). F1 reflects data-cleaner quality.
$p$ Method Linear model 2-layer MLP
Test acc. F1 Test acc. F1
40 RHG $55.64\pm 0.45$ $36.5\pm 0.4$ $54.64\pm 0.29$ $36.4\pm 0.2$
IAPTT-GM $60.4\pm 0.22$ $54.3\pm 0.4$ $59.37\pm 0.25$ $52.5\pm 0.4$
G-PBGD $73.89\pm 0.13$ $65.9\pm 0.3$ $72.54\pm 0.32$ $65.8\pm 0.3$
V-PBGD $65.34\pm 0.33$ $59.2\pm 0.2$ $64.77\pm 0.2$ $57.2\pm 0.2$
PSZO-MinSel $\mathbf{82.85\pm 0.32}$ $\mathbf{72.1\pm 0.4}$ $\mathbf{85.31\pm 0.24}$ $\mathbf{73.2\pm 0.5}$
60 RHG $51.24\pm 0.19$ $29.8\pm 0.1$ $52.94\pm 0.42$ $28.5\pm 0.5$
IAPTT-GM $59.52\pm 0.38$ $43.2\pm 0.5$ $58.72\pm 0.35$ $44.2\pm 0.3$
G-PBGD $64.56\pm 0.52$ $50.2\pm 0.5$ $62.65\pm 0.42$ $51.7\pm 0.4$
V-PBGD $63.14\pm 0.43$ $48.7\pm 0.5$ $62.63\pm 0.53$ $50.4\pm 0.4$
PSZO-MinSel $\mathbf{79.50\pm 0.18}$ $\mathbf{55.2\pm 0.1}$ $\mathbf{80.70\pm 0.28}$ $\mathbf{55.9\pm 0.2}$
80 RHG $48.14\pm 0.22$ $21.5\pm 0.2$ $50.32\pm 0.34$ $21.9\pm 0.3$
IAPTT-GM $56.94\pm 0.44$ $27.5\pm 0.3$ $58.24\pm 0.36$ $26.1\pm 0.3$
G-PBGD $46.84\pm 0.32$ $28.9\pm 0.3$ $47.70\pm 0.44$ $27.3\pm 0.3$
V-PBGD $61.23\pm 0.32$ $33.5\pm 0.3$ $60.42\pm 0.42$ $34.9\pm 0.4$
PSZO-MinSel $\mathbf{69.45\pm 0.45}$ $\mathbf{35.2\pm 0.4}$ $\mathbf{70.27\pm 0.48}$ $\mathbf{44.1\pm 0.3}$
7 Conclusion and Future Work
This paper studies minima-selection in bilevel optimization and proposes a method that is both theoretically grounded and implementable. Under the condition, exploiting the manifold structure of the lower-level solution set, allows to introduce a dimension-aware viewpoint: the difficulty of BLO with minima-selection is governed by the intrinsic dimension of the minimizer manifold, rather than the ambient dimension of the lower-level variable. The proposed superquantile–Gibbs (SQ-G) relaxation turns minima-selection into a sampling-based procedure with an explicit selection rule. Overall, the results suggest that minima-selection can be handled in a structured way by combining geometric information (the minimizer manifold), stochastic approximation (Gibbs measures), and risk-type criteria (superquantiles).
Looking ahead, several directions appear particularly promising. On the algorithmic side, a natural next step is to move from zeroth-order schemes to hyper-gradient methods, in which one first selects a point on the minimizer manifold (using ideas similar to the Gibbs-based selection studied here) to maximize [^22] For the optimistic formulation, one can also minimize the upper-level function over the lower-level solution manifold. the upper-level objective, and then differentiates through this choice via implicit differentiation. In parallel, it is important to improve the efficiency of the Gibbs sampling oracle. Our current framework relies on a standard Langevin Monte Carlo implementation; geometry-aware Langevin samplers tailored to the identified manifold structure could sharpen the dependence on both the intrinsic dimension k and the ambient dimension $d$. Such developments would make the proposed PSZO-MinSel algorithm more practical for large-scale applications.
Appendix A Definitions and Technical Lemmas
This appendix collects the background components used throughout the paper. We first fix notation and recall a few basic Riemannian notions for the minimizer manifold (distance, second fundamental form, injectivity radius, and curvature) that are repeatedly used in Section˜3. We then briefly summarize standard differentiability results for bilevel hyper-objectives. Finally, we verify that our assumptions imply the technical conditions required to apply [38, Theorem 3.6] in the discussion of Gibbs-measure convergence.
Additional Notations. We use the notation $(\cdot)_{+}$ to denote the positive part of a scalar or vector, defined as $(\cdot)_{+}:=\max\{\cdot,0\}$.
A.1 Riemannian Notions
In this subsection we fix notation for basic Riemannian objects and record definitions that will be used repeatedly in later proofs. Any compact manifold $\mathcal{S}$ of class $\mathcal{C}^{3}$ (indeed $\mathcal{C}^{1}$) without boundary admits a Riemannian metric [^23] A fundamental theorem in differential geometry states that every paracompact manifold $M$ of class $\mathcal{C}^{k}$ with $k\geq 1$ can be endowed with a Riemannian metric.. Given a Riemannian metric $\mathsf{g}$ on $\mathcal{S}$, define
[…]
By compactness of $\mathcal{S}$, there exists at least one minimizing path $\Gamma^{*}$.
Definition A.1 (Second fundamental form).
Let $\mathcal{S}\subset\mathbb{R}^{d}$ be a $C^{2}$ embedded submanifold endowed with the Riemannian metric induced by the Euclidean inner product. For each $x\in\mathcal{S}$, let $T_{x}\mathcal{S}$ and $T_{x}\mathcal{S}^{\perp}$ denote the tangent space and its orthogonal complement, respectively. The second fundamental form of $\mathcal{S}$ at $x$ is the bilinear map
$\mathrm{I\!I}_{\mathcal{S}}(x):T_{x}\mathcal{S}\times T_{x}\mathcal{S}\to T_{x}\mathcal{S}^{\perp},\qquad\mathrm{I\!I}_{\mathcal{S}}(x)[u,v]:=\bigl(\nabla^{\mathbb{R}^{d}}_{u}v\bigr)^{\perp},$
where $\nabla^{\mathbb{R}^{d}}$ is the Euclidean (Levi–Civita) connection, i.e., the usual directional derivative on $\mathbb{R}^{d}$: for a smooth vector field $v$ defined near $x$, $\nabla^{\mathbb{R}^{d}}_{u}v=Dv(x)[u]$, and $(\cdot)^{\perp}$ denotes orthogonal projection onto $T_{x}\mathcal{S}^{\perp}$.
We measure its size by the operator norm
$\bigl\|\mathrm{I\!I}_{\mathcal{S}}(x)\bigr\|_{\mathrm{op}}:=\sup\bigl\{\|\mathrm{I\!I}_{\mathcal{S}}(x)[u,u]\|:u\in T_{x}\mathcal{S},\ \|u\|=1\bigr\},$
where $\|\cdot\|_{\mathrm{op}}$ denotes the operator norm induced by the Euclidean norm $\|\cdot\|$. This quantity coincides with the maximal normal curvature of $\mathcal{S}$ at $x$. In particular, if $\gamma$ is a unit-speed geodesic in $\mathcal{S}$, then
$\|\ddot{\gamma}(t)\|=\bigl\|\mathrm{I\!I}_{\mathcal{S}}(\gamma(t))[\dot{\gamma}(t),\dot{\gamma}(t)]\bigr\|\leq\bigl\|\mathrm{I\!I}_{\mathcal{S}}(\gamma(t))\bigr\|_{\mathrm{op}}.$
Section˜2 requires that, for each $\theta\in\Theta$, the second fundamental form $\mathrm{I\!I}_{S(\theta)}$ is uniformly bounded in this sense, with a constant $\mathsf{C}>0$ independent of $\theta$. For further background, see also [35, Appendix D].
Definition A.2 (Injectivity radius).
Injectivity radius at a point $p\in\mathcal{S}$ is defined as
$\operatorname{inj}(p)\;=\;\sup\bigl\{\,r>0\;\bigm|\;\exp_{p}\!\colon\,\mathbb{B}_{d-\texttt{k}}(p;r)\subset T_{p}\mathcal{S}\,\to\,\mathbb{B}_{\mathsf{g}}(p;r)\subset\mathcal{S}\text{ is a \emph{diffeomorphism}}\bigr\},$
where $\mathbb{B}_{d-\texttt{k}}(p;r)\subset T_{p}\mathcal{S}$ is the open ball of radius $r$ in the tangent space at $p$, $\exp_{p}$ is the Riemannian exponential map at $p$, $\mathbb{B}_{\mathsf{g}}(p;r)\subset\mathcal{S}$ is the geodesic ball of radius $r$ in $\mathcal{S}$. Equivalently, $\operatorname{inj}(p)$ is the distance from $p$ to the first place where the exponential map fails to be one-to-one. The injectivity radius of the whole manifold is the infimum of the pointwise injectivity radii:
$\operatorname{inj}(\mathcal{S})\;=\;\inf_{p\in\mathcal{S}}\,\operatorname{inj}(p).$
In words, $\operatorname{inj}(\mathcal{S})$ is the largest radius $r$ such that every point in $\mathcal{S}$ has a geodesic ball of radius $r$ on which the exponential map at that point is a diffeomorphism.
Definition A.3 (Sectional Curvature).
In a Riemannian manifold $(\mathcal{S},\mathsf{g})$, the sectional curvature measures how curved the manifold is in each two-dimensional “slice" (or “plane") of the tangent space at a point. Concretely 1) Pick a point $p\in\mathcal{S}$. 2) Pick a 2-dimensional subspace $A_{p}\subset T_{p}\mathcal{S}$. (In other words, $A_{p}$ is a plane in the tangent space at $p$.) The sectional curvature $K(p,A_{p})$ is defined using the Riemann curvature tensor [^24] In a Riemannian manifold $(\mathcal{S},\mathsf{g})$ with Levi-Civita connection $\nabla$, the Riemann curvature tensor is a $(1,3)$-tensor $R$ that measures how much the covariant derivative $\nabla$ fails to commute. $R$. If $X$ and $Y$ are any two linearly independent vectors spanning $A_{p}$, then
$K(p,A_{p})=\frac{\mathsf{g}(R(X,Y)Y,X)}{\mathsf{g}(X,X)\mathsf{g}(Y,Y)-\mathsf{g}(X,Y)^{2}}.$
Equivalently, if you choose an orthonormal basis $\{\,e_{1},e_{2}\}$ of $A_{p}$ with respect to $\mathsf{g}$ (so $\mathsf{g}(e_{i},e_{j})=\delta_{ij}$), then the denominator becomes $1$, and one obtains the simpler formula
$K(p,A_{p})=\mathsf{g}(R(e_{1},e_{2})\,e_{2},\,e_{1}).$
A.2 Differentiability of Hyper-objective
The authors of [50, 14] reformulate BLO via the classical value-function approach and assume that a PŁ-type condition (or an error bound) holds for the penalized objective $\sigma f+g$, uniformly over $\sigma\in[0,\sigma_{0}]$ where $\sigma_{0}>0$ is a constant. More precisely, Kwon et al. [50] impose a proximal error bound for $\sigma f+g$, whereas Chen et al. [14] assume a global PŁ condition for $\sigma f+g$; Chen et al. [14, Proposition C.1] show these two assumptions are equivalent.
Under this global regularity, both works establish a Danskin-type result implying that the hyper-objective $F_{\max}$ is differentiable. However, their conclusion also entails that every choice $x\in\mathcal{S}(\theta)$ yields the same hyper-gradient $\nabla F_{\max}(\theta)$, so minima-selection becomes redundant. Finally, requiring a uniform PŁ condition (or error bound) for $\sigma f+g$ is quite stringent and may fail even when $f$ and $g$ individually satisfy PŁ-type properties.
A.3 Discussion on The Validity of Assumptions in [38, Theorem 3.6]
The assumptions required by [38, Theorem 3.6] follow from our standing conditions. First, their positive definiteness requirement for $g(\theta,\cdot)$ holds because the condition implies that $g(\theta,\cdot)$ is locally Morse–Bott [73, Section 2.3], and hence the Hessian of $g$ is positive definite in directions normal to the minimizer manifold $\mathcal{S}(\theta)$. Second, their tail condition [38, Assumption 3.1] is ensured by our coercivity assumption on $g$.
Finally, Sections˜2 and 2 verify the geometric conditions in [38, Assumption 2.2] (smooth, boundaryless, and connected minimizer manifolds). It is worth noting that the orientability of the manifold is not necessary in our setting. We define the volume measure $\mathcal{M}$ in a coordinate-free manner based on the Riemannian metric $g_{\mathcal{S}}$. Orientability, a stronger condition, ensures the existence of a global top-form rather than just a density, which is not required in our framework.
Appendix B Proofs of Results in Section˜3
This appendix contains deferred proofs for results in Section˜3 concerning the structure and regularity of the minimizer sets $\{\mathcal{S}(\theta)\}_{\theta\in\Theta}$.
B.1 Proof of Section˜3.1
Proof.
Consider the decreasing sublevel sets $E_{\varepsilon}:=\{x:\ g(x)-g_{\star}\leq\varepsilon\}$; if none is contained in $\mathcal{U}(\theta)$, compactness yields a limit point in $\mathcal{S}(\theta)\cap(\mathbb{R}^{d}\setminus\mathcal{U}(\theta))$, a contradiction.
We provide a detailed proof here. Fix $\theta$ and write $g(x):=g(\theta,x)$, $g_{\star}:=\min_{z}g(z)$, and $\mathcal{S}:=\arg\min_{z}g(z)$. Let $\mathcal{U}=\mathcal{U}(\theta)$ be the neighborhood in (9). Consider the nested sublevel sets
$E_{j}:=\{x:\ g(x)\leq g_{\star}+1/j\},\qquad j\in\mathbb{N}.$
By continuity of $g$, each $E_{j}$ is closed and $E_{j+1}\subseteq E_{j}$, and clearly $\bigcap_{j\geq 1}E_{j}=\mathcal{S}$.
By coercivity of $g$ (from Section˜2), $E_{1}$ is compact, hence the family $\{E_{j}\}$ is a nested sequence of nonempty compact sets. If the claim were false, then for every $j$ we could pick $x_{j}\in E_{j}\setminus\mathcal{U}$. By compactness of $E_{1}$, there exists a subsequence $x_{j_{k}}\to\bar{x}\in E_{1}$. Since the sets are nested, $\bar{x}\in E_{\ell}$ for every fixed $\ell$, so $\bar{x}\in\bigcap_{\ell}E_{\ell}=\mathcal{S}$. On the other hand, $\mathbb{R}^{d}\setminus\mathcal{U}$ is closed, hence $\bar{x}\notin\mathcal{U}$, contradicting $\mathcal{S}\subset\mathcal{U}$.
Therefore, there exists $j_{0}$ such that $E_{j_{0}}\subset\mathcal{U}$. Setting $\tilde{\epsilon}:=1/j_{0}$ yields $\{x:\ g(x)-g_{\star}\leq\tilde{\epsilon}\}\subset\mathcal{U},$ as claimed. ∎
B.2 Proof of Section˜3.1
The following lemma formalizes a simple but useful observation: if a map is $L_{G}$–Lipschitz on every pair of points that are sufficiently close (i.e., within a fixed radius $r$), then the same Lipschitz constant extends globally. This will be used in the proof of Section˜3.1.
Lemma B.1.
For a given $r>0$, if $G:\mathbb{R}^{d_{1}}\to\mathbb{R}^{d_{2}}$ is a functional that satisfies $\|G(x)-G(y)\|\leq L_{G}\|x-y\|$ for $x,y\in\mathbb{R}^{d_{1}}$ with $\|x-y\|\leq r$, then $G$ satisfies $L_{G}$-Lipschitzness over $\mathbb{R}^{d_{1}}$.
Proof.
We aim to show that $\|G(x)-G(y)\|\leq L_{G}\|x-y\|$ for all $x,y\in\mathbb{R}^{d_{1}}$. If $\|x-y\|\leq r$, our proof is complete. Assume $\|x-y\|=R$ where $R>r$. Define $z_{t}=(1-t)\cdot x+t\cdot y$ for $t\in[0,1]$. Let $w_{k}:=z_{k/({2R/r})}$ for k=[1:2R/r]k=[1:{2R/r}]. Notice that $\|w_{k}-w_{k-1}\|=\|x-y\|/{2R/r}\leq r/2$. Then
$\|G(x)-G(y)\|=\left\|\sum_{k=1}^{{2R/r}}G(w_{k})-G(w_{k-1})\right\|\leq\sum_{k=1}^{{2R/r}}\left\|G(w_{k})-G(w_{k-1})\right\|$
$\leq L_{G}\sum_{k=1}^{{2R/r}}\left\|z_{k}-z_{k-1}\right\|=L_{G}\sum_{k=1}^{{2R/r}}\frac{k}{{2R/r}}\cdot\left\|x-y\right\|=L_{G}\left\|x-y\right\|.$ (36)
For completeness, we provide a proof of Section˜3.1:
Proof.
For any $\theta_{1},\theta_{2}\in\Theta$ satisfying $\|\theta_{1}-\theta_{2}\|\leq\delta$ with $\delta\leq\tilde{\epsilon}/(2L_{g,1}(\mathsf{D}))$ (where $\tilde{\epsilon}$ is defined in Section˜3.1), we obtain the bound:
$\mathrm{dist}(\mathcal{S}(\theta_{1}),\mathcal{S}(\theta_{2}))\leq\frac{L_{g,2}}{\mu}\|\theta_{1}-\theta_{2}\|.$
Selecting $x_{1}\in\arg\max_{x\in\mathcal{S}(\theta_{1})}f(\theta_{1},x)$ and $x_{2}\in\arg\min_{y\in\mathcal{S}(\theta_{2})}\mathrm{dist}(y,x_{1})$, we derive
$F_{\max}(\theta_{1})-F_{\max}(\theta_{2})=f(\theta_{1},x_{1})-F_{\max}(\theta_{2})\leq f(\theta_{1},x_{1})-f(\theta_{2},x_{2})$
$\leq|f(\theta_{1},x_{1})-f(\theta_{2},x_{1})|+|f(\theta_{2},x_{1})-f(\theta_{2},x_{2})|$
$\overset{(a)}{\leq}L_{f,1}\|\theta_{1}-\theta_{2}\|+L_{f,1}\|x_{1}-x_{2}\|$
$\overset{(b)}{\leq}L_{f,1}\|\theta_{1}-\theta_{2}\|+L_{f,1}\mathrm{dist}(\mathcal{S}(\theta_{1}),\mathcal{S}(\theta_{2}))$
$\overset{(c)}{\leq}L_{f,1}\|\theta_{1}-\theta_{2}\|+L_{f,1}\cdot\frac{L_{g,2}}{\mu}\cdot\|\theta_{1}-\theta_{2}\|.$ (37)
Step (a) follows from $L_{f,1}$-Lipschitzness of $f$. Step (b) follows from
$\|x_{1}-x_{2}\|=\mathrm{dist}(x_{1},\mathcal{S}(\theta_{2}))\leq\max_{x\in\mathcal{S}(\theta_{1})}\mathrm{dist}(x,\mathcal{S}(\theta_{2}))\leq\mathrm{dist}(\mathcal{S}(\theta_{1}),\mathcal{S}(\theta_{2})).$
Step (c) follows from Section˜3.1 under the condition $\delta\leq\tilde{\epsilon}/(2L_{g,1}(\mathsf{D}))$.
A similar argument shows that $F_{\max}(\theta_{2})-F_{\max}(\theta_{1})\leq(L_{f,1}+L_{f,1}\cdot\frac{L_{g,2}}{\mu})\|\theta_{1}-\theta_{2}\|$. Thus, for all $\theta_{1},\theta_{2}\in\Theta$ with $\|\theta_{1}-\theta_{2}\|\leq\delta$,
$|F_{\max}(\theta_{1})-F_{\max}(\theta_{2})|\leq\left(L_{f,1}+L_{f,1}\frac{L_{g,2}}{\mu}\right)\cdot\|\theta_{1}-\theta_{2}\|.$
In other words, $F_{\max}$ is locally Lipschitz on $\Theta$ with the same Lipschitz constant and a radius $\delta$ that do not depend on the base point $\theta$. As explained in Section˜B.2, such a uniform local Lipschitz property on $\Theta$ implies that $F_{\max}$ is globally (uniformly) Lipschitz on $\Theta$ with the same constant, which proves the claim. ∎
B.3 Proof of Section˜3.2
This subsection proves that all minimizer manifolds $\mathcal{S}(\theta)$ share a common intrinsic dimension. The key step is to show that the Hessian rank (equivalently, the normal-space dimension) cannot change under small perturbations of $\theta$, and hence is constant on $\Theta$. We provide the proof of Section˜3.2 in the following.
Proof.
Fix $\theta\in\Theta$. By [73, Corollary 2.13] and the path-connectedness of $\mathcal{S}(\theta)$, the rank of $\partial_{x}^{2}g(\theta,x)$ is constant over $\mathcal{S}(\theta)$. Denote it by
$\mathrm{r}(\theta):=\mathop{\rm rank}\nolimits\bigl(\partial_{x}^{2}g(\theta,x)\bigr),\qquad x\in\mathcal{S}(\theta),$
then the dimension of $\mathcal{S}(\theta)$ is $\texttt{k}(\theta)=d-\mathrm{r}(\theta)$.
We show that $\mathrm{r}(\theta)$ is constant on $\Theta$ by contradiction. Suppose not. Then there exist $\theta_{n}\to\bar{\theta}$ with $\mathrm{r}(\theta_{n})\neq\mathrm{r}(\bar{\theta})$. Pick $x_{n}\in\mathcal{S}(\theta_{n})$. By the Lipschitz continuity of the solution mapping (Section˜3.1), there exists $\bar{x}\in\mathcal{S}(\bar{\theta})$ such that $\|x_{n}-\bar{x}\|\to 0$. Since $\partial_{x}^{2}g$ is continuous in $(\theta,x)$ (Section˜2),
$\bigl\|\partial_{x}^{2}g(\theta_{n},x_{n})-\partial_{x}^{2}g(\bar{\theta},\bar{x})\bigr\|\to 0.$
Moreover, implies a local Morse–Bott structure for $g(\theta,\cdot)$ [73, Corollary 2.13], hence the nonzero eigenvalues of $\partial_{x}^{2}g(\bar{\theta},x)$ are bounded away from 0 uniformly for $x$ near $\mathcal{S}(\bar{\theta})$. Therefore, for $n$ large enough, the perturbation above cannot change the number of nonzero eigenvalues, and thus
$\mathop{\rm rank}\nolimits\bigl(\partial_{x}^{2}g(\theta_{n},x_{n})\bigr)=\mathop{\rm rank}\nolimits\bigl(\partial_{x}^{2}g(\bar{\theta},\bar{x})\bigr)=\mathrm{r}(\bar{\theta}).$
Since the rank is constant on each $\mathcal{S}(\theta_{n})$, the left-hand side equals $\mathrm{r}(\theta_{n})$, contradicting $\mathrm{r}(\theta_{n})\neq\mathrm{r}(\bar{\theta})$. Hence $\mathrm{r}(\theta)$ is constant on $\Theta$, and so is $\texttt{k}(\theta)=d-\mathrm{r}(\theta)$. ∎
B.4 Proof of Section˜3.2
Using the Lipschitz continuity of the solution mapping established in Section˜3.1, we first show that the family of sets $\{\mathcal{S}(\theta)\}_{\theta\in\Theta}$ is uniformly bounded. [^25] A collection of sets $\{\mathcal{B}(\theta)\}_{\theta\in\Theta}$ is uniformly bounded if there exists a ball $\mathbb{B}_{d}(0;R)\subset\mathbb{R}^{d}$ such that $\mathcal{B}(\theta)\subset\mathbb{B}_{d}(0;R)$ for all $\theta\in\Theta$. This lemma will be used in the proof of Property 2 in Section˜3.2 and in the proof of Section˜3.2.
Lemma B.2.
A collection of sets $\{\mathcal{A}(\theta)\}_{\theta\in\Theta}$ is uniformly bounded if $\mathcal{A}(\theta)$ is a compact and non-empty set for every $\theta\in\Theta$, $\Theta$ is a compact set, and $\mathcal{A}(\theta)$ is continuous in Hausdorff distance.
Proof.
We aim to show that $\sup_{\theta\in\Theta}\sup_{x\in\mathcal{A}(\theta)}\|x\|$ is finite. Let define $M(\theta):=\sup_{x\in\mathcal{A}(\theta)}\|x\|$. Since $\mathcal{A}(\theta)$ is compact and non-empty, $M(\theta)$ is finite for each $\theta\in\Theta$. We will show that $M(\theta)$ is continuous and since $\Theta$ is compact, $M(\theta)$ attains its maximum value on $\Theta$.
The mapping $\theta\mapsto\mathcal{A}(\theta)$ is continuous in the Hausdorff metric. Hence given $\|\theta_{1}-\theta_{2}\|\leq\delta$, for any $x\in\mathcal{A}(\theta_{2})$, there exists $y\in\mathcal{A}(\theta_{1})$ such that $\|x-y\|<\epsilon$. Then
$\|x\|=\|y+(x-y)\|\leq\|y\|+\|x-y\|<\|y\|+\epsilon.$
Taking the supremum over all $y\in\mathcal{A}(\theta_{1})$:
$\|x\|<M(\theta_{1})+\varepsilon.$
Since this holds for all $x\in\mathcal{A}(\theta_{2})$, we have:
$M(\theta_{2})=\sup_{x\in\mathcal{A}(\theta_{2})}\|x\|\leq M(\theta_{1})+\varepsilon.$
By similar argument, $M(\theta_{1})\leq M(\theta_{n})+\varepsilon$. Then
$|M(\theta_{2})-M(\theta_{1})|\leq\varepsilon.$
We then provide a proof for Section˜3.2 in the following:
Proof.
Property 1. Since we know that $g(\theta,\cdot)$ is coercive, it follows that for each $\theta\in\Theta$, the diameter of $\mathcal{S}(\theta)$ is finite. We aim to prove that $\sup_{\theta\in\Theta}\text{diam}(\mathcal{S}(\theta))$ is also bounded. Given Section˜2, from Section˜3.1, we deduce that $\mathcal{S}(\theta)$ is continuous with respect to $\theta$ in the Hausdorff distance. Using this property, we will further establish that the diameter of $\mathcal{S}(\theta)$ is a continuous function of $\theta$. We aim to show that for every $\epsilon>0$, there exists $\delta\in\mathbb{R}^{m}_{+}$ such that $|\text{diam}(\mathcal{S}(\theta+\delta))-\text{diam}(\mathcal{S}(\theta))|\leq\epsilon$.
[…]
$\text{diam}(\mathcal{S}(\theta))-\text{diam}(\mathcal{S}(\theta+\delta))=\sup_{x,y\in\mathcal{S}(\theta)}\|x-y\|-\sup_{x^{\prime},y^{\prime}\in\mathcal{S}(\theta+\delta)}\|x^{\prime}-y^{\prime}\|$
$\overset{(a)}{\leq}\sup_{x,y\in\mathcal{S}(\theta)}\{\|x-y\|-\|\tilde{x}-\tilde{y}\|\}$
$\overset{(b)}{\leq}\sup_{x,y\in\mathcal{S}(\theta)}\{\|x-\tilde{x}\|+\|\tilde{y}-y\|\}$
$\leq\sup_{x,y\in\mathcal{S}(\theta)}\text{dist}(\mathcal{S}(\theta),\mathcal{S}(\theta+\delta))\leq\epsilon,$ (38)
where (a) comes from $(\tilde{x},\tilde{y})\in\mathcal{S}(\theta+\delta)\times\mathcal{S}(\theta+\delta)$. (b) follows by $\|x-y\|-\|\tilde{x}-\tilde{y}\|\leq\|x-y-\tilde{x}+\tilde{y}\|\leq\|x-\tilde{x}\|+\|\tilde{y}-y\|$. The last inequality holds based on the continuity of the solution sets in Hausdorff distance with respect to $\theta$, i.e., there exist $\delta\in\mathbb{R}_{+}^{m}$ such that $\text{dist}(\mathcal{S}(\theta+\delta),\mathcal{S}(\theta))\leq\epsilon$.
Similarly, for the reverse case, we have
$\text{diam}(\mathcal{S}(\theta+\delta))-\text{diam}(\mathcal{S}(\theta))\leq\epsilon$
Therefore, for every $\epsilon>0$ there exists $\delta\in\mathbb{R}^{m}_{+}$ such that
$|\text{diam}(\mathcal{S}(\theta+\delta))-\text{diam}(\mathcal{S}(\theta))|\leq\epsilon.$
According to Weierstrass’s extreme value theorem, for the continuous function $\text{diam}(\mathcal{S}(\theta))$, there exists a $\theta_{0}\in\Theta$ such that $\text{diam}(\mathcal{S}(\theta_{0}))$ equals the supremum of $\text{diam}(\mathcal{S}(\theta))$ over all $\theta\in\Theta$. We denote $\text{diam}(\mathcal{S}(\theta_{0}))$ by $\mathsf{D}$.
Property 2. We will show that there exists a constant $\mathsf{C}_{0}>0$ such that $|\mathrm{Sec}(\mathcal{S}(\theta))|\leq\mathsf{C}_{0}$ for all $\theta\in\Theta$. In general, for every $\theta\in\Theta$, the sectional curvature of a compact manifold is bounded both above and below as stated in the following result.
Lemma B.3 ([9], Section 9.3).
If $\mathcal{M}$ is a compact Riemannian manifold, then there exist $k,K\in\mathbb{R}$ such that for any plane section $A_{p}$ at any point $p\in\mathcal{M}$, $k\leq K(A_{p})\leq K$.
For any fixed $\theta\in\Theta$, the manifold $\mathcal{S}(\theta)$ is a compact, boundaryless $C^{2}$ submanifold (by Section˜2). As a consequence of general Riemannian geometry (see Section˜B.4), the sectional curvature $\mathrm{Sec}(\mathcal{S}(\theta))$ is bounded on $\mathcal{S}(\theta)$. The sectional curvature $K_{\mathcal{S}(\theta)}(x,A_{x})$ is a continuous function of both the position $x\in\mathcal{S}(\theta)$ and the upper-level parameter $\theta\in\Theta$. This continuity holds because the manifold metrics, coordinate functions, and tangent spaces depend continuously and smoothly on $\theta$ (stemming from the regularity assumptions on $g$ in Section˜2).
Since the parameter domain $\Theta$ is compact, the entire collection of manifolds $\{\mathcal{S}(\theta)\}_{\theta\in\Theta}$ is uniformly bounded (Section˜B.4) and varies continuously, implying that the space of all possible geometric configurations (i.e., all points $x$ and all planes $A_{x}$) forms a compact set in the corresponding product space $\Theta\times\mathcal{S}(\theta)$. By the Weierstrass Extreme Value Theorem, a continuous function on a compact set attains its maximum. Therefore, the absolute value of the sectional curvature attains a global maximum, $\mathsf{C}_{0}$, establishing the uniform bound $|\mathrm{Sec}(\mathcal{S}(\theta))|\leq\mathsf{C}_{0}$ for every $\theta\in\Theta$.
Property 3. Given the boundaryless and compactness properties of $\mathcal{S}$ as well as the $\mathcal{C}^{2}$ property (see Section˜2), we can apply the following result to establish Property 3.
Proposition B.4 ([71], Lemma 6.4.7).
Let $\mathcal{S}$ be a compact, boundaryless Riemannian manifold of dimension k with a $\mathcal{C}^{2}$ metric. Then if $\mathrm{Sec}(\mathcal{S})\leq\mathsf{C}_{0}$ for $\mathsf{C}_{0}>0$,
$\mathrm{inj}(\mathcal{S})\geq\min\left\{\frac{\pi}{\mathsf{C}_{0}},\frac{1}{2}\ell(\mathcal{S})\right\},$
where $\ell(\mathcal{S})$ is the infimum of lengths of closed geodesics in $\mathcal{S}$.
In order to give a lower bound on $\ell(\mathcal{S})$, we know that for any unit-speed closed geodesic $\gamma\in\mathcal{S}$, the extrinsic curvature of $\gamma$ is given by $|\mathsf{g}(\dot{\gamma},\dot{\gamma})|\leq\mathsf{C}$ where $\mathsf{C}$ is the uniform bound on second fundamental form in Section˜2, then
$\ell(\gamma)\geq\frac{2\pi}{\mathsf{C}}.$
Thus $\ell(\mathcal{S})\geq\frac{2\pi}{\mathsf{C}}$.
From Property 2, we have $|\mathrm{Sec}(\mathcal{S}(\theta))|\leq\mathsf{C}_{0}$ for every $\theta\in\Theta$. Additionally, since $\mathsf{C}$ uniformly bounds the second fundamental form of $\{\mathcal{S}(\theta)\}_{\theta\in\Theta}$, the injectivity radius satisfies $\mathrm{inj}(\mathcal{S}(\theta))\geq\min\left\{\frac{\pi}{\mathsf{C}_{0}},\frac{\pi}{\mathsf{C}}\right\}$. ∎
B.5 Proof of Section˜3.2
To establish the uniform bounds on the limiting density $\rho_{\theta}$ in (7), it is necessary to control the $\theta$-dependent normalizing factor, which involves the Riemannian volume $\mathcal{M}\!\left(\mathcal{S}(\theta)\right)$. We therefore first show that the volumes of the minimizer manifolds are uniformly bounded away from 0 and $\infty$ over $\theta\in\Theta$.
Lemma B.5.
Under Section˜2, there exists a constant $0<\mathsf{b}<1$ such that
$\mathsf{b}\ \leq\ \mathcal{M}\bigl(\mathcal{S}(\theta)\bigr)\ \leq\ \mathsf{b}^{-1},\qquad\forall\,\theta\in\Theta,$
where $\mathcal{M}$ denotes the Riemannian volume measure on $(\mathcal{S}(\theta),\mathsf{g})$.
Proof.
For each fixed $\theta$, $\mathcal{S}(\theta)$ is a compact k-dimensional Riemannian manifold, hence $0<\mathcal{M}(\mathcal{S}(\theta))<\infty$ by the standard coordinate formula for the Riemannian volume element; see [53, Section “Riemannian volume”].
It remains to make the bounds uniform over $\Theta$. Define $V(\theta)\;:=\;\mathcal{M}\bigl(\mathcal{S}(\theta)\bigr).$ By Section˜B.4, the family $\{\mathcal{S}(\theta)\}_{\theta\in\Theta}$ is uniformly bounded, i.e., there exists $\mathsf{D}<\infty$ such that $\mathcal{S}(\theta)\subset\mathbb{B}_{d}(0;\mathsf{D})$ for all $\theta\in\Theta$. Since $\theta\mapsto\mathcal{S}(\theta)$ is continuous in Hausdorff distance and the manifolds are $\mathcal{C}^{1}$ (indeed $\mathcal{C}^{2}$) with the induced metric varying continuously in $(\theta,x)$ on $\Theta\times\mathbb{B}_{d}(0;\mathsf{D})$, the volume functional $V(\theta)$ depends continuously on $\theta$.
Because $\Theta$ is compact and $V(\theta)>0$ for all $\theta$, the extreme value theorem yields
$0<V_{\min}:=\min_{\theta\in\Theta}V(\theta)\quad\text{and}\quad V_{\max}:=\max_{\theta\in\Theta}V(\theta)<\infty.$
Setting $\mathsf{b}:=\min\{V_{\min},\,V_{\max}^{-1}\}$ gives $\mathsf{b}\leq V(\theta)\leq\mathsf{b}^{-1}$ for all $\theta\in\Theta$. ∎
Now we provide the proof of Section˜3.2:
Proof.
To establish the boundedness of the density $\rho_{\theta}$, it suffices to show that $|\partial_{r}^{2}g(\theta,y(u,r))|$ is bounded. Since the condition ensures that $g(\theta,\cdot)$ is locally Morse-Bott (see [73, Section 2.3]), we have $\partial_{r}^{2}g(\theta,y(u,r))\big|_{r=0}\succeq\mu\cdot I_{(d-\texttt{k})\times(d-\texttt{k})}$.
Note that $y(u,r)$ is linear in $r$ and we have
$\partial_{r}g(\theta,y(u,r))=\partial_{r}y(u,r)\partial_{x}g(\theta,y(u,r))=\mathsf{N}\partial_{x}g(\theta,y(u,r)),$
where $\mathsf{N}:=[v_{\texttt{k}+1}(u),\ldots,v_{d}(u)]^{T}$. Then
$\partial_{r}^{2}g(\theta,y(u,r))=\mathsf{N}\,\partial_{x}^{2}g(\theta,y(u,r))\mathsf{N}^{T}.$
When $r=0$, $y(u,r)=u$, then for every $u\in\mathcal{S}(\theta)$, $\partial_{r}^{2}g(\theta,y(u,r))|_{r=0}=\mathsf{N}\,\partial_{x}^{2}g(\theta,u)\mathsf{N}^{T}$.
From Section˜2, $\partial_{x}^{2}g(\theta,x)$ is continuous with respect to $(\theta,x)$. Since $\Theta$ is a compact set, the collection $\{\mathcal{S}(\theta)\}_{\theta\in\Theta}$ is uniformly bounded (see Section˜B.4). By Weierstrass’s extreme value theorem, $\|\partial_{x}^{2}g(\theta,u)\|$ achieves its supremum at some $(\theta_{0},u_{0})\in\Theta\times\mathcal{S}(\theta_{0})$, denoted by $\|\partial_{x}^{2}g(\theta_{0},u_{0})\|=\mathsf{M}_{g}$. Then $\partial_{r}^{2}g(\theta,y(u,r))|_{r=0}\preceq\mathsf{M}_{g}I_{(d-\texttt{k})\times(d-\texttt{k})}$
Now, recall the definition of the density $\rho_{\theta}(u)$:
$\rho_{\theta}(u)=\frac{\mathrm{d}\mu^{0}_{\theta}}{\mathrm{d}\mathcal{M}_{\theta}}=\frac{\left|\partial_{r}^{2}g(\theta,y(u,r))|_{r=0}\right|^{-\frac{1}{2}}}{\int_{\mathcal{S}(\theta)}\left|\partial_{r}^{2}g(\theta,y(v,r))|_{r=0}\right|^{-\frac{1}{2}}\mathrm{d}\mathcal{M}_{\theta}(v)}.$
Using the lower and upper bounds on $\partial_{r}^{2}g(\theta,y(u,r))|_{r=0}$, we obtain:
$\frac{\mathsf{M}_{g}^{-\frac{d-\texttt{k}}{2}}}{\mu^{-\frac{d-\texttt{k}}{2}}\mathcal{M}_{\theta}(\mathcal{S}(\theta))}\leq|\rho_{\theta}(u)|\leq\frac{\mu^{-\frac{d-\texttt{k}}{2}}}{\mathsf{M}_{g}^{-\frac{d-\texttt{k}}{2}}\mathcal{M}_{\theta}(\mathcal{S}(\theta))}.$
From Section˜B.5, we have $\mathsf{b}\leq\mathcal{M}_{\theta}(\mathcal{S}(\theta))\leq\mathsf{b}^{-1}$. Finally, we define $\kappa:=\mathsf{b}\cdot\left(\frac{\mu}{\mathsf{M}_{g}}\right)^{\frac{d-\texttt{k}}{2}}$. ∎
Appendix C Proofs of Results in Section˜4
This appendix provides proofs for the geometric and measure-approximation results in Section˜3. We first establish a uniform comparison between the volume of small geodesic balls on $\mathcal{S}(\theta)$ and Euclidean balls, and then use Wasserstein stability to control the superquantile functional under perturbations of the underlying measure.
C.1 Proof of Section˜4.1
We begin by showing that, for a k-dimensional ($\mathcal{C}^{2}$) embedded Riemannian manifold, the Riemannian volume of a sufficiently small geodesic ball is comparable to the k-dimensional Lebesgue measure of an Euclidean ball of the same radius. We already used this result to prove ˜4.2.
Proof.
Let $(\mathcal{S},\mathsf{g})$ be a k-dimensional compact Riemannian manifold. By compactness, $\mathcal{S}$ is geodesically complete. Assume:
  1. 1.
    $\operatorname{inj}(\mathcal{S})\geq r_{0}>0$ (uniform lower bound on the injectivity radius),
  2. 2.
    $|\mathrm{Sec}(x)|\leq\mathsf{C}_{0}$ for all $x\in\mathcal{S}$ (uniform bound on sectional curvature).
Step 1: Expressing the volume of a geodesic ball in normal coordinates. For a Riemannian manifold $(\mathcal{S},\mathsf{g})$, the volume measure $\mathcal{M}$ is locally given by the Riemannian volume form. In a coordinate chart $\{x^{i}\}$, one has
$\mathrm{d}\mathcal{M}(x)\;=\;\sqrt{\det\bigl(\mathsf{g}_{ij}(x)\bigr)}\,\mathrm{d}x^{1}\cdots\mathrm{d}x^{\texttt{k}},$
so that, for any measurable $U$,
$\int_{U}f(x)\,\mathrm{d}\mathcal{M}(x)\;=\;\int_{U}f(x)\,\sqrt{\det\bigl(\mathsf{g}_{ij}(x)\bigr)}\,\mathrm{d}x^{1}\cdots\mathrm{d}x^{\texttt{k}},$
where $\mathsf{g}_{ij}(x)$ is the coordinate matrix of the metric $\mathsf{g}$.
Fix a point $p\in\mathcal{S}$ and $0<r<r_{0}$. By [71, Theorem 5.4.2], the exponential map
$\exp_{p}:\mathbb{B}_{\texttt{k}}(0;r)\subset T_{p}\mathcal{S}\;\to\;\mathbb{B}_{\mathsf{g}}(p;r)\subset\mathcal{S}$
is a diffeomorphism, where we identify $T_{p}\mathcal{S}\cong\mathbb{R}^{\texttt{k}}$ via an orthonormal basis and $\mathbb{B}_{\texttt{k}}(0;r)$ is the Euclidean ball of radius $r$.
Using $\exp_{p}$ as a change of variables, we write the volume of the geodesic ball as
$\mathcal{M}\bigl(\mathbb{B}_{\mathsf{g}}(p;r)\bigr)\;=\;\int_{\mathbb{B}_{\mathsf{g}}(p;r)}\mathrm{d}\mathcal{M}(x)\;=\;\int_{\mathbb{B}_{\texttt{k}}(0;r)}\bigl(\exp_{p}^{*}\mathrm{d}\mathcal{M}\bigr)(v),$ (39)
where $v=(v^{1},\dots,v^{\texttt{k}})$ are coordinates on $T_{p}\mathcal{S}\cong\mathbb{R}^{\texttt{k}}$ and $\mathrm{d}v$ is the standard Lebesgue measure on $\mathbb{R}^{\texttt{k}}$.
In normal coordinates at $p$, the exponential map $\exp_{p}$ is precisely the coordinate map, so $\exp_{p}^{*}\mathrm{d}\mathcal{M}$ has the expression
$(\exp_{p}^{*}\mathrm{d}\mathcal{M})(v)\;=\;\sqrt{\det\bigl(\mathsf{g}_{ij}(v)\bigr)}\,\mathrm{d}v^{1}\cdots\mathrm{d}v^{\texttt{k}},$
where $\mathsf{g}_{ij}(v)$ is the metric tensor in these normal coordinates. Thus (39) becomes
$\mathcal{M}\bigl(\mathbb{B}_{\mathsf{g}}(p;r)\bigr)\;=\;\int_{\mathbb{B}_{\texttt{k}}(0;r)}\sqrt{\det\!\bigl(\mathsf{g}_{ij}(v)\bigr)}\;\mathrm{d}v.$ (40)
Step 2: Quadratic control of the metric in normal coordinates, uniformly in $p$. In a normal coordinate neighborhood of $p$, the metric admits the expansion (see, e.g., [71, Section 5.4])
$\mathsf{g}_{ij}(v)\;=\;\delta_{ij}-\frac{1}{3}R_{ikjl}(p)\,v^{k}v^{l}+\mathcal{O}(\|v\|^{3}),$ (41)
where $R_{ikjl}(p)$ are the components of the Riemann curvature tensor at $p$, and $\delta_{ij}$ is the Kronecker delta.
The sectional curvature bound $|\mathrm{Sec}|\leq\mathsf{C}_{0}$ implies that the curvature tensor is uniformly bounded on $\mathcal{S}$: there exists a constant $\widetilde{\mathsf{C}}>0$, depending only on $\mathsf{C}_{0}$ and k, such that
$|R_{ikjl}(x)|\leq\widetilde{\mathsf{C}}\quad\text{for all }x\in\mathcal{S}\text{ and all indices }i,k,j,l.$
(For example, this follows from norm equivalence on the finite-dimensional space of algebraic curvature tensors.) After possibly increasing the constant, we may (and do) assume $\widetilde{\mathsf{C}}=\mathsf{C}_{0}$ for notational simplicity, so that $|R_{ikjl}(x)|\;\leq\;\mathsf{C}_{0}$ for all $x\in\mathcal{S}$.
Since $\mathcal{S}$ is compact, higher-order derivatives of the metric are also bounded, and the $\mathcal{O}(\|v\|^{3})$ term in (41) can be dominated uniformly in $p$. Hence there exist constants $r_{1}\in(0,r_{0}]$ and $C>0$, depending only on $(\texttt{k},\mathsf{C}_{0},r_{0})$, such that for all $p\in\mathcal{S}$ and all $\|v\|\leq r_{1}$,
$\bigl|\mathsf{g}_{ij}(v)-\delta_{ij}\bigr|\leq C\,\mathsf{C}_{0}\,\|v\|^{2}\quad\text{for all }i,j.$ (42)
Let $\|\cdot\|_{\mathrm{op}}$ denote the operator norm on matrices with respect to the Euclidean norm. Using the elementary bound $\|\mathsf{g}(v)-I\|_{\mathrm{op}}\leq\texttt{k}\;\max_{i,j}\bigl|\mathsf{g}_{ij}(v)-\delta_{ij}\bigr|$, we obtain from (42) that
$\|\mathsf{g}(v)-I\|_{\mathrm{op}}\;\leq\;C^{\prime}\,\mathsf{C}_{0}\,\|v\|^{2},\qquad\|v\|\leq r_{1},$ (43)
for some constant $C^{\prime}>0$ depending only on $(\texttt{k},\mathsf{C}_{0},r_{0})$.
Step 3: Explicit small-radius condition in terms of $r_{0}$ and $\mathsf{C}_{0}$. Define
$r_{*}\;:=\;\min\Bigl\{\,r_{1},\,\frac{1}{2\sqrt{C^{\prime}\,\mathsf{C}_{0}}}\,\Bigr\}.$
Then for every $v$ with $\|v\|\leq r_{*}$, (43) yields $\|\mathsf{g}(v)-I\|_{\mathrm{op}}\leq C^{\prime}\,\mathsf{C}_{0}\,r_{*}^{2}\leq{1}/{4}$. Thus all eigenvalues of the symmetric matrix $\mathsf{g}(v)$ lie in the interval [3/4, 5/4][3/4,\,5/4]. In particular,
$\frac{3}{4}\,I\preceq\mathsf{g}(v)\preceq\frac{5}{4}\,I\,\Longrightarrow\,\Bigl(\frac{3}{4}\Bigr)^{\texttt{k}/2}\leq\sqrt{\det\bigl(\mathsf{g}_{ij}(v)\bigr)}\leq\Bigl(\frac{5}{4}\Bigr)^{\texttt{k}/2},\quad\text{for all }\|v\|\leq r_{*}.$ (44)
Observe that $r_{*}$ depends only on $(\texttt{k},\mathsf{C}_{0},r_{0})$. In particular, there exists a numerical constant $c_{\texttt{k}}>0$, such that
$r_{*}\;\geq\;\min\Bigl\{\,r_{0},\ \frac{c_{\texttt{k}}}{\sqrt{\mathsf{C}_{0}}}\,\Bigr\}.$
Thus, up to adjusting the constant, we may think of the admissible radius as
$0<r\;\leq\;\min\Bigl\{\,r_{0},\ \frac{c_{\texttt{k}}}{\sqrt{\mathsf{C}_{0}}}\,\Bigr\},$
which makes the dependence on the injectivity radius and sectional curvature bound explicit.
Step 4: Volume comparison. Combining (40) and (44), for any $0<r\leq r_{*}$ we get
$\Bigl(\frac{3}{4}\Bigr)^{\texttt{k}/2}{\texttt{Leb}}_{\texttt{k}}\bigl(\mathbb{B}_{\texttt{k}}(p;r)\bigr)\leq\mathcal{M}\bigl(\mathbb{B}_{\mathsf{g}}(p;r)\bigr)$
$=\int_{\mathbb{B}_{\texttt{k}}(0;r)}\sqrt{\det\bigl(\mathsf{g}_{ij}(v)\bigr)}\,\mathrm{d}v\leq\Bigl(\frac{5}{4}\Bigr)^{\texttt{k}/2}{\texttt{Leb}}_{\texttt{k}}\bigl(\mathbb{B}_{\texttt{k}}(p;r)\bigr).$
where ${\texttt{Leb}}_{\texttt{k}}$ denotes Lebesgue measure on $\mathbb{R}^{\texttt{k}}$. This is the desired comparison, with constants $\mathsf{c}_{L}:=({3}/{4})^{\texttt{k}/2}$ and $\mathsf{c}_{H}:=({5}/{4})^{\texttt{k}/2}$ valid for all radii
$0<r\leq r_{*}\;\text{ and hence for all }\;0<r\leq\min\Bigl\{\,r_{0},\ \frac{c_{\texttt{k}}}{\sqrt{\mathsf{C}_{0}}}\,\Bigr\}.$
C.2 Proof of Section˜4.2
We next provide the proof of Section˜4.2 which is essential for our approximation result in ˜4.5.
Proof.
Recall
$\phi_{\lambda,\delta}(\theta,\beta)=\beta+\frac{1}{\delta}\,\operatorname{E}_{X\sim\mu_{\theta}^{\lambda}}\!\bigl[(f(\theta,X)-\beta)_{+}\bigr].$
By Section˜2, $f(\theta,\cdot)$ is $L_{f,1}$-Lipschitz. Since $t\mapsto(t-\beta)_{+}$ is $1$-Lipschitz, the map
$G_{\beta}(x):=(f(\theta,x)-\beta)_{+}$
is $L_{f,1}$-Lipschitz in $x$. Hence, by the Kantorovich–Rubinstein duality for $\mathbb{W}_{1}$,
$\bigl|\phi_{\lambda,\delta}(\theta,\beta)-\phi_{0,\delta}(\theta,\beta)\bigr|$ $=\frac{1}{\delta}\,\Bigl|\operatorname{E}_{\mu_{\theta}^{\lambda}}\!\bigl[G_{\beta}(X)\bigr]-\operatorname{E}_{\mu_{\theta}^{0}}\!\bigl[G_{\beta}(X)\bigr]\Bigr|\leq\frac{L_{f,1}}{\delta}\,\mathbb{W}_{1}\!\bigl(\mu_{\theta}^{\lambda},\mu_{\theta}^{0}\bigr).$ (45)
Finally, to prove the statement of Lemma 4.2, we have
$\min_{\beta^{\prime}}\phi_{0,\delta}(\theta,\beta^{\prime})\leq\phi_{0,\delta}(\theta,\beta)$
$\leq\phi_{\lambda,\delta}(\theta,\beta)+|\phi_{0,\delta}(\theta,\beta)-\phi_{\lambda,\delta}(\theta,\beta)|$
$\overset{(a)}{\leq}\phi_{\lambda,\delta}(\theta,\beta)+\frac{L_{f,1}}{\delta}\cdot\mathbb{W}_{1}(\mu^{\lambda}_{g}(\theta),\mu^{0}_{g}(\theta))),$ (46)
where (a) follows from (45). Then, taking the minimum over $\beta$ at the r.h.s. of (46), we get $\min_{\beta^{\prime}}\phi_{0,\delta}(\theta,\beta^{\prime})-\min_{\beta}\phi_{\lambda,\delta}(\theta,\beta)\leq\frac{L_{f,1}}{\delta}(\mu^{\lambda}_{g}(\theta),\mu^{0}_{g}(\theta)))$. Interchanging $\min_{\beta}\phi_{\lambda,\delta}(\theta,\beta)$ and $\min_{\beta^{\prime}}\phi_{0,\delta}(\theta,\beta^{\prime})$, one obtains
$|\min_{\beta}\phi_{\lambda,\delta}(\theta,\beta)-\min_{\beta^{\prime}}\phi_{0,\delta}(\theta,\beta^{\prime})|\leq L_{f,1}\delta^{-1}\mathbb{W}_{1}(\mu^{\lambda}_{g}(\theta),\mu^{0}_{g}(\theta))).$ (47)
Appendix D Proofs of Results in Section˜5
This appendix collects technical results for the oracle analysis in Section˜5. We begin by deriving moment and tail bounds for the Gibbs measure; these bounds are then used to prove Section˜5.1. Next, we establish accuracy guarantees for the PSGD-based inner estimator. Finally, we present the complete proof of ˜5.5, separating the standard zeroth-order argument from our new control of the minima-selection error.
D.1 Proof of Section˜5.1
We start by presenting the following technical lemma, which is an essential prerequisite for the proof of Section˜5.1. It establishes a technical result ensuring the finiteness of the moments of the Gibbs measure $\mu_{\theta}^{\lambda}$.
Lemma D.1.
Under Section˜2, for
$\lambda\leq\min\left\{\frac{(2\pi)^{d/2}\mathsf{M}_{g}^{-\frac{d}{2}}}{2c},\frac{2\mathsf{D}^{2}}{n+d}\right\},$
we have for every $n\in\mathbb{N}\cup\{0\}$
$\mathbb{E}_{X\sim\mu_{\theta}^{\lambda}}[\|X\|^{n}\cdot\bm{1}\left[X\notin\mathbb{B}_{d}(0;2\mathsf{D})\right]]\leq\mathsf{W}_{n}\cdot\lambda^{-\frac{d}{2}+1}e^{-\frac{\mathsf{D}^{2}}{\lambda}}.$
Here,
$\mathsf{W}_{n}:=\mathsf{M}_{g}^{\frac{d}{2}}\cdot 2^{-d/2+2}\pi^{d/2+n-1}\cdot 2\mu_{\mathsf{qg}}^{-1}\cdot\max\{(d+n)/2,1\}\cdot\mathsf{D}^{d+n-1}$
with $\mathsf{M}_{g}$ is defined in Section˜3.2 and $c>0$ is a constant.
Proof.
Assume without loss of generality that $g$ does not depend on $\theta$, $\min_{x}g(x)=0$, and $0\in\mathcal{S}=\operatorname*{arg\,min}_{x}g(x)$. We first note that by switching to spherical coordinates, one obtains for any $n\geq 0$ the bound
$\int_{x:\|x\|>R}\|x\|^{n}e^{-c\|x\|^{2}}\mathrm{d}x=$
$\int_{0}^{\pi}\cdots\int_{0}^{\pi}\int_{0}^{2\pi}\int_{r>R}e^{-cr^{2}}r^{d+n-1}\sin^{d-2}(\phi_{1})\sin^{d-3}(\phi_{2})\ldots\sin(\phi_{d-2})\mathrm{d}r\mathrm{d}\phi_{1}\cdots\mathrm{d}\phi_{d-1}$
$\leq\int_{0}^{\pi}\cdots\int_{0}^{\pi}\int_{0}^{2\pi}\int_{r>R}e^{-cr^{2}}r^{d+n-1}\mathrm{d}r\mathrm{d}\phi_{1}\cdots\mathrm{d}\phi_{d-1}$
$\leq 2\pi^{d-1}\int_{R}^{+\infty}e^{-cr^{2}}r^{d+n-1}\mathrm{d}r\leq 2\pi^{d+n-1}c^{-\frac{d+n}{2}}\cdot\Gamma\left(\frac{d+n}{2},cR^{2}\right),$ (48)
where in the last inequality we use the definition of the upper incomplete Gamma function, i.e., $\Gamma(s,x):=\int_{x}^{+\infty}t^{s-1}e^{-t}\mathrm{d}t$. By Gautschi’s inequality [^26] By Gautschi-type’s inequality, $\Gamma(x,s)\leq sx^{s-1}e^{-x}$ for $x>s$ and $s\geq 1$. Note that for $s=1$, $\Gamma(s,x)=e^{-x}$., for $cR^{2}>(d+n)/2$ and $d+n\geq 2$,
$\int_{x:\|x\|>R}\|x\|^{n}e^{-c\|x\|^{2}}\mathrm{d}x\leq 2\pi^{d+n-1}\cdot c^{-\frac{d+n}{2}}\max\{(d+n)/2,1\}(cR^{2})^{\frac{d+n}{2}-1}e^{-cR^{2}}$
$\leq 2\pi^{d+n-1}\cdot c^{-1}\max\{(d+n)/2,1\}R^{d+n-2}e^{-cR^{2}}.$ (49)
Since
$\mathbb{E}_{X\sim\mu_{\theta}^{\lambda}}[\|X\|^{n}\cdot\bm{1}\left[X\notin\mathbb{B}_{d}(0;2\mathsf{D})\right]]=\frac{1}{Z_{\lambda}}\int_{\|x\|\geq 2\mathsf{D}}\|x\|^{n}e^{-g(x)/\lambda}\mathrm{d}x,$
and noting that, if $\mathcal{S}\subseteq\mathbb{B}_{d}(0;\mathsf{D})$, then for $x\notin\mathbb{B}_{d}(0;2\mathsf{D})$ we have
$\|x\|-\mathsf{D}\leq\mathrm{dist}(x,\mathcal{S}),$
the quadratic growth of $g$ outside $\mathbb{B}_{d}(0;2\mathsf{D})$ implies
$\frac{\mu_{\mathsf{qg}}}{2}(\|x\|-\mathsf{D})^{2}\leq g(x),\quad\forall\,\|x\|\geq 2\mathsf{D}.$ (50)
Hence,
$\mathbb{E}_{X\sim\mu_{\theta}^{\lambda}}[\|X\|^{n}\cdot\bm{1}\left[X\notin\mathbb{B}_{d}(0;2\mathsf{D})\right]]\leq\frac{1}{Z_{\lambda}}\int_{\|x\|\geq 2\mathsf{D}}\|x\|^{n}e^{-\frac{\mu_{\mathsf{qg}}(\|x\|-\mathsf{D})^{2}}{2\lambda}}\mathrm{d}x.$
Applying (49) gives for $\lambda<2\mathsf{D}^{2}/(n+d)$,
$\mathbb{E}_{X\sim\mu_{\theta}^{\lambda}}[\|X\|^{n}\cdot\bm{1}\left[X\notin\mathbb{B}_{d}(0;2\mathsf{D})\right]]\leq$
$\frac{1}{Z_{\lambda}}\cdot 2\pi^{d+n-1}\cdot\left(\frac{2\lambda}{\mu_{\mathsf{qg}}}\right)\cdot\max\{(d+n)/2,1\}\cdot\mathsf{D}^{d+n-2}e^{-\frac{\mathsf{D}^{2}}{\lambda}}.$ (51)
For a lower bound on $Z_{\lambda}$, we use the estimate (see [38, Remark A.21]): for a compact set $U\subset\mathbb{R}^{d}$ and $x_{0}\in\mathcal{S}$,
$\int_{U}e^{-g(x)/\lambda}\mathrm{d}x\geq\lambda^{\frac{d}{2}}\left[(2\pi)^{d/2}\det\left(\frac{\partial^{2}g(x_{0})}{\partial x^{2}}\right)^{-\frac{1}{2}}-c\lambda\right].$
Since $\det\left(\frac{\partial^{2}g(x_{0})}{\partial x^{2}}\right)\leq\mathsf{M}_{g}^{d}$ (by Section˜3.2), it follows that
$Z_{\lambda}\geq\lambda^{\frac{d}{2}}\left[(2\pi)^{d/2}\mathsf{M}_{g}^{-\frac{d}{2}}-c\lambda\right].$
Thus, if
$\lambda\leq\min\left\{\frac{(2\pi)^{d/2}\mathsf{M}_{g}^{-\frac{d}{2}}}{2c},\frac{2\mathsf{D}^{2}}{n+d}\right\},$
we have
$Z_{\lambda}\geq\lambda^{\frac{d}{2}}\cdot(2\pi)^{d/2}\mathsf{M}_{g}^{-\frac{d}{2}}/2.$
Substituting this into (51) completes the proof. ∎
We then prove Section˜5.1 which states that every global minimizer of $\phi_{\lambda,\delta}(\theta,\cdot)$ lies in a bounded interval.
Proof.
Upper bound on $\beta^{*}(\theta)$: Pick a $\beta^{*}(\theta)\in\operatorname*{arg\,min}_{\beta}\phi_{\lambda,\delta}(\theta,\beta)$. It follows that
$\beta^{*}(\theta)\leq\beta^{*}(\theta)+\delta^{-1}\mathbb{E}[\max\{0,f(\theta,X)-\beta^{*}(\theta)\}]$
$=\min_{\beta}\phi_{\lambda,\delta}(\theta,\beta)$ (52)
In the following, we give an upper bound on $\min_{\beta}\phi_{\lambda,\delta}(\theta,\beta)$. First, note that similar to the proof of Section˜4.2, we have
$|\min_{\beta}\phi_{0,\delta}(\theta,\beta)-\min_{\beta}\phi_{\lambda,\delta}(\theta,\beta)|\leq\frac{L_{f,1}}{\delta}\cdot\mathbb{W}_{1}(\mu_{\theta}^{\lambda},\mu_{\theta}^{0}).$
Using this bound along with the result of Section˜2.1.2, we obtain
$\beta^{*}(\theta)\leq\min_{\beta}\phi_{\lambda,\delta}(\theta,\beta)\leq\min_{\beta}\phi_{0,\delta}(\theta,\beta)+|\min_{\beta}\phi_{0,\delta}(\theta,\beta)-\min_{\beta}\phi_{\lambda,\delta}(\theta,\beta)|$
$\leq\min_{\beta}\phi_{0,\delta}(\theta,\beta)+\frac{L_{f,1}}{\delta}\cdot K_{0}\lambda^{\frac{1}{2}}.$ (53)
Then we have for $Y\sim\mu_{\theta}^{0}$,
$\min_{\beta}\phi_{0,\delta}(\theta,\beta)=\text{SQ}_{1-\delta}(f(\theta,Y))\leq\max_{x\in\mathcal{S}(\theta)}f(\theta,x),$
where the last inequality comes from the definition of superquantile and the fact that $\text{supp}(\mu_{\theta}^{0})=\mathcal{S}(\theta)$. By Equation˜5, $\max_{x\in\mathcal{S}(\theta)}f(\theta,x)\leq L_{f,0}(\mathsf{D})$. Hence, we have $\min_{\beta}\phi_{0,\delta}(\theta,\beta)\leq L_{f,0}(\mathsf{D})$ and then from (53), we obtain
$\beta^{*}(\theta)\leq L_{f,0}(\mathsf{D})+\frac{L_{f,1}}{\delta}\cdot K_{0}\lambda^{1/2}$ (54)
Lower bound on $\beta^{*}(\theta)$: By the definition of the $(1-\delta)$-quantile of the random variable $f(\theta,X)$ under $X\sim\mu_{\theta}^{\lambda}$, it holds that
$\mathbb{P}_{X\sim\mu_{\theta}^{\lambda}}[f(\theta,X)<\beta^{*}(\theta)]>1-\delta.$
Because $\delta\leq 1/2$, we in particular have
$\mathbb{P}_{X\sim\mu_{\theta}^{\lambda}}[f(\theta,X)<\beta^{*}(\theta)]>\frac{1}{2}.$
W.L.O.G., we assume that $\mathcal{S}(\theta)\subseteq\mathbb{B}_{d}(0;\mathsf{D})$ with sufficiently large $\mathsf{D}$. Then
$\frac{1}{2}<\mathbb{P}_{X\sim\mu_{\theta}^{\lambda}}[f(\theta,X)<\beta^{*}(\theta)]\leq$
$\mathbb{P}[f(\theta,X)<\beta^{*}(\theta),X\in\mathbb{B}_{d}(0;2\mathsf{D})]+\mathbb{P}[f(\theta,X)<\beta^{*}(\theta),X\notin\mathbb{B}_{d}(0;2\mathsf{D})]$
$\leq\mathbb{P}[f(\theta,X)<\beta^{*}(\theta),X\in\mathbb{B}_{d}(0;2\mathsf{D})]+\mathbb{P}[X\notin\mathbb{B}_{d}(0;2\mathsf{D})],$
Next, we use the result in Section˜D.1 for $n=0$, to give the following bound:
$\mathbb{P}_{X\sim\mu_{\theta}^{\lambda}}[X\notin\mathbb{B}_{d}(0;2\mathsf{D})]\leq\mathsf{W}_{0}\cdot\lambda^{-\frac{d}{2}+1}e^{-\frac{\mathsf{D}^{2}}{\lambda}}.$
Then
$\mathbb{P}[f(\theta,X)<\beta^{*}(\theta),X\in\mathbb{B}_{d}(0;2\mathsf{D})]\geq$
$\frac{1}{2}-\mathsf{W}_{0}\cdot\lambda^{-\frac{d}{2}+1}e^{-\frac{\mathsf{D}^{2}}{\lambda}}.$ (55)
For the case $d\leq 2$, when $\lambda\leq\frac{\mathsf{D}^{2}}{\log(4\mathsf{W}_{0})}$, we have $\mathbb{P}[f(\theta,X)<\beta^{*}(\theta),X\in\mathbb{B}_{d}(0;2\mathsf{D})]\geq\frac{1}{4}$. For the case $d>2$, when
$\lambda\leq\frac{2\mathsf{D}^{2}}{(d-2)\cdot W_{\mathrm{Lam}}\left(\frac{2}{d-2}\left(\frac{4\mathsf{W}_{0}}{\mathsf{D}^{\,d-2}}\right)^{\frac{2}{d-2}}\right)},$
where $W_{\mathrm{Lam}}$ is the Lambert function, we have
$\mathbb{P}[f(\theta,X)<\beta^{*}(\theta),X\in\mathbb{B}_{d}(0;2\mathsf{D})]\geq\frac{1}{4}.$
Thus, there exists some $x_{0}\in\mathbb{B}_{d}(0;2\mathsf{D})$ such that $f(\theta,x_{0})\leq\beta^{*}(\theta)$. We conclude from Equation˜5, $L_{f,0}(2\mathsf{D})\leq\min_{x\in\mathbb{B}_{d}(0;\mathsf{D})}f(\theta,x)\leq\beta^{*}(\theta)$. This completes the proof. ∎
D.2 Proof of Section˜5.1
This subsection proves the mean-square accuracy bound for the PSGD estimator of $\tilde{F}_{{\texttt{SQ-G}}}(\theta)$. This estimate is the key input in the zeroth-order convergence proof.
For completeness, we state the following bounds, which are used in the convergence analysis of PSGD (see Algorithm˜1). The function $\phi_{\lambda,\delta}(\theta,\beta)$ is $(1+\delta^{-1})$-Lipschitz on $\beta$ since, given $\theta\in\Theta$, for every $\beta_{1},\beta_{2}\in\mathbb{R}$,
$|\phi_{\lambda,\delta}(\theta,\beta_{1})-\phi_{\lambda,\delta}(\theta,\beta_{2})|\leq|\beta_{1}-\beta_{2}|+\frac{1}{\delta}\cdot\mathbb{E}[\left|h(f(\theta,X)-\beta_{1})-h(f(\theta,X)-\beta_{2})\right|]$
$\leq\left[1+\frac{1}{\delta}\right]\cdot|\beta_{1}-\beta_{2}|.$ (56)
Here the last inequality comes from $h(t):=\max\{t,0\}$ is 1-Lipschitz. Moreover,
$|\tilde{\partial}_{\beta}\phi_{\lambda,\delta}(\theta,\beta;X)|=\left|1+\frac{1}{\delta}\hat{\partial}_{\beta}[\max\{f(\theta,X)-\beta,0\}]\right|^{2}\leq|1+\delta^{-1}|:=\sigma.$
Here $\hat{\partial}_{\beta}[\max\{f(\theta,X)-\beta,0\}]$ is the sub-gradient of $\max\{f(\theta,X)-\beta,0\}$ and is bounded uniformly by $1$. We then provide the proof of Section˜5.1 in the following:
Proof.
To give an upper bound on $\mathbb{E}[|\tilde{\psi}(\theta)-\tilde{F}_{{\texttt{SQ-G}}}(\theta)|^{2}\mid\theta]$, by the triangle inequality, we have
$\mathbb{E}[|\phi_{\lambda,\delta}(\theta,\hat{\beta};\tilde{X}_{I})-\tilde{F}_{\texttt{SQ-G}}(\theta)|^{2}\mid\theta]\leq$
$2\mathbb{E}[|\phi_{\lambda,\delta}(\theta,\hat{\beta};\tilde{X}_{I})-\phi_{\lambda,\delta}(\theta,\hat{\beta})|^{2}\mid\theta]+2\mathbb{E}[|\phi_{\lambda,\delta}(\theta,\hat{\beta})-\tilde{F}_{\texttt{SQ-G}}(\theta)|^{2}\mid\theta]$
$=2\underbrace{\mathbb{E}[|\phi_{\lambda,\delta}(\theta,\hat{\beta};\tilde{X}_{I})-\phi_{\lambda,\delta}(\theta,\hat{\beta})|^{2}\mid\theta]}_{(i)}+2\underbrace{\mathbb{E}[(\phi_{\lambda,\delta}(\theta,\hat{\beta})-\min_{\beta}\phi_{\lambda,\delta}(\theta,\beta))^{2}\mid\theta]}_{(ii)},$ (57)
where the last equation follows from $\tilde{F}_{\texttt{SQ-G}}(\theta)=\min_{\beta}\phi_{\lambda,\delta}(\theta,\beta)$. We will give the upper bounds on the terms (i) and (ii) in the following.
Upper bound on (i) in (57): We use the following chain of inequalities:
[…]
where the last inequality comes from the fact that $\{h(f(\theta,\tilde{X}^{(i)})-\hat{\beta})\}_{i\in I}$ are independent random variables given $\theta$ and $\hat{\beta}$.
Now we give a constant upper bound on (58) in the following:
$\mathbb{E}\left[\left(h(f(\theta,\tilde{X}^{(i)})-\hat{\beta})-\mathbb{E}[h(f(\theta,X)-\hat{\beta})\mid\theta,\hat{\beta}]\right)^{2}\mid\theta,\hat{\beta}\right$]
$\leq\mathbb{E}\left[\left(h(f(\theta,\tilde{X}^{(i)})-\hat{\beta})-h(f(\theta,X)-\hat{\beta})\right)^{2}\mid\theta,\hat{\beta}\right$]
$\leq\mathbb{E}\left[\left(f(\theta,\tilde{X}^{(i)})-f(\theta,X)\right)^{2}\mid\theta,\hat{\beta}\right$]
$\leq 4\mathbb{E}\left[(f(\theta,X))^{2}\mid\theta,\hat{\beta}\right$]
$\overset{(a)}{\leq}4\mathbb{E}\left[(C_{f}\|X\|^{n_{f}}+D_{f})^{2}\mid\theta,\hat{\beta}\right$]
$\leq 8C_{f}^{2}\mathbb{E}[\|X\|^{2n_{f}}]+8D_{f}^{2}$
$\overset{(b)}{\leq}\tilde{C}_{f}:=8C_{f}^{2}\cdot\left((2\mathsf{D})^{2n_{f}}+\mathsf{W}_{2n_{f}}\cdot\lambda^{-\frac{d}{2}}e^{-\frac{\mathsf{D}^{2}}{\lambda}}\right)+8D_{f}^{2},$ (59)
where step (a) follows from Section˜2, and step (b) uses Section˜D.1.
Then from (58), we get
$\mathbb{E}[|\phi_{\lambda,\delta}(\theta,\hat{\beta};\tilde{X}_{I})-\phi_{\lambda,\delta}(\theta,\hat{\beta})|^{2}\mid\theta]\leq\frac{\tilde{C}_{f}}{\delta^{2}|I|}.$ (60)
Upper bound on (ii) in (57): Because $\phi_{\lambda,\delta}(\theta,\cdot)$ is $\sigma$-Lipschitz in $\beta$ and the feasible interval for $\beta$ is $\mathrm{B},\mathrm{B}]$, we have the deterministic range bound
$0\leq\phi_{\lambda,\delta}(\theta,\hat{\beta})-\min_{\beta}\phi_{\lambda,\delta}(\theta,\beta)\leq R:=2\sigma\,\mathrm{B}.$
Given that (i) $\phi_{\lambda,\delta}(\theta,\cdot)$ is convex and Lipschitz, (ii) the feasible set is bounded, and (iii) the stochastic subgradient $\tilde{\partial}_{\beta}\phi_{\lambda,\delta}(\theta,\beta;X)$ satisfies $|\tilde{\partial}_{\beta}\phi_{\lambda,\delta}(\theta,\beta;X)|\leq\sigma:=1+\delta^{-1}$, we invoke the convergence guarantees for convex Lipschitz objectives [67, 51]. Accordingly, the output $\hat{\beta}$ of PSGD (see Algorithm˜1) satisfies
$\mathbb{P}\!\left(\phi_{\lambda,\delta}(\theta,\hat{\beta})-\min_{\beta}\phi_{\lambda,\delta}(\theta,\beta)\leq\epsilon_{v}\;\middle|\;\theta\right)\;\geq\;1-\Delta.$
after at most $\mathcal{O}(\mathrm{B}^{2}\sigma^{2}\epsilon_{v}^{-2}\log^{2}(\Delta^{-1}))$ iterations.
From $0\leq\phi_{\lambda,\delta}(\theta,\hat{\beta})-\min_{\beta}\phi_{\lambda,\delta}(\theta,\beta)\leq R$ and the tail bound we get
$\mathbb{E}\!\left[\left(\phi_{\lambda,\delta}(\theta,\hat{\beta})-\min_{\beta}\phi_{\lambda,\delta}(\theta,\beta)\right)^{2}\,\middle|\;\theta\right]\;\leq\;\epsilon_{v}^{2}(1-\Delta)+R^{2}\Delta\;\leq\;\epsilon_{v}^{2}+R^{2}\Delta.$
Choosing $\Delta:=\min\{1,\,\epsilon_{v}^{2}/R^{2}\}$ yields
$\mathbb{E}\left[\left(\phi_{\lambda,\delta}(\theta,\hat{\beta})-\min_{\beta}\phi_{\lambda,\delta}(\theta,\beta)\right)^{2}\,\middle|\;\theta\right]\;\leq\;2\,\epsilon_{v}^{2},$ (61)
at the cost of
$L\;\geq\;C\cdot\frac{\mathrm{B}^{2}}{\delta^{2}\epsilon_{v}^{2}}\cdot\log^{2}\left(\frac{1}{\epsilon_{v}\delta}\right).$
Thus, by substituting (61) and (60) into (57), we obtain
$\mathbb{E}[|\tilde{\psi}(\theta)-\tilde{F}_{{\texttt{SQ-G}}}(\theta)|^{2}\mid\theta]\leq 2\epsilon_{v}^{2}+\frac{2\tilde{C}_{f}}{\delta^{2}|I|}.$ (62)
By picking $\delta=\epsilon_{v}^{\texttt{k}}$ and $|I|\geq 4\tilde{C}_{f}\epsilon_{v}^{-2\texttt{k}-2}$ and
$L\;\geq\;\mathcal{O}\left(\texttt{k}^{2}\cdot\frac{\mathrm{B}^{2}}{\epsilon_{v}^{2(\texttt{k}+1)}}\cdot\log^{2}\left(\frac{1}{\epsilon_{v}}\right)\right),$
we get $\mathbb{E}[|\tilde{\psi}(\theta)-\tilde{F}_{{\texttt{SQ-G}}}(\theta)|^{2}\mid\theta]\leq\epsilon_{v}^{2}.$
D.3 Proof of ˜5.5
We now prove ˜5.5. The argument follows the standard projected zeroth-order analysis for the smoothed objective, and the only additional step is to bound the error introduced by replacing $F_{\max}$ with $\tilde{\psi}$ using Equation˜26 and Section˜5.1.
Proof.
The proof has two parts. We first follow the standard convergence analysis for projected zeroth-order methods applied to the smoothed objective of $F_{\max}$. The only nonstandard step is then to control the additional error induced by replacing $F_{\max}$ with $\tilde{\psi}$ in the gradient estimator, where we invoke Section˜5.1.
Step 1: Standard zeroth-order analysis for the smoothed problem. We define the smoothed objective $\psi_{\rho}(\theta):=\mathbb{E}_{u\sim\text{uniform}(\mathbb{S}^{m-1})}[F_{\max}(\theta+\rho u)]$, where $u$ is uniformly distributed on the unit sphere $\mathbb{S}^{m-1}$. As shown in [68, Theorem 1 and Lemma 2], for any $L_{\mathsf{f},1}$-Lipschitz function $\mathsf{f}:\mathbb{R}^{m}\to\mathbb{R}$, the smoothed version $\mathsf{f}_{\rho}(\theta):=\mathbb{E}_{u\sim\text{uniform}(\mathbb{S}^{m-1})}[\mathsf{f}(\theta+\rho u)]$ satisfies:
  • $\bullet$
    $\mathsf{f}_{\rho}$ is a $\rho\sqrt{m}L_{\mathsf{f},1}$-approximation of $\mathsf{f}$, $\mathsf{f}_{\rho}$ is a $L_{f,1}$-Lipschitz function,
  • $\bullet$
    $\mathsf{f}_{\rho}$ is differentiable with gradient Lipschitz constant $L_{\mathsf{f},2}=\sqrt{m}L_{\mathsf{f},1}\rho^{-1}$,
  • $\bullet$
    $\nabla\mathsf{f}_{\rho}(\theta)\in\partial_{\rho}\mathsf{f}(\theta)$, the $\rho$-subdifferential of $\mathsf{f}$, as shown in [57, Theorem 3.1].
In our scenario, by using Section˜3.1, $F_{\max}$ is Lipschitz continuous with the constant $L_{F_{\max},1}:=L_{f,1}(1+L_{g,2}/\mu)$. By applying the aforementioned findings on $F_{\max}$, it follows that $\psi_{\rho}$ is $(L_{\psi,2}=\sqrt{m}L_{F_{\max},1}\rho^{-1})$-smooth. Furthermore, the gradient $\nabla\psi_{\rho}(\theta)$ is contained in the $\rho$-subdifferential set of $F_{\max}$ at the point $\theta$. Therefore, if a point $\theta$ satisfies $\|\mathcal{G}_{\Theta}(\theta,\nabla\psi_{\rho};\eta)\|\leq\epsilon$, then $\theta$ is an $(\epsilon,\rho,\eta)$-generalized Goldstein stationary point of $F_{\max}$ (see Section˜5.2). In what follows, we analyze the convergence of Algorithm 2 to a critical point of $\psi_{\rho}$ in terms of the gradient mapping.
Given the $(L_{\psi,2}=\sqrt{m}L_{F_{\max},1}\rho^{-1})$-smoothness of $\psi_{\rho}(\theta)$, we have
$\psi_{\rho}(\theta_{i+1})$ $\leq\psi_{\rho}(\theta_{i})+\langle\nabla\psi_{\rho}(\theta_{i}),\theta_{i+1}-\theta_{i}\rangle+\frac{L_{\psi,2}}{2}\|\theta_{i+1}-\theta_{i}\|^{2}$ (63)
Substituting $\theta_{i+1}-\theta_{i}=-\tilde{\eta}_{i}\mathcal{G}_{\Theta}(\theta_{i},\tilde{\mathsf{g}}_{\rho}(\theta_{i});\tilde{\eta}_{i})$, we have
$\psi_{\rho}(\theta_{i+1})\leq\psi_{\rho}(\theta_{i})-\tilde{\eta}_{i}\langle\nabla\psi_{\rho}(\theta_{i}),\mathcal{G}_{\Theta}(\theta_{i},\tilde{\mathsf{g}}_{\rho}(\theta_{i});\tilde{\eta}_{i})\rangle+\frac{L_{\psi,2}}{2}\tilde{\eta}_{i}^{2}\|\mathcal{G}_{\Theta}(\theta_{i},\tilde{\mathsf{g}}_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}$ (64)
Expanding the inner product and rearranging terms
$\psi_{\rho}(\theta_{i+1})$ $\leq\psi_{\rho}(\theta_{i})-\tilde{\eta}_{i}\langle\tilde{\mathsf{g}}_{\rho}(\theta_{i}),\mathcal{G}_{\Theta}(\theta_{i},\tilde{\mathsf{g}}_{\rho}(\theta_{i});\tilde{\eta}_{i})\rangle+\frac{L_{\psi,2}}{2}\tilde{\eta}_{i}^{2}\|\mathcal{G}_{\Theta}(\theta_{i},\tilde{\mathsf{g}}_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}$
$\quad-\tilde{\eta}_{i}\langle\nabla\psi_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i}),\mathcal{G}_{\Theta}(\theta_{i},\tilde{\mathsf{g}}_{\rho}(\theta_{i});\tilde{\eta}_{i})\rangle$ (65)
Further simplifying and using the fact that $\|\mathcal{G}_{\Theta}(\theta,v;\eta)\|^{2}\leq\langle\mathcal{G}_{\Theta}(\theta,v;\eta),v\rangle$ for every $v\in\Theta$, we have
$\psi_{\rho}(\theta_{i+1})\leq\psi_{\rho}(\theta_{i})-\left[\tilde{\eta}_{i}-\frac{L_{\psi,2}}{2}\tilde{\eta}_{i}^{2}\right]\cdot\|\mathcal{G}_{\Theta}(\theta_{i},\tilde{\mathsf{g}}_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}$
$-\tilde{\eta}_{i}\langle\nabla\psi_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i}),\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\rangle$
$\quad-\tilde{\eta}_{i}\langle\nabla\psi_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i}),\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})-\mathcal{G}_{\Theta}(\theta_{i},\tilde{\mathsf{g}}_{\rho}(\theta_{i});\tilde{\eta}_{i})\rangle$
$\leq\psi_{\rho}(\theta_{i})-\left[\tilde{\eta}_{i}-\frac{L_{\psi,2}}{2}\tilde{\eta}_{i}^{2}\right]\cdot\|\mathcal{G}_{\Theta}(\theta_{i},\tilde{\mathsf{g}}_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}$
$-\tilde{\eta}_{i}\langle\nabla\psi_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i}),\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\rangle+\tilde{\eta}_{i}\|\nabla\psi_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i})\|^{2},$ (66)
where the last inequality comes from the Cauchy-Schwartz inequality and $\|\mathcal{G}_{\Theta}(\theta,v_{1};\eta)-\mathcal{G}_{\Theta}(\theta,v_{2};\eta)\|\leq\|v_{1}-v_{2}\|$. Let us define $\mathsf{g}_{\rho}(\theta):=\frac{1}{b_{u}}\sum_{t=1}^{b_{u}}\mathsf{g}_{\rho}(\theta,u_{t})$ where $\mathsf{g}_{\rho}(\theta,u_{t}):=\frac{m}{2\rho}\cdot[F_{\max}(\theta+\rho u_{t})-F_{\max}(\theta-\rho u_{t})]u_{t}$. Then
$\psi_{\rho}(\theta_{i+1})\leq\psi_{\rho}(\theta_{i})-\left[\tilde{\eta}_{i}-\frac{L_{\psi,2}}{2}\tilde{\eta}_{i}^{2}\right]\cdot\|\mathcal{G}_{\Theta}(\theta_{i},\tilde{\mathsf{g}}_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}$
$-\tilde{\eta}_{i}\langle\nabla\psi_{\rho}(\theta_{i})-\mathsf{g}_{\rho}(\theta_{i}),\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\rangle$
$-\tilde{\eta}_{i}\langle\mathsf{g}_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i}),\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\rangle+\tilde{\eta}_{i}\|\nabla\psi_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i})\|^{2},$ (67)
By picking $\tilde{\eta}_{i}\leq 1/L_{\psi,2}$, we have
$\|\mathcal{G}_{\Theta}(\theta_{i},\tilde{\mathsf{g}}_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}\leq\frac{2[\psi_{\rho}(\theta_{i})-\psi_{\rho}(\theta_{i+1})]}{\tilde{\eta}_{i}}$
$-2\langle\nabla\psi_{\rho}(\theta_{i})-\mathsf{g}_{\rho}(\theta_{i}),\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\rangle$
$-2\langle\mathsf{g}_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i}),\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\rangle+2\|\nabla\psi_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i})\|^{2}.$ (68)
Then by using Young’s inequality ($\langle{\bf a},{\bf b}\rangle\leq c\|{\bf a}\|^{2}/2+\|{\bf b}\|^{2}/(2c)$ for the vectors ${\bf a},{\bf b}$ and positive constant $c$, we have
$\|\mathcal{G}_{\Theta}(\theta_{i},\tilde{\mathsf{g}}_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}\leq\frac{2[\psi_{\rho}(\theta_{i})-\psi_{\rho}(\theta_{i+1})]}{\tilde{\eta}_{i}}$
$-2\langle\nabla\psi_{\rho}(\theta_{i})-\mathsf{g}_{\rho}(\theta_{i}),\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\rangle+4\|\mathsf{g}_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i})\|^{2}$
$+4^{-1}\|\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}+2\|\nabla\psi_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i})\|^{2}.$ (69)
Using [54, Lemma D1], $\mathbb{E}_{u_{t}}[\mathsf{g}_{\rho}(\theta_{i})\mid\theta_{i}]=\nabla\psi_{\rho}(\theta_{i})$. Taking expectation given $\theta_{i}$
$\mathbb{E}[\|\mathcal{G}_{\Theta}(\theta_{i},\tilde{\mathsf{g}}_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}\mid\theta_{i}]\leq$
$\frac{2[\psi_{\rho}(\theta_{i})-\mathbb{E}[\psi_{\rho}(\theta_{i+1})\mid\theta_{i}]]}{\tilde{\eta}_{i}}+4\,\mathbb{E}[\|\mathsf{g}_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}]+$
$\quad+4^{-1}\mathbb{E}[\|\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}\mid\theta_{i}]+2\mathbb{E}[\|\nabla\psi_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}].$ (70)
The term involving $\nabla\psi_{\rho}(\theta_{i})-\mathsf{g}_{\rho}(\theta_{i})$ vanishes because $\mathbb{E}_{u_{t}}[\mathsf{g}_{\rho}(\theta_{i})\mid\theta_{i}]=\nabla\psi_{\rho}(\theta_{i})$ and $\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})$ is deterministic given $\theta_{i}$. Moreover, by $(a+b)^{2}\leq 2(a^{2}+b^{2})$,
$\mathbb{E}[\|\nabla\psi_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}]\leq 2\,\mathbb{E}[\|\nabla\psi_{\rho}(\theta_{i})-\mathsf{g}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}]+2\,\mathbb{E}[\|\mathsf{g}_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}].$
Substituting this into (70), we obtain
$\mathbb{E}[\|\mathcal{G}_{\Theta}(\theta_{i},\tilde{\mathsf{g}}_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}\mid\theta_{i}]\leq$
$\frac{2[\psi_{\rho}(\theta_{i})-\mathbb{E}[\psi_{\rho}(\theta_{i+1})\mid\theta_{i}]]}{\tilde{\eta}_{i}}+4\,\mathbb{E}[\|\nabla\psi_{\rho}(\theta_{i})-\mathsf{g}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}]$
$\qquad+8\,\mathbb{E}[\|\mathsf{g}_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}]+4^{-1}\mathbb{E}[\|\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}\mid\theta_{i}].$ (71)
In order to give a bound on $\mathbb{E}[\|\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}|\theta_{i}]$, we have
$\mathbb{E}[\|\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}|\theta_{i}]\leq 2\mathbb{E}[\|\mathcal{G}_{\Theta}(\theta_{i},\tilde{\mathsf{g}}_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}|\theta_{i}]$
$+2\mathbb{E}[\|\mathcal{G}_{\Theta}(\theta_{i},\tilde{\mathsf{g}}_{\rho}(\theta_{i});\tilde{\eta}_{i})-\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}|\theta_{i}]$
$\overset{(a)}{\leq}2\mathbb{E}[\|\mathcal{G}_{\Theta}(\theta_{i},\tilde{\mathsf{g}}_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}|\theta_{i}]+2\mathbb{E}[\|\tilde{\mathsf{g}}_{\rho}(\theta_{i})-\nabla\psi_{\rho}(\theta_{i})\|^{2}|\theta_{i}]$
$\leq 2\mathbb{E}[\|\mathcal{G}_{\Theta}(\theta_{i},\tilde{\mathsf{g}}_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}|\theta_{i}]$
$\qquad+4\mathbb{E}[\|\nabla\psi_{\rho}(\theta_{i})-\mathsf{g}_{\rho}(\theta_{i})\|^{2}|\theta_{i}]+4\mathbb{E}[\|\tilde{\mathsf{g}}_{\rho}(\theta_{i})-\mathsf{g}_{\rho}(\theta_{i})\|^{2}|\theta_{i}].$ (72)
where (a) follows by $\|\mathcal{G}_{\Theta}(\theta,v_{1};\eta)-\mathcal{G}_{\Theta}(\theta,v_{2};\eta)\|\leq\|v_{1}-v_{2}\|$. Incorporating (71) in (72), we get
$\mathbb{E}[\|\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}|\theta_{i}]\leq\frac{4[\psi_{\rho}(\theta_{i})-\mathbb{E}[\psi_{\rho}(\theta_{i+1})\mid\theta_{i}]]}{\tilde{\eta}_{i}}$
$+12\,\mathbb{E}[\|\nabla\psi_{\rho}(\theta_{i})-\mathsf{g}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}]+20\,\mathbb{E}[\|\mathsf{g}_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}]$
$\qquad+2^{-1}\mathbb{E}[\|\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}\mid\theta_{i}].$ (73)
Finally, we get
$\mathbb{E}[\|\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}|\theta_{i}]\leq$
$\frac{8[\psi_{\rho}(\theta_{i})-\mathbb{E}[\psi_{\rho}(\theta_{i+1})\mid\theta_{i}]]}{\tilde{\eta}_{i}}+24\,\mathbb{E}[\|\nabla\psi_{\rho}(\theta_{i})-\mathsf{g}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}]$
$\qquad+40\,\mathbb{E}[\|\mathsf{g}_{\rho}(\theta_{i})-\tilde{\mathsf{g}}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}].$ (74)
Up to this point, the analysis is standard for constrained zeroth-order methods and does not use the specific structure of $F_{\max}$.
Step 2: Controlling the directional and oracle errors. We first bound the variance coming from the random directions. Since the variables $\{\mathsf{g}_{\rho}(\theta_{i},u_{i,t})\}_{t=1}^{b_{u}}$ are i.i.d. and satisfy $\mathbb{E}_{u_{t}}[\mathsf{g}_{\rho}(\theta_{i},u_{i,t})\mid\theta_{i}]=\nabla\psi_{\rho}(\theta_{i})$, we have
$\mathbb{E}[\|\nabla\psi_{\rho}(\theta_{i})-\mathsf{g}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}]$
$=\frac{1}{b_{u}^{2}}\sum_{t=1}^{b_{u}}\mathbb{E}[\|\mathsf{g}_{\rho}(\theta_{i},u_{i,t})-\nabla\psi_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}]$
$\leq\frac{1}{b_{u}}\,\mathbb{E}[\|\mathsf{g}_{\rho}(\theta_{i},u)\|^{2}\mid\theta_{i}]$
$\leq\frac{m^{2}L_{F_{\max},1}^{2}}{b_{u}},$ (75)
where we used
$\|\mathsf{g}_{\rho}(\theta_{i},u)\|=\frac{m}{2\rho}\,|F_{\max}(\theta_{i}+\rho u)-F_{\max}(\theta_{i}-\rho u)|\leq mL_{F_{\max},1}.$
We now invoke Section˜5.1 to bound the oracle term in (D.3):
$\mathbb{E}[\|\tilde{\mathsf{g}}_{\rho}(\theta_{i})-\mathsf{g}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}]$
$\leq\frac{m^{2}}{\rho^{2}b_{u}}\sum_{t=1}^{b_{u}}\mathbb{E}[|\tilde{\psi}(\theta_{i}+\rho u_{i,t})-F_{\max}(\theta_{i}+\rho u_{i,t})|^{2}$
$\hskip 91.04872pt+|\tilde{\psi}(\theta_{i}-\rho u_{i,t})-F_{\max}(\theta_{i}-\rho u_{i,t})|^{2}]$
$\leq\frac{m^{2}}{\rho^{2}}\cdot\mathcal{O}\left(\epsilon_{v}^{2}\right),$ (76)
where the last inequality comes from Section˜5.1 with $|I|\geq 4\tilde{C}_{f}\epsilon_{v}^{-2-2\texttt{k}}$, and
$L=\mathcal{O}\left(\frac{\texttt{k}^{2}\mathrm{B}^{2}}{\epsilon_{v}^{2+2\texttt{k}}}\cdot\log^{2}(\epsilon_{v}^{-1})\right).$
By choosing $\tilde{\eta}_{i}=\eta\leq 1/L_{\psi,2}=m^{-1/2}\rho L_{F_{\max},1}^{-1}$ and integrating (D.3) and (76) into (D.3), followed by averaging over $\theta_{i}$, we obtain
$\frac{1}{N}\sum_{t=1}^{N}\mathbb{E}[\|\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho};\eta)\|^{2}]\leq\frac{8\psi_{\rho}(\theta_{0})-8\mathbb{E}[\psi_{\rho}(\theta_{N})]}{N\eta}+\mathcal{O}\left(\frac{m^{2}}{b_{u}}\right)$
$\qquad\qquad+\frac{40m^{2}}{\rho^{2}}\cdot\mathcal{O}\left(\epsilon_{v}^{2}\right)$
$\leq\mathcal{O}\left(\frac{m^{\frac{1}{2}}\cdot(\psi_{\rho}(\theta_{0})-\mathbb{E}[\psi_{\rho}(\theta_{N})])}{N\rho}+\frac{m^{2}}{b_{u}}+\frac{m^{2}}{\rho^{2}}\cdot\epsilon_{v}^{2}\right)$
$\leq\mathcal{O}\left(\frac{m\cdot\text{diam}(\Theta)}{N\rho^{2}}+\frac{m^{2}}{b_{u}}+\frac{m^{2}}{\rho^{2}}\cdot\epsilon_{v}^{2}\right),$ (77)
where the final inequality results from the Lipschitz continuity of $F_{\max}$ with constant $(\sqrt{m}\rho^{-1}L_{F_{\max},1})$ and the boundedness of domain $\Theta$.
Let us define $\hat{\theta}=\operatorname*{arg\,min}_{i\in[1:N]}\mathbb{E}[\|\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho};\eta)\|^{2}]$. Then by Jensen’s inequality, we have
$\mathbb{E}[\|\mathcal{G}_{\Theta}(\hat{\theta},\nabla\psi_{\rho};\eta)\|]=\mathcal{O}\left(\frac{m^{\frac{1}{2}}(\text{diam}(\Theta))^{\frac{1}{2}}}{\sqrt{N}\rho}+\frac{m}{\sqrt{b_{u}}}+\frac{m}{\rho}\cdot\epsilon_{v}\right).$ (78)
Let us pick $\epsilon_{v}=\rho\epsilon m^{-1}$, $N\geq m\cdot\text{diam}(\Theta)\cdot\epsilon^{-2}\rho^{-2}$, and $b_{u}\geq m^{2}\epsilon^{-2}$. Then $\mathbb{E}[\|\mathcal{G}_{\Theta}(\hat{\theta},\nabla\psi_{\rho};\eta)\|]=\mathcal{O}\left(\epsilon\right)$. We also pick $b_{u}=\mathcal{O}(m^{2}\epsilon^{-2})$. Then the total number of samples from Gibbs measure is
$N\times(b_{u}L+b_{u}|I|)=\mathcal{O}\left(\frac{m^{5+2\texttt{k}}\cdot\text{diam}(\Theta)\cdot(\texttt{k}^{2}\mathrm{B}^{2}+\tilde{C}_{f})\cdot\log^{2}\left(m\epsilon^{-1}\rho^{-1}\right)}{\epsilon^{2\texttt{k}+6}\rho^{2\texttt{k}+4}}\right).$ (79)
Appendix E Computational Complexity of Gibbs Oracle
In this section, we modify the result of Section˜5 for the setting where the Gibbs samples are generated approximately by LMC rather than drawn exactly from the Gibbs distribution.
E.1 Runtime Complexity of Algorithm˜1 by Using LMC
We first replace Algorithm˜1 by Algorithm˜3, where the only change is that the exact samples $\{\tilde{X}^{l}\}_{l=0}^{L-1}$ from the Gibbs distribution $\mu_{\theta}^{\lambda}$ are replaced by approximate samples $\{\tilde{X}_{l}^{n}\}_{l=0}^{L-1}$, each $\tilde{X}_{l}^{n}$ being produced by running an $n$-step Langevin Monte Carlo (LMC) algorithm (see Equation˜30). Our goal is to establish an updated version of Section˜5.1 tailored to Algorithm˜3.
We pick $\lambda>0$ and
Algorithm 3 $\text{PSGD-LMC}(\theta,\beta_{0})$
Input: $\theta$, $\beta_{0}$
1:for $l$ = 0 to L1L-1 do
2:  $\beta^{l+1}=\text{Proj}_{[-\mathrm{B},\mathrm{B}]}(\beta^{l}-\eta_{l}\tilde{\partial}_{\beta}\phi_{\lambda,\delta}(\theta,\beta^{l};\tilde{X}^{n}_{i}))$ $\{\tilde{X}^{n}_{i}\}_{l=0}^{L-1}$ are i.i.d samples of $L$ parallel LMC methods.]
3:end for
[…]
define the following approximate objective:
$\tilde{\phi}_{\lambda,\delta}(\theta,\beta):=\beta+\frac{1}{\delta}\mathbb{E}_{X\sim\hat{\mu}_{n}}\left[\max\{f(\theta,X)-\beta,0\}\right],$ (80)
where $\hat{\mu}_{n}$ denotes the empirical distribution induced by $n$-step LMC output. Since
$\bigl|\tilde{\phi}_{\lambda,\delta}(\theta,\beta)-\phi_{\lambda,\delta}(\theta,\beta)\bigr|\leq\frac{L_{f,1}}{\delta}\,\mathbb{W}_{1}(\hat{\mu}_{n},\mu_{\theta}^{\lambda}),$
we can invoke the LMC Wasserstein bound from Section˜5.3 to obtain
$\bigl|\tilde{\phi}_{\lambda,\delta}(\theta,\beta)-\phi_{\lambda,\delta}(\theta,\beta)\bigr|\leq\frac{L_{f,1}}{\delta}\cdot 2C_{PI}^{1/2}C^{1/2}\,\varepsilon.$ (81)
Setting $\epsilon_{v}:=\varepsilon/\delta$ with $\delta=\epsilon_{v}^{\texttt{k}}$ and running LMC for $n=\tilde{\mathcal{O}}\bigl(d\,\epsilon_{v}^{-6-6\texttt{k}}\bigr)$ iterations yields
$\bigl|\tilde{\phi}_{\lambda,\delta}(\theta,\beta)-\phi_{\lambda,\delta}(\theta,\beta)\bigr|=\mathcal{O}(\epsilon_{v}).$
Since (81) holds for every $\beta$, the same argument as in Section˜4.2 implies
$\bigl|\min_{\beta}\tilde{\phi}_{\lambda,\delta}(\theta,\beta)-\min_{\beta}\phi_{\lambda,\delta}(\theta,\beta)\bigr|=\mathcal{O}(\epsilon_{v}).$
We then define the LMC–based estimator of $\tilde{F}_{{\texttt{SQ-G}}}(\theta)$ by
$\hat{\psi}(\theta):=\phi_{\lambda,\delta}(\theta,\tilde{\beta};\hat{X}^{n}_{I})=\tilde{\beta}+\frac{1}{\delta|I|}\sum_{i\in I}\max\{f(\theta,\hat{X}^{n}_{i})-\tilde{\beta},0\},$ (82)
where $\hat{X}^{n}_{I}:=\{\hat{X}^{n}_{i}\}_{i\in I}$ are i.i.d. nn–step LMC samples from $\hat{\mu}_{n}$.
We then restate section˜5.1 by using the updates of algorithm˜3 as follows:
Lemma E.1.
Given $\theta\in\Theta$, by using Algorithm˜3 ($\text{PSGD-LMC}(\theta,\beta_{0})$), we obtain
$\mathbb{E}[|\hat{\psi}(\theta)-\tilde{F}_{{\texttt{SQ-G}}}(\theta)|^{2}\mid\theta]=\mathcal{O}(\epsilon_{v}^{2}),$ (83)
with $|I|\geq\tilde{C}_{f}\epsilon_{v}^{-2\texttt{k}-2}$ and the number of iterations of PSGD as follows
$L\times n\geq\tilde{\mathcal{O}}\left(\frac{d\texttt{k}^{2}\mathrm{B}^{2}\log^{2}(\epsilon_{v}^{-1})}{\epsilon_{v}^{8+8\texttt{k}}}\right).$
Here $\tilde{C}_{f}$ is a constant, and $\mathrm{B}$ is defined in Section˜5.1.
Proof.
The proof mirrors that of Section˜5.1; we highlight only the modifications due to the use of LMC samples and the function $\tilde{\phi}_{\lambda,\delta}$.
Step 1: PSGD error for $\tilde{\phi}_{\lambda,\delta}$. Fix $\theta\in\Theta$, and let $\tilde{\beta}$ denote the output of $\mathrm{PSGD\mbox{-}LMC}(\theta,\beta_{0})$. As in the proof of Section˜5.1, $\tilde{\phi}_{\lambda,\delta}(\theta,\cdot)$ is convex, $\sigma$–Lipschitz with $\sigma=1+\delta^{-1}$, and optimized over the interval $\mathrm{B},\mathrm{B}]$. Therefore, the high–probability PSGD guarantee (18) applies to $\tilde{\phi}_{\lambda,\delta}$ verbatim: for any $\epsilon_{v}>0$ and $\Delta\in(0,1)$,
$\mathbb{P}\Bigl(\tilde{\phi}_{\lambda,\delta}(\theta,\tilde{\beta})-\min_{\beta\in[-\mathrm{B},\mathrm{B}]}\tilde{\phi}_{\lambda,\delta}(\theta,\beta)\leq\epsilon_{v}\;\Big|\;\theta\Bigr)\;\geq\;1-\Delta,$
after at most
$L\;=\;\mathcal{O}\bigl(\mathrm{B}^{2}\sigma^{2}\epsilon_{v}^{-2}\log^{2}(\Delta^{-1})\bigr).$
Choosing $\delta=\epsilon_{v}^{\texttt{k}}$ so that $\sigma=\mathcal{O}(\epsilon_{v}^{-\texttt{k}})$ and taking $\Delta:=\min\{1,\epsilon_{v}^{2}/R^{2}\}$ as in the proof of Section˜5.1, we obtain
$\mathbb{E}\Bigl[\bigl(\tilde{\phi}_{\lambda,\delta}(\theta,\tilde{\beta})-\min_{\beta}\tilde{\phi}_{\lambda,\delta}(\theta,\beta)\bigr)^{2}\Bigm|\theta\Bigr]\;\leq\;C\,\epsilon_{v}^{2},$ (84)
for some constant $C>0$, using
$L=\mathcal{O}\bigl(\texttt{k}^{2}\,\mathrm{B}^{2}\epsilon_{v}^{-2-2\texttt{k}}\log^{2}(\epsilon_{v}^{-1})\bigr).$ (85)
Step 2: From $\tilde{\phi}_{\lambda,\delta}$ to $\phi_{\lambda,\delta}$ and $\tilde{F}_{{\texttt{SQ-G}}}$. Recall that $\tilde{F}_{{\texttt{SQ-G}}}(\theta)=\min_{\beta}\phi_{\lambda,\delta}(\theta,\beta)$. For the same $\tilde{\beta}$ as above,
$\phi_{\lambda,\delta}(\theta,\tilde{\beta})-\tilde{F}_{{\texttt{SQ-G}}}(\theta)$ $=\phi_{\lambda,\delta}(\theta,\tilde{\beta})-\min_{\beta}\phi_{\lambda,\delta}(\theta,\beta)$
$\leq\bigl|\phi_{\lambda,\delta}(\theta,\tilde{\beta})-\tilde{\phi}_{\lambda,\delta}(\theta,\tilde{\beta})\bigr|+\bigl(\tilde{\phi}_{\lambda,\delta}(\theta,\tilde{\beta})-\min_{\beta}\tilde{\phi}_{\lambda,\delta}(\theta,\beta)\bigr)$
$\quad+\bigl|\min_{\beta}\tilde{\phi}_{\lambda,\delta}(\theta,\beta)-\min_{\beta}\phi_{\lambda,\delta}(\theta,\beta)\bigr|.$
By the uniform approximation bound (81) and its consequence for minimizers (see the discussion after (81)), both $\bigl|\phi_{\lambda,\delta}(\theta,\beta)-\tilde{\phi}_{\lambda,\delta}(\theta,\beta)\bigr|$ and $\bigl|\min_{\beta}\tilde{\phi}_{\lambda,\delta}(\theta,\beta)-\min_{\beta}\phi_{\lambda,\delta}(\theta,\beta)\bigr|$ are $\mathcal{O}(\epsilon_{v})$, uniformly in $\beta$, provided $n$ is chosen as below. Combining these with (84) and using $(a+b+c)^{2}\leq 3(a^{2}+b^{2}+c^{2})$ gives
$\mathbb{E}\Bigl[\bigl(\phi_{\lambda,\delta}(\theta,\tilde{\beta})-\tilde{F}_{{\texttt{SQ-G}}}(\theta)\bigr)^{2}\Bigm|\theta\Bigr]\;\leq\;C_{2}\,\epsilon_{v}^{2},$ (86)
for some constant $C_{2}>0$, provided the LMC chain length satisfies
$n=\tilde{\mathcal{O}}\bigl(d\,\epsilon_{v}^{-6-6\texttt{k}}\bigr),$
as in the Wasserstein–convergence bound used in Section˜5.3.
Step 3: Bounding the MSE of $\hat{\psi}(\theta)$. By Jensen’s inequality and the triangle inequality,
$\mathbb{E}\bigl[|\hat{\psi}(\theta)-\tilde{F}_{{\texttt{SQ-G}}}(\theta)|^{2}\mid\theta\bigr]\leq 3\,\underbrace{\mathbb{E}\bigl[|\hat{\psi}(\theta)-\tilde{\psi}(\theta)|^{2}\mid\theta\bigr]}_{(1)}$
$+3\,\underbrace{\mathbb{E}\bigl[|\tilde{\psi}(\theta)-\phi_{\lambda,\delta}(\theta,\tilde{\beta})|^{2}\mid\theta\bigr]}_{(2)}+3\,\underbrace{\mathbb{E}\bigl[|\phi_{\lambda,\delta}(\theta,\tilde{\beta})-\tilde{F}_{{\texttt{SQ-G}}}(\theta)|^{2}\mid\theta\bigr]}_{(3)}.$ (87)
We bound the three terms as follows.
1. Conditioning on $\theta$ and $\tilde{\beta}$, using independence across $i\in I$,
$\mathbb{E}\bigl[|\hat{\psi}(\theta)-\tilde{\psi}(\theta)|^{2}\mid\theta\bigr]$ $=\mathbb{E}\Bigl[\Bigl|\frac{1}{\delta|I|}\sum_{i\in I}\bigl(h(f(\theta,\hat{X}^{n}_{i})-\tilde{\beta})-h(f(\theta,\tilde{X}^{i})-\tilde{\beta})\bigr)\Bigr|^{2}\Bigm|\theta\Bigr]$ (88)
$\leq\frac{1}{\delta^{2}|I|}\sum_{i\in I}\mathbb{E}\Bigl[\bigl(h(f(\theta,\hat{X}^{n}_{i})-\tilde{\beta})-h(f(\theta,\tilde{X}^{i})-\tilde{\beta})\bigr)^{2}\Bigm|\theta\Bigr]$ (89)
$\leq\frac{L_{f,1}^{2}}{\delta^{2}}\,\mathbb{W}_{2}^{2}(\hat{\mu}_{n},\mu_{\theta}^{\lambda}),$ (90)
[…]
$\mathbb{E}\bigl[|\hat{\psi}(\theta)-\tilde{\psi}(\theta)|^{2}\mid\theta\bigr]\;\leq\;C_{3}\,\epsilon_{v}^{2}$ (91)
for some constant $C_{3}>0$.
2. Exactly as in (59)–(60), we have
$\mathbb{E}\bigl[|\tilde{\psi}(\theta)-\phi_{\lambda,\delta}(\theta,\tilde{\beta})|^{2}\mid\theta\bigr]\;\leq\;\frac{\tilde{C}_{f}}{\delta^{2}|I|},$ (92)
for some constant $\tilde{C}_{f}>0$. With $\delta=\epsilon_{v}^{\texttt{k}}$ and $|I|\geq\tilde{C}_{f}\epsilon_{v}^{-2\texttt{k}-2}$, this is $\mathcal{O}(\epsilon_{v}^{2})$.
3. This is exactly (86):
$\mathbb{E}\bigl[|\phi_{\lambda,\delta}(\theta,\tilde{\beta})-\tilde{F}_{{\texttt{SQ-G}}}(\theta)|^{2}\mid\theta\bigr]\;\leq\;C_{2}\,\epsilon_{v}^{2}.$ (93)
Substituting (91)–(93) into (87), we obtain
$\mathbb{E}\bigl[|\hat{\psi}(\theta)-\tilde{F}_{{\texttt{SQ-G}}}(\theta)|^{2}\mid\theta\bigr]\;\leq\;C\,\epsilon_{v}^{2},$
for some constant $C>0$ independent of $\theta$ and $\epsilon_{v}$. Finally, combining (85) with $n=\tilde{\mathcal{O}}(d\,\epsilon_{v}^{-6-6\texttt{k}})$ yields
$L\times n\;=\;\tilde{\mathcal{O}}\!\left(\frac{d\,\texttt{k}^{2}\,\mathrm{B}^{2}\log^{2}(\epsilon_{v}^{-1})}{\epsilon_{v}^{8+8\texttt{k}}}\right),$
which completes the proof. ∎
E.2 Proof of ˜5.7
First, we change Algorithm˜2 to Algorithm˜4 by replacing the samples of Gibbs measure with samples of the output of parallel LMCs and changing PSGD to PSGD-LMC (see Algorithm˜3) in lines 4 and 5. Relative to the proof of ˜5.5, the directional-variance term in (D.3) is unchanged; the only new ingredient is the bound on the oracle term $\mathbb{E}[\|\mathsf{g}_{\rho}(\theta_{i})-\hat{\mathsf{g}}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}]$. We provide the proof of ˜5.7 in the following:
Algorithm 4 Projected Stochastic Zeroth-order with LMC (PSZO-LMC)
Initialization: $\beta_{0}$, $\theta_{0}$, $\lambda,\delta<1/2$, and integers $N,n$.
1:for $k$ = 0 to N1N-1 do
2:  Query i.i.d. samples of $|I|$ parallel LMC methods, denoted by $\hat{X}_{I}^{n}=\{\hat{X}_{i}^{n}\}_{i\in I}$.
3:  Sample $\{u_{k,t}\}_{t=1}^{b_{u}}$ independently and uniformly from a unit sphere.
4:  $\beta_{k,t}^{+}=\text{PSGD-LMC}(\theta_{k}+\rho u_{k,t},\beta_{0}),$ $\forall t\in[1:b_{u}]$.
5:  $\beta_{k,t}^{-}=\text{PSGD-LMC}(\theta_{k}-\rho u_{k,t},\beta_{0}),$ $\forall t\in[1:b_{u}]$.
6:  $\hat{\psi}(\theta_{k}+\rho u_{k,t})=\phi_{\lambda,\delta}(\theta_{k}+\rho u_{k,t},\beta_{k,t}^{+};\hat{X}_{I}^{N})$.
7:  $\hat{\psi}(\theta_{k}-\rho u_{k,t})=\phi_{\lambda,\delta}(\theta_{k}-\rho u_{k,t},\beta_{k,t}^{-};\hat{X}_{I}^{N})$
8:  $\hat{\mathsf{g}}_{\rho}(\theta_{k})=\frac{m}{2\rho}\cdot\frac{1}{b_{u}}\sum_{t=1}^{b_{u}}\left[\hat{\psi}(\theta_{k}+\rho u_{k,t})-\hat{\psi}(\theta_{k}-\rho u_{k,t})\right]u_{k,t}$
9:  $\theta_{k+1}=\text{Proj}_{\Theta}(\theta_{k}-\tilde{\eta}_{k}\hat{\mathsf{g}}_{\rho}(\theta_{k}))$
10:end for
Proof.
The argument mirrors the proof of ˜5.5. We briefly recall the zeroth-order part and then highlight the only new ingredient, namely the control of the LMC-induced error via Section˜E.1.
Step 1: Zeroth-order analysis (same as Theorem 5.5). As in the proof of ˜5.5, we work with the smoothed objective $\psi_{\rho}(\theta):=\mathbb{E}_{u\sim\mathrm{Unif}(\mathbb{S}^{m-1})}\big[F_{\max}(\theta+\rho u)\big].$ Using the smoothness of $\psi_{\rho}$ and the projected gradient-mapping update, the standard analysis of constrained zeroth-order methods yields (see section˜D.3 in Section˜D.3)
$\mathbb{E}\big[\|\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\tilde{\eta}_{i})\|^{2}\mid\theta_{i}\big]$ $\leq\frac{8\big(\psi_{\rho}(\theta_{i})-\mathbb{E}[\psi_{\rho}(\theta_{i+1})\mid\theta_{i}]\big)}{\tilde{\eta}_{i}}$
$\qquad\qquad+24\,\mathbb{E}\big[\|\nabla\psi_{\rho}(\theta_{i})-\mathsf{g}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}\big]$
$\qquad\qquad+40\,\mathbb{E}\big[\|\mathsf{g}_{\rho}(\theta_{i})-\hat{\mathsf{g}}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}\big],$ (94)
where
$\mathsf{g}_{\rho}(\theta_{i}):=\frac{1}{b_{u}}\sum_{t=1}^{b_{u}}\mathsf{g}_{\rho}(\theta_{i},u_{i,t}),\qquad\mathsf{g}_{\rho}(\theta,u):=\frac{m}{2\rho}\big[F_{\max}(\theta+\rho u)-F_{\max}(\theta-\rho u)\big]u,$
and $\hat{\mathsf{g}}_{\rho}(\theta_{i})$ is the stochastic gradient estimator used in Algorithm 4. Up to (E.2), the analysis is identical to Theorem 5.5; in particular,
$\mathbb{E}\big[\|\nabla\psi_{\rho}(\theta_{i})-\mathsf{g}_{\rho}(\theta_{i})\|^{2}\mid\theta_{i}\big]\leq\frac{m^{2}L_{F_{\max},1}^{2}}{b_{u}}.$
The only new task is to control $\mathbb{E}[\|\hat{\mathsf{g}}_{\rho}(\theta)-\mathsf{g}_{\rho}(\theta)\|^{2}\mid\theta]$ when $\hat{\mathsf{g}}_{\rho}$ is built from the LMC-based oracle $\hat{\psi}$.
Step 2: Bounding the LMC-induced error via section˜E.1. In Algorithm 4, the proxy gradient in line 8 is
$\hat{\mathsf{g}}_{\rho}(\theta)=\frac{m}{2\rho}\cdot\frac{1}{b_{u}}\sum_{t=1}^{b_{u}}\big[\hat{\psi}(\theta+\rho u_{t})-\hat{\psi}(\theta-\rho u_{t})\big]u_{t},$ (95)
[…]
$\mathbb{E}\big[\|\hat{\mathsf{g}}_{\rho}(\theta)-\mathsf{g}_{\rho}(\theta)\|^{2}\mid\theta\big]\leq$
$\;\frac{m^{2}}{\rho^{2}}\,\mathbb{E}\Big[\big|\hat{\psi}(\theta+\rho u_{t})-F_{\max}(\theta+\rho u_{t})\big|^{2}+\big|\hat{\psi}(\theta-\rho u_{t})-F_{\max}(\theta-\rho u_{t})\big|^{2}\Bigm|\theta\Big].$ (96)
Thus it suffices to bound $\mathbb{E}[|\hat{\psi}(\theta)-F_{\max}(\theta)|^{2}\mid\theta]$.
By Section˜E.1, for any $\theta\in\Theta$ we have
$\mathbb{E}\big[|\hat{\psi}(\theta)-\tilde{F}_{{\texttt{SQ-G}}}(\theta)|^{2}\mid\theta\big]=\mathcal{O}(\epsilon_{v}^{2}),$ (97)
provided that
[…]
with $\tilde{C}_{f}$ a constant and $\mathrm{B}$ as in Lemma 5.1. Moreover, by the uniform approximation result for the superquantile–Gibbs relaxation (see ˜4.2), we have
$\big|\tilde{F}_{{\texttt{SQ-G}}}(\theta)-F_{\max}(\theta)\big|=\mathcal{O}(\epsilon_{v})\qquad\text{for all }\theta\in\Theta.$ (98)
Combining (97) and (98) and using $(a+b)^{2}\leq 2(a^{2}+b^{2})$, we obtain
$\mathbb{E}\big[|\hat{\psi}(\theta)-F_{\max}(\theta)|^{2}\mid\theta\big]$ $\leq 2\,\mathbb{E}\big[|\hat{\psi}(\theta)-\tilde{F}_{{\texttt{SQ-G}}}(\theta)|^{2}\mid\theta\big]+2\,\big|\tilde{F}_{{\texttt{SQ-G}}}(\theta)-F_{\max}(\theta)\big|^{2}$
$=\mathcal{O}(\epsilon_{v}^{2}).$ (99)
Substituting (99) into (E.2) yields
$\mathbb{E}\big[\|\hat{\mathsf{g}}_{\rho}(\theta)-\mathsf{g}_{\rho}(\theta)\|^{2}\mid\theta\big]=\mathcal{O}\!\left(\frac{m^{2}}{\rho^{2}}\epsilon_{v}^{2}\right).$ (100)
Step 3: Convergence rate and choice of parameters. Plugging (100) into (E.2), choosing a constant stepsize $\tilde{\eta}_{i}=\eta\leq m^{-1/2}\rho L_{F_{\max},1}^{-1}$, and summing $i=0,\dots,N-1$ exactly as in the proof of Theorem 5.5 gives
$\frac{1}{N}\sum_{i=0}^{N-1}\mathbb{E}\big[\|\mathcal{G}_{\Theta}(\theta_{i},\nabla\psi_{\rho}(\theta_{i});\eta)\|^{2}\big]\;=\;\mathcal{O}\!\left(\frac{m\,\mathrm{diam}(\Theta)}{N\rho^{2}}+\frac{m^{2}}{b_{u}}+\frac{m^{2}}{\rho^{2}}\epsilon_{v}^{2}\right).$
Let $\hat{\theta}$ be the best iterate in terms of the gradient-mapping norm. By Jensen’s inequality,
$\mathbb{E}\big[\|\mathcal{G}_{\Theta}(\hat{\theta},\nabla\psi_{\rho}(\hat{\theta});\eta)\|\big]=\mathcal{O}\!\left(\sqrt{\frac{m\,\mathrm{diam}(\Theta)}{N\rho^{2}}}+\frac{m}{\sqrt{b_{u}}}+\frac{m}{\rho}\epsilon_{v}\right).$
Setting $\epsilon_{v}:=\rho\epsilon m^{-1}$ and taking $N\gtrsim m\,\mathrm{diam}(\Theta)\,\rho^{-2}\epsilon^{-2}$ and $b_{u}\gtrsim m^{2}\epsilon^{-2}$ yields
$\mathbb{E}\big[\|\mathcal{G}_{\Theta}(\hat{\theta},\nabla\psi_{\rho}(\hat{\theta});\eta)\|\big]=\mathcal{O}(\epsilon),$
so $\hat{\theta}$ is an $(\epsilon,\rho,\eta)$–generalized Goldstein stationary point of $F_{\max}$.
Each outer iteration of Algorithm 4 uses $b_{u}$ two-point evaluations of the oracle $\hat{\psi}$. Each call to $\hat{\psi}$ in turn runs PSGD-LMC with $|I|$ parallel LMC chains of length $n$. Consequently, the total number of lower-level gradient evaluations is $N\times n\times\bigl(b_{u}L+b_{u}|I|\bigr),$ and the total number of upper-level function evaluations is $N\times b_{u}\times|I|$. Substituting the choices of $N$, $n$, $L$, and $|I|$ from Section˜E.1 and the LMC mixing bound (Section˜5.3), with the choice $\epsilon_{v}=\rho\epsilon m^{-1}$, gives
$N\times n\times(b_{u}L+b_{u}|I|)$ $=\mathcal{O}\!\left(\frac{d\,m^{11+8\texttt{k}}\,[\texttt{k}^{2}\,\mathrm{B}^{2}+\tilde{C}_{f}]\log^{2}(\epsilon^{-1})}{\epsilon^{8\texttt{k}+12}\rho^{8\texttt{k}+10}}\right),$
$N\times b_{u}\times|I|$ $=\mathcal{O}\left(\frac{m^{2\texttt{k}+5}}{\rho^{2\texttt{k}+4}\epsilon^{2\texttt{k}+6}}\right),$
which is the claimed complexity bounds. ∎
Appendix F Compute-matched Comparisons under a Fixed Wall-clock Budget
This appendix complements the iteration-based results in Section˜6 by comparing methods under a fixed wall-clock time budget. As emphasized in Section˜6, PSZO-MinSel invokes a Gibbs–superquantile oracle implemented via PSGD and parallel-chain Langevin Monte Carlo, so the cost per outer iteration can differ substantially from gradient-based baselines.
Timing protocol and selection. Toy-example methods are implemented in JAX, while data hyper-cleaning methods are implemented in PyTorch; all experiments are executed on the same machine described in Section˜6. To obtain reliable timings, we exclude one-time compilation by running a short warm-up pass before timing and synchronize the device when recording timestamps. Runtimes are measured as wall-clock seconds elapsed from the start of the timed run; for JAX experiments we time a JIT-compiled step function and record timestamps with device synchronization to avoid asynchronous timing artifacts. As in Section˜6, for each method we select the best-so-far iterate within a budget $T$, $\hat{\theta}(T)\in\arg\min\{\mathcal{E}_{n}:\ t_{n}\leq T\},$ where $\mathcal{E}_{n}$ is $F(\theta_{n})$ for the toy example and the validation loss in (34) for data hyper-cleaning. All tables below report performance evaluated at $\hat{\theta}(T)$.
F.1 Toy example: fixed-runtime results
Per-iteration timing. Table˜5 reports the mean wall-clock time per outer iteration (post warm-up) for the toy example. This highlights the substantial per-iteration cost gap between PSZO-MinSel and gradient-based baselines.
$d$ V-PBGD G-PBGD RHG IAPTT-GM PSZO-MinSel
2 0.022 0.018 0.026 0.027 6.74
5 0.021 0.018 0.027 0.032 7.90
10 0.022 0.018 0.036 0.045 8.10
20 0.021 0.017 0.045 0.046 9.70
Table 5: Mean wall-clock time per outer iteration (ms) on the toy example with $\texttt{k}=d-1$.
F.1.1 Optimistic formulation: dimension sweep under fixed time
Table˜6 replicates Table˜1 under a fixed wall-clock budget $T_{\mathrm{toy}}=10$ seconds and reports the best-so-far absolute error $|\hat{\theta}(T_{\mathrm{toy}})-\theta^{\star}|$.
$d$ V-PBGD G-PBGD RHG IAPTT-GM PSZO-MinSel
[…]
Table 6: Compute-matched analogue of Table˜1: best-so-far absolute error $|\hat{\theta}(T_{\mathrm{toy}})-\theta^{\star}|$ under a fixed wall-clock budget $T_{\mathrm{toy}}=10$ s.
Under equal time, PSZO-MinSel retains a clear advantage across dimensions, consistent with the iteration-based conclusions in Section˜6.
F.1.2 Optimistic formulation: constant manifold dimension under fixed time
Table˜7 replicates Table˜2 under the same budget $T_{\mathrm{toy}}=10$ seconds, with fixed manifold dimension $\texttt{k}=1$.
$d$ V-PBGD G-PBGD RHG IAPTT-GM PSZO-MinSel
[…]
Table 7: Compute-matched analogue of Table˜2: best-so-far absolute error $|\hat{\theta}(T_{\mathrm{toy}})-\theta^{\star}|$ under $T_{\mathrm{toy}}=10$ s, with fixed $\texttt{k}=1$.
With fixed intrinsic dimension, PSZO-MinSel remains stable as $d$ increases, reinforcing that performance is governed primarily by the manifold dimension rather than the ambient dimension.
F.2 Data hyper-cleaning: fixed-runtime results
Per-iteration timing.
Table˜8 reports the mean wall-clock time per outer iteration (post warm-up) on MNIST for representative models.
Model V-PBGD G-PBGD RHG IAPTT-GM PSZO-MinSel
Linear 0.015 0.016 15.6 0.068 13.5
2-layer MLP 0.022 0.026 18.4 0.087 17.1
Table 8: Mean wall-clock time per outer iteration (s) on MNIST data hyper-cleaning (post warm-up).
RHG and PSZO-MinSel have much higher per-outer costs in this task, making fixed-runtime evaluation particularly important.
Fixed-runtime performance.
Table˜9 replicates Table˜4 under a fixed budget $T_{\mathrm{hc}}=1800$ seconds. For each method, we select the iterate achieving the lowest validation loss within $T_{\mathrm{hc}}$ and report the corresponding test accuracy and F1-score.
Table 9: Compute-matched analogue of Table˜4 under a fixed wall-clock budget $T_{\mathrm{hc}}=1800$ seconds (pollute rate $p\in\{40,60,80\}\%$): test accuracy (%) and F1 score (%), evaluated at the best-so-far iterate within the time budget (mean $\pm$ 95% CI over 10 runs).
$p$ Method Linear model 2-layer MLP
Test acc. F1 Test acc. F1
40 RHG $60.50\pm 0.32$ $45.5\pm 0.4$ $60.54\pm 0.29$ $42.3\pm 0.4$
IAPTT-GM $\mathbf{90.4\pm 0.45}$ $80.5\pm 0.5$ $\mathbf{93.76\pm 0.33}$ $76.9\pm 0.3$
G-PBGD $83.55\pm 0.57$ $71.3\pm 0.4$ $82.64\pm 0.62$ $72.3\pm 0.4$
V-PBGD $89.43\pm 0.23$ $\mathbf{83.2\pm 0.3}$ $86.56\pm 0.10$ $\mathbf{80.2\pm 0.2}$
PSZO-MinSel $88.5\pm 0.51$ $78.2\pm 0.4$ $90.3\pm 0.44$ $79.2\pm 0.2$
60 RHG $62.74\pm 0.69$ $33.8\pm 0.4$ $63.34\pm 0.42$ $35.5\pm 0.3$
IAPTT-GM $79.42\pm 0.38$ $45.4\pm 0.3$ $78.82\pm 0.45$ $50.5\pm 0.4$
G-PBGD $69.48\pm 0.42$ $40.2\pm 0.2$ $66.65\pm 0.42$ $46.5\pm 0.6$
V-PBGD $73.14\pm 0.25$ $48.7\pm 0.3$ $72.63\pm 0.28$ $52.4\pm 0.3$
PSZO-MinSel $\mathbf{80.50\pm 0.30}$ $\mathbf{57.3\pm 0.4}$ $\mathbf{82.50\pm 0.40}$ $\mathbf{56.5\pm 0.4}$
80 RHG $50.2\pm 0.4$ $24.5\pm 0.4$ $54.5\pm 0.5$ $20.9\pm 0.3$
IAPTT-GM $65.4\pm 0.54$ $30.1\pm 0.4$ $64.54\pm 0.42$ $32.9\pm 0.2$
G-PBGD $67.5\pm 0.33$ $30.9\pm 0.8$ $62.60\pm 0.34$ $30.4\pm 0.4$
V-PBGD $64.23\pm 0.22$ $32.5\pm 0.2$ $61.22\pm 0.23$ $30.9\pm 0.4$
PSZO-MinSel $\mathbf{70.34\pm 0.45}$ $\mathbf{36.7\pm 0.5}$ $\mathbf{68.43\pm 0.38}$ $\mathbf{39.1\pm 0.4}$
At low noise ($p=40\%$), PSZO-MinSel is slightly below the strongest baselines but remains comparable under the same runtime budget; as the pollute rate increases ($p=60\%,80\%$), PSZO-MinSel achieves the best performance, consistent with the benefits of trajectory-based selection under heavier corruption.